### v2

# Locally homogeneous Axiom A flows II: geometric structures for Anosov subgroups

Abstract.
Given a non-compact semisimple real Lie group $G$ and an Anosov subgroup $\Gamma$, we utilize the correspondence between $\mathbb{R}$-valued additive characters on Levi subgroups $L$ of $G$ and $\mathbb{R}$-affine homogeneous line bundles over $G/L$ to systematically construct families of non-empty domains of proper discontinuity for the $\Gamma$-action. If $\Gamma$ is torsion-free, the analytic dynamical systems on the quotients are Axiom A, and assemble into a single partially hyperbolic multiflow. Each Axiom A system admits global analytic stable/unstable foliations with non-wandering set a single basic set on which the flow is conjugate to Sambarino’s refraction flow, establishing that all refraction flows arise in this fashion. Furthermore, the $\mathbb{R}$-valued additive character is regular if and only if the associated Axiom A system admits a compatible pseudo-Riemannian metric and contact structure, which we relate to the Poisson structure on the dual of the Lie algebra of $G$.
1. Introduction
The concept of a $\Theta$-Anosov subgroup $\Gamma<G$ of a semisimple real Lie group $G$ (i.e., the image of a $\Theta$-Anosov representation with respect to a subset $\Theta\subset\Delta$ of a simple system) was introduced by Labourie [Lab06] and generalized by Guichard-Wienhard [GW12]. It proved to be so rich in examples and interesting general properties that it has become its own field of research. As the name suggests, Labourie’s original approach was intimately related to uniformly hyperbolic smooth dynamical systems (a.k.a. Anosov flows), but modern definitions [KLP17, GGKW17, BPS19, KP22] no longer appeal to flows, instead casting the $\Theta$-Anosov property in terms of growth rates of singular or eigenvalues.
If $G$ has real rank one, a subgroup $\Gamma<G$ is $\Theta$-Anosov if and only if it is convex cocompact [GW12, Thm. 5.15], and if $\Gamma$ is torsion-free the rank one Riemannian locally symmetric space is a geometric structure for $\Gamma$. The unit tangent bundle of the latter carries the geodesic flow, featuring a similar uniformly hyperbolic behavior as an Anosov flow (Smale’s Axiom A, see Definition 2.4).
In higher rank, the construction of geometric structures associated to (torsion-free) $\Theta$-Anosov subgroups $\Gamma<G$ is a key challenge in the field [Kas18, Wie18] with many results (e.g. [GW12, KLP18a, GGKW17, CS24a, NR24]), yet a systematic construction which always produces non-empty locally homogeneous manifolds has so far remained elusive. In this article, we provide such a construction. The obtained manifolds carry analytic Axiom A flows or partially hyperbolic multiflows, respectively, reaffirming the original link between Anosov subgroups and smooth hyperbolic dynamics.
In Part I [DMS24] of this series, we showed that every torsion-free projective Anosov subgroup $\Gamma<\mathrm{SL}(d,\mathbb{R})$ is the monodromy group of a locally homogeneous real analytic contact Axiom A dynamical system $(\mathcal{M},\phi^{t})$ with a unique basic hyperbolic set $\mathcal{K}\subset\mathcal{M}$. With the modern toolbox of smooth hyperbolic flows now applicable, we proved exponential mixing for all Gibbs measures (assuming $\Gamma<\mathrm{SL}(d,\mathbb{R})$ is irreducible) and established meromorphic continuation of Ruelle zeta functions. The analog of $(\mathcal{M},\phi^{t})$ when $\Gamma<G$ is a torsion-free $\Theta$-Anosov subgroup is a family of Axiom A systems $(\mathcal{M}_{\mathbf{b}},\phi_{\mathbf{b}}^{t})$ indexed by additive characters $\mathbf{b}:L_{\Theta}\rightarrow\mathbb{R}$ in an open convex cone in $\mathrm{Hom}(L_{\Theta},\mathbb{R})\simeq\mathbb{R}^{\lvert\Theta\rvert}$, where $L_{\Theta}<G$ is the standard Levi subgroup associated to $\Theta\subset\Delta$.
It is well-known (see e.g. [GW12, Sec. 9], [GGKW17, Sec. 3]) that every $\Theta$-Anosov subgroup $\Gamma<G$ becomes projective Anosov after composition with a suitable linear representation $G\rightarrow\mathrm{SL}(d,\mathbb{R})$, the latter of which are organized by the integer lattice of fundamental weights. If the differential $d_{e}\mathbf{b}$ at the identity lies at suitable points in the weight lattice, the Axiom A system $\mathcal{M}_{\mathbf{b}}$ we will construct also appears as an invariant submanifold in the corresponding locally homogeneous $(\mathrm{SL}(d,\mathbb{R}),\mathbb{L})$-manifold constructed in Part I [DMS24]. But, even if $G=\mathbf{G}(\mathbb{R})$ where $\mathbf{G}$ is a semisimple algebraic group defined over $\mathbb{\mathbb{Q}},$ a generic character $\mathbf{b}\in\mathrm{Hom}(L_{\Theta},\mathbb{R})$ is a real analytic, non-algebraic homomorphism and we cannot obtain $\mathcal{M}_{\mathbf{b}}$ from Part I by using linear representations.
Sambarino [Sam24] used additive characters on $L_{\Theta}$ to construct families of so-called refraction flows associated to any $\Theta$-Anosov subgroup, and we will show that every refraction flow is bi-Hölder conjugate to the restriction of the Axiom A flow $\phi_{\mathbf{b}}^{t}$ to its unique basic hyperbolic set $\mathcal{K}_{\mathbf{b}}\subset\mathcal{M}_{\mathbf{b}}$, in particular arriving at the same open convex cone of admissible parameters $\mathbf{b}\in\mathrm{Hom}(L_{\Theta},\mathbb{R})$ as Sambarino.
Our starting point is the observation that the differential $d_{e}\mathbf{b}$ canonically extends to a linear form $\beta:\mathfrak{g}\rightarrow\mathbb{R},$ and the coadjoint $G$-orbit of $\beta$ is generically isomorphic to $G/L_{\Theta}.$ Simultaneously, $\mathbf{b}:L_{\Theta}\rightarrow\mathbb{R}$ defines a homogeneous $\mathbb{R}$-affine line bundle $\mathbb{L}_{\mathbf{b}}$ over this coadjoint orbit, and it is in the total space of $\mathbb{L}_{\mathbf{b}}$ where we shall construct a non-empty domain of proper discontinuity for the $\Gamma$-action. We now turn to the details of this construction and the statements of our main results.
Throughout, $G$ denotes a non-compact real semisimple Lie group with Lie algebra ${\bf\mathfrak{g}}$ and $\Theta\subset\Delta$ a non-empty subset of simple restricted roots defining associated opposite proper parabolic subgroups $P_{\Theta}^{+},P_{\Theta}^{-}<G$ with standard Levi intersection $L_{\Theta}=P_{\Theta}^{+}\cap P_{\Theta}^{-}$.
1.1. Dynamical $(G,X)$-structures
Here we give a short description of the homogeneous spaces studied in this paper and the attached geometric and dynamical structures. Their construction can be thought of as reverse-engineering the unit tangent bundle $T^{1}\mathbb{X}$ of a rank-one symmetric space $\mathbb{X}$ from the projections $T^{1}\mathbb{X}\to\partial_{\infty}\mathbb{X}$ sending a unit tangent vector $v\in T^{1}_{x}\mathbb{X}$ to the future and past endpoints $v^{\pm}\in\partial_{\infty}\mathbb{X}$ of the geodesic initiated by $v$. We replace the visual boundary $\partial_{\infty}\mathbb{X}$ with a pair of opposite flag manifolds ±$\mathcal{F}^{\pm}\simeq G/P^{\pm}_{\Theta}$. The image in $\partial_{\infty}{\mathbb{X}}\times\partial_{\infty}{\mathbb{X}}$ of the projection $v\mapsto(v^{+},v^{-})$ will be replaced with the transverse flag space ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}\subset{\mathcal{F}}^{+}\times{\mathcal{F}}^{-}$, the unique open $G$-orbit in +×${\mathcal{F}}^{+}\times{\mathcal{F}}^{-}\simeq G/P^{+}_{\Theta}\times G/P^{+}_{\Theta}$ which can be equivalently described as the homogeneous space +×${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}\simeq G/L_{\Theta}$.
1.1.1. Hyperbolic homogeneous flows
By a homogeneous flow on a homogeneous space $X=G/H$ we mean a flow $\phi^{t}:X\to X$ defined for all $t\in\mathbb{R}$ that commutes with the $G$-action. We will study the case where the orbit space of $\phi^{t}$ is the homogeneous space $G/L_{\Theta}$, leading us to consider subgroups $H_{\mathbf{b}}:=\ker\mathbf{b}$ for some non-zero additive character $\mathbf{b}\in\mathrm{Hom}(L_{\Theta},\mathbb{R})$. From the structure of Levi subgroups (see Section 3.1.1), we find a canonical choice of an element $Z_{\mathbf{b}}\in\mathfrak{z}(\mathfrak{l}_{\Theta})$ in the center of the Lie algebra of $L_{\Theta}$ such that $d_{e}\mathbf{b}(Z_{\mathbf{b}})=1$. This allows us to equip the homogeneous space ${\mathbb{L}}_{\mathbf{b}}:=G/H_{\mathbf{b}}$ with the flow
$\phi^{t}_{\mathbf{b}}:\left\{\begin{array}[]{ccc}{\mathbb{L}}_{\mathbf{b}}&\to&{\mathbb{L}}_{\mathbf{b}}\\ gH_{\mathbf{b}}&\mapsto&g\exp(tZ_{\mathbf{b}})H_{\mathbf{b}},\end{array}\right.$
which is a right action of the one-parameter subgroup $A_{\mathbf{b}}:=\exp(\mathbb{R}Z_{\mathbf{b}})<Z(L_{\Theta})$.
From the embedding of $G/L_{\Theta}$ into $G/P^{+}_{\Theta}\times G/P^{-}_{\Theta}$, we obtain a decomposition
${\rm T}{\mathbb{L}}_{\mathbf{b}}=E^{-}\oplus E^{0}\oplus E^{+},$ (1.1)
where the rank one bundle $E^{0}$ is tangent to the flow and $E^{\pm}$ is sent isomorphically onto $T(G/P^{\pm}_{\Theta})$ by the differential of the projection ${\mathbb{L}}_{\mathbf{b}}=G/H_{\mathbf{b}}\to G/L_{\Theta}$ (see Section 8.3.2).
Definition 1.1.
The pair $({\mathbb{L}}_{\mathbf{b}},\phi_{\mathbf{b}}^{t})$ will be referred to as a hyperbolic homogeneous flow.
1.1.2. Multiflows
Hyperbolic homogeneous flows with a given orbit space $G/L_{\Theta}$ form a family indexed by $\mathrm{Hom}(L_{\Theta},\mathbb{R})$. The Levi subgroup $L_{\Theta}$ splits as a direct product $L=A_{\Theta}M_{\Theta}$, where $A_{\Theta}<Z(L_{\Theta})$ is isomorphic to the abelian Lie group $\mathbb{R}^{|\Theta|}$, $M_{\Theta}$ is a reductive subgroup, and the restriction map $\mathrm{Hom}(L_{\Theta},\mathbb{R})\to\mathrm{Hom}(A_{\Theta},\mathbb{R})$ is an isomorphism. Equivalently, these subgroups can be described as factors in the Langlands decompositions $P^{\pm}_{\Theta}=M_{\Theta}A_{\Theta}N^{\pm}_{\Theta}$. The homogeneous space $\mathbb{W}_{\Theta}:=G/M_{\Theta}$ comes with an action of $A_{\Theta}\simeq\mathbb{R}^{|\Theta|}$ on the right given by $(gM_{\Theta})\cdot h:=ghM_{\Theta}$ for $h\in A_{\Theta}$. For $\mathbf{b}\in\mathrm{Hom}(L_{\Theta},\mathbb{R})\setminus\{0\}$, the normal subgroup $M_{\Theta}\lhd H_{\mathbf{b}}$ satisfies $H_{\mathbf{b}}/M_{\Theta}\simeq A_{\Theta}/A_{\mathbf{b}}\simeq\mathbb{R}^{|\Theta|-1}$, and defining the projection $p_{\mathbf{b}}:\mathbb{W}_{\Theta}=G/M_{\Theta}\to G/H_{\mathbf{b}}={\mathbb{L}}_{\mathbf{b}},$ we find that
$p_{\mathbf{b}}(x\cdot h)=\phi_{\mathbf{b}}^{\mathbf{b}(h)}\big{(}p_{\mathbf{b}}(x)\big{)}\,\qquad\forall\,x\in\mathbb{W}_{\Theta},h\in A_{\Theta}.$
Definition 1.2.
The right $A_{\Theta}$-action on $\mathbb{W}_{\Theta}=G/M_{\Theta}$ will be referred to as a multiflow.
Dynamically, the fact that $p_{\mathbf{b}}:\mathbb{W}_{\Theta}\to{\mathbb{L}}_{\mathbf{b}}$ is a principal $H_{\mathbf{b}}/M_{\Theta}\simeq A_{\Theta}/A_{\mathbf{b}}\simeq\mathbb{R}^{|\Theta|-1}$-bundle means that ${\mathbb{L}}_{\mathbf{b}}$ is obtained from $\mathbb{W}_{\Theta}$ by “quotienting out all but one flow direction.”
Perhaps surprisingly, for a $\Theta$-Anosov subgroup $\Gamma<G,$ we will see that there exists a non-empty domain of proper discontinuity $\widetilde{\mathcal{N}}\subset\mathbb{W}_{\Theta}$ such that the $\mathbb{R}^{\lvert\Theta\rvert-1}\simeq A_{\Theta}/A_{\mathbf{b}}$-action is proper on the $\Gamma$-quotient $\Gamma\backslash\widetilde{\mathcal{N}}$ and the quotient $\Gamma\backslash\widetilde{\mathcal{N}}/\mathbb{R}^{\lvert\Theta\rvert-1}$ is an Axiom A system.
1.1.3. $(G,X)$-structures for Anosov subgroups
Given a $G$-homogeneous space $X=G/H$ and a manifold $\mathcal{M}$, a $(G,X)$-structure on $\mathcal{M}$ is a maximal atlas of $X$-valued charts on $\mathcal{M}$ such that the transition maps are $G$-valued and locally constant. A $(G,X)$-structure on $\mathcal{M}$ can be equivalently defined by a holonomy homomorphism $\rho_{\mathrm{hol}}:\pi_{1}(\mathcal{M})\to G$ and a $\rho_{\mathrm{hol}}$-equivariant local diffeomorphism $\mathrm{Dev}:\widetilde{\mathcal{M}}\to X$ called the developing map, where $\widetilde{\mathcal{M}}$ is any universal cover of ${\mathcal{M}}$ realizing $\pi_{1}(\mathcal{M})$ as the group of deck transformations.
Starting with a discrete subgroup $\Gamma<G$, one way of producing a $(G,X)$-manifold is to find a non-empty open $\Gamma$-invariant subset $U\subset X$ on which the $\Gamma$-action is free and properly discontinuous. The quotient manifold $\mathcal{M}_{\Gamma}:=\Gamma\backslash U$ carries a $(G,X)$-structure which we call uniformized (this is referred to as “Type $\mathrm{U}$” in [Kas18]).
Any $(G,X)$-manifold inherits all local structures on $X$ preserved by $G$. In the case of $X={\mathbb{L}}_{\mathbf{b}}$ for some $\mathbf{b}\in\mathrm{Hom}(L_{\Theta},\mathbb{R})\setminus\{0\}$, this means that a $(G,{\mathbb{L}}_{\mathbf{b}})$-manifold comes with a local flow, i.e. a vector field. Similarly, a $(G,\mathbb{W}_{\Theta})$-structure defines a local $A_{\Theta}$-action, i.e. a family of $|\Theta|$ commuting vector fields.
Definition 1.3.
A $(G,X)$-structure with $X={\mathbb{L}}_{\mathbf{b}}$ or $X=\mathbb{W}_{\Theta}$ is called dynamically complete if the induced vector fields are complete.
In the uniformized case, this means that the open subset $U\subset{\mathbb{L}}_{\mathbf{b}}$ (resp. $U\subset\mathbb{W}_{\Theta}$) is $\phi_{\mathbf{b}}^{t}$-invariant (resp. $A_{\Theta}$-invariant). The history of $(G,X)$-structures associated to Anosov subgroups is rich and well documented, see [Kas18, Wie18, CS24a] and references therein. To date and not including [DMS24], dynamical $(G,X)$-structures for higher rank Anosov subgroups do exist (e.g. [GW08]) but are rather rare among known examples. We now explain that dynamical $(G,X)$-structures with strong hyperbolicity properties exist in abundance.
1.2. Results
We now consider a $\Theta$-Anosov subgroup $\Gamma<G$. The dynamics of the $\Gamma$-action on the flow space ${\mathbb{L}}_{\mathbf{b}}$ will depend on the character $\mathbf{b}\in\mathrm{Hom}(L_{\Theta},\mathbb{R})$, and especially on its relation to the $\Theta$-limit cone $\mathcal{L}_{\Theta}(\Gamma)\subset{\bf\mathfrak{a}}_{\Theta}^{+}$ (the projection of the Benoist limit cone $\mathcal{L}(\Gamma)$ to the Lie algebra ${\bf\mathfrak{a}}_{\Theta}$ of $A_{\Theta}$, see Section 7.2). Consider the dual cone
$\mathcal{L}_{\Theta}(\Gamma)^{*}=\left\{\beta\in{\bf\mathfrak{a}}_{\Theta}^{*}\,\middle|\,\beta|_{\mathcal{L}_{\Theta}(\Gamma)}\geq 0\right\},$
whose interior $\mathrm{Int}(\mathcal{L}_{\Theta}(\Gamma)^{*})$ is non-empty because $\mathcal{L}_{\Theta}(\Gamma)$ is acute, and identify the differential $d_{e}\mathbf{b}\in\mathfrak{l}_{\Theta}^{*}$ with its restriction to ${\bf\mathfrak{a}}_{\Theta}\subset\mathfrak{l}_{\Theta}$, where $\mathfrak{l}_{\Theta}$ is the Lie algebra of $L_{\Theta}$.
1.2.1. Uniformized $(G,X)$-structures
Our first result establishes the existence of $(G,X)$-structures and hyperbolic homogeneous flows, generalizing [DMS24, Thm. A(1)].
Theorem A.
Let $\Gamma<G$ be $\Theta$-Anosov, and let $\mathbf{b}\in\mathrm{Hom}(L_{\Theta},\mathbb{R})$. If $d_{e}\mathbf{b}\in\mathrm{Int}(\mathcal{L}_{\Theta}(\Gamma)^{*})$, then there is a non-empty $\Gamma\times\phi_{\mathbf{b}}^{t}$-invariant open set ~$\widetilde{\mathcal{M}}_{\mathbf{b}}\subset{\mathbb{L}}_{\mathbf{b}}$ on which $\Gamma$ acts properly discontinuously, yielding a dynamically complete uniformized $(G,{\mathbb{L}}_{\mathbf{b}})$-manifold $\mathcal{M}_{\mathbf{b}}:=\Gamma\backslash\widetilde{\mathcal{M}_{\mathbf{b}}}$ when $\Gamma$ is torsion-free.
It is to be noted that the action of $\Gamma$ on ${\mathbb{L}}_{\mathbf{b}}$ is almost never proper, the only notable exception being the case of a group $G$ of real rank one (then ${\mathbb{L}}_{\mathbf{b}}\simeq T^{1}{\mathbb{X}}$). The fact that the multiflow space $\mathbb{W}_{\Theta}$ fibers over ${\mathbb{L}}_{\mathbf{b}}$ immediately yields the following consequence:
Corollary B.
Let $\Gamma<G$ be $\Theta$-Anosov. There is a non-empty $\Gamma\times A_{\Theta}$-invariant open set $\widetilde{\mathcal{N}}\subset\mathbb{W}_{\Theta}$ on which $\Gamma$ acts properly discontinuously, yielding a dynamically complete uniformized $(G,\mathbb{W}_{\Theta})$-manifold $\mathcal{N}:=\Gamma\backslash\widetilde{\mathcal{N}}$ when $\Gamma$ is torsion-free.
For each $\mathbf{b}\in\mathrm{Hom}(L_{\Theta},\mathbb{R})$ with $d_{e}\mathbf{b}\in\mathrm{Int}(\mathcal{L}_{\Theta}(\Gamma)^{*})$, the projection $\mathcal{N}\rightarrow\mathcal{M}_{\mathbf{b}}$ induced by the canonical projection $\mathbb{W}_{\Theta}\to{\mathbb{L}}_{\mathbf{b}}$ is a right principal $\mathbb{R}^{\lvert\Theta\rvert-1}\simeq A_{\Theta}/A_{\mathbf{b}}$-bundle.
The quotient manifolds $\mathcal{M}_{\mathbf{b}}$ are usually non-compact, but we will see that in each case the closure of the set of periodic orbits is compact and equal to the non-wandering set of the flow. A helpful example is that of quasi-Fuchsian surface groups $\Gamma<\mathrm{PSL}(2,\mathbb{C}),$ where the locally homogeneous manifolds are non-compact but have compact cores.
In real rank one, or more generally when $|\Theta|=1$, the spaces ${\mathbb{L}}_{\mathbf{b}}$ and $\mathbb{W}_{\Theta}$ are identical. In higher rank, the situation when $\Theta$ is the entire set $\Delta$ of simple restricted roots stands out because of the properness of the $G$-action on $\mathbb{W}_{\Delta}$. In other situations, a $\Theta$-Anosov subgroup $\Gamma$ typically does not act properly on $\mathbb{W}_{\Theta}$. The recent preprint [KT24] proving the sharpness conjecture gives a systematic analysis for cocompact actions of general discrete subgroups which includes some examples of Anosov subgroups.
The homogeneous space $\mathbb{W}_{\Theta}=G/M_{\Theta}$ often appears in the study of Anosov subgroups in the form of a Hopf parametrization $\mathbb{W}_{\Theta}\simeq G/L_{\Theta}\times\mathfrak{a}_{\Theta}$ with a cocycle defining the $G$-action on the right hand side, similarly for ${\mathbb{L}}_{\mathbf{b}}\simeq G/L_{\Theta}\times\mathbb{R}.$ The novelty in Theorem A and Corollary B is the discovery of the domains of proper discontinuity ~$\widetilde{\mathcal{M}}_{\mathbf{b}}\subset{\mathbb{L}}_{\mathbf{b}}$, $\widetilde{\mathcal{N}}\subset\mathbb{W}_{\Theta}.$
1.2.2. Dynamics of quotient flows
Next, we turn to the dynamics of the flows defined by $(G,{\mathbb{L}}_{\mathbf{b}})$-structures obtained in Theorem A, establishing Smale’s Axiom A. For the definition of the latter and other dynamical terminology used in the following, see Section 2.1.
Theorem C.
Let $\Gamma<G$ be a torsion-free $\Theta$-Anosov subgroup, and let $\mathbf{b}\in\mathrm{Hom}(L_{\Theta},\mathbb{R})$ be such that $d_{e}\mathbf{b}\in\mathrm{Int}(\mathcal{L}_{\Theta}(\Gamma)^{*})$. Consider the $(G,{\mathbb{L}}_{\mathbf{b}})$-manifold ${\mathcal{M}}_{\mathbf{b}}$ with its flow $\phi_{\mathbf{b}}^{t}$ obtained from Theorem A.
  1. (1)
    The real analytic flow $\phi^{t}_{\mathbf{b}}:{\mathcal{M}}_{\mathbf{b}}\to{\mathcal{M}}_{\mathbf{b}}$ is an Axiom A flow.
  2. (2)
    The non-wandering set of $\phi^{t}_{\mathbf{b}}$ consists of a unique basic set $\mathcal{K}_{\mathbf{b}}\subset{\mathcal{M}}_{\mathbf{b}}$, and the splitting of $T{\mathcal{M}}_{\mathbf{b}}|_{\mathcal{K}_{\mathbf{b}}}$ into stable/neutral/unstable subbundles provided by the Axiom A property is given by first descending the $G$-invariant splitting (1.1) to the global real analytic $\phi_{\mathbf{b}}^{t}$-invariant splitting $T{\mathcal{M}}_{\mathbf{b}}=E_{\Gamma,\mathbf{b}}^{-}\oplus E_{\Gamma,\mathbf{b}}^{0}\oplus E_{\Gamma,\mathbf{b}}^{+}$, and then restricting to the basic hyperbolic set $\mathcal{K}_{\mathbf{b}}.$
  3. (3)
    The real analytic $\phi_{\mathbf{b}}^{t}$-invariant $1$-form $\tau_{\mathbf{b}}$ which vanishes on $E_{\Gamma,\mathbf{b}}^{-}\oplus E_{\Gamma,\mathbf{b}}^{+}$ and is equal to $1$ on the generator of $\phi_{\mathbf{b}}^{t}$ is a contact $1$-form if and only if $\mathrm{B}_{\bf\mathfrak{g}}(d_{e}\mathbf{b},w_{\alpha})\neq 0$ for every fundamental weight $w_{\alpha}$ with $\alpha\in\Theta$, where $\mathrm{B}_{\bf\mathfrak{g}}$ is the Killing form.
For each character $\mathbf{b}\in\mathrm{Hom}(L_{\Theta},\mathbb{R})$ as in Theorem C, Sambarino [Sam24] defined a model of the Gromov geodesic flow of $\Gamma$ given by the so-called refraction flow $\psi_{\mathbf{b}}^{t}$ (see Section 7.3). Our locally homogeneous Axiom A flows are extensions of refraction flows:
Theorem D.
In the situation of Theorem C, the restriction of the Axiom A flow $\phi^{t}_{\mathbf{b}}:{\mathcal{M}}_{\mathbf{b}}\to{\mathcal{M}}_{\mathbf{b}}$ to its basic set $\mathcal{K}_{\mathbf{b}}\subset\mathcal{M}_{\mathbf{b}}$ is bi-Hölder-conjugate to the refraction flow $\psi_{\mathbf{b}}^{t}$.
Theorems C and D are direct generalizations of [DMS24, Theorems A, B], which treated projective Anosov subgroups of $\operatorname{SL}(d,\mathbb{R})$. More examples will be considered in Section 11. One of the differences when moving to this general setting is the condition imposed on the parameter $\mathbf{b}$ for the flow to be contact. From the perspective of refraction flows, this condition is superfluous and our results imply that any refraction flow can be realized as the restriction of a contact Axiom A flow to its non-wandering set (see Remark 9.2).
1.2.3. Ruelle zeta functions
Let $\lambda:G\to{\bf\mathfrak{a}}^{+}$ denote the Jordan projection with respect to a chosen positive Weyl chamber (see Section 3.7.1). Given a $\Theta$-Anosov subgroup $\Gamma<G$ and $\mathbf{b}\in\mathrm{Hom}(L_{\Theta},\mathbb{R})$ as in Theorem C with $d_{e}{\mathbf{b}}$ canonically extended to a linear form on ${\bf\mathfrak{g}}$, the non-trivial primitive periodic orbits of the flow $\phi_{\mathbf{b}}^{t}$ are in bijection with primitive hyperbolic conjugacy classes in $\Gamma$ and the period of $[\gamma]$ is equal to $d_{e}{\mathbf{b}}(\lambda(\gamma)).$ Since the flow is real analytic, results of Fried [Fri95] imply the meromorphic continuation of Ruelle zeta functions twisted by linear representations, which immediately yields:
Theorem E.
Let $k\in\mathbb{N}$ and $\rho:\Gamma\rightarrow\mathrm{GL}(k,\mathbb{C})$ a homomorphism. In the situation of Theorem C, the formal Euler product
$\zeta_{\mathbf{b},\rho}(s)=\prod_{[\gamma]\in[\Gamma]_{\mathrm{prim}}}\mathrm{det}\left(\mathbf{1}_{k}-e^{-sd_{e}{\mathbf{b}}(\lambda(\gamma))}\rho([\gamma])\right)^{-1}$
coverges for $\mathrm{Re}(s)\gg 1$ and admits a meromorphic continuation to $s\in\mathbb{C}.$ Here $[\Gamma]_{\mathrm{prim}}$ is the set of conjugacy classes of non-trivial primitive elements in $\Gamma$ and the appearance of $\rho([\gamma])$ is only through its conjugation invariant characteristic polynomial.
Setting $k=1$ and $\rho=1$ recovers the usual Ruelle zeta function. Alternatively, instead of [Fri95], we can also directly apply results of Dyatlov-Guillarmou [DG18a] and Borns-Weil-Shen [BWS21] giving the meromorphic continuation of Ruelle zeta functions with arbitrary $C^{\infty}$-weights. In the setting $k=1$ and $\rho=1$, Theorem E was proved for projective Anosov closed surface subgroups by Pollicott-Sharp [PS24], and shortly afterwards for all projective Anosov subgroups in [DMS24], where it was deduced analogously from [DMS24, Theorems A, B] using [DG18a, BWS21]. Theorem E gives the full unconditional result for all $\Theta$-Anosov subgroups $\Gamma<G$ and what we call admissible parameters $\mathbf{b}\in\mathrm{Hom}(L_{\Theta},\mathbb{R})$.
When $\Gamma<G$ is Zariski-dense, Chow-Sarkar [CS24b] recently elaborated the estimates necessary to prove a type of local non-integrability condition (uniformly in the admissible parameter $\mathbf{b}$) for Sambarino’s refraction flows, leading to a proof that the usual Ruelle zeta functions $\zeta_{\mathbf{b}}$ have a zero/pole-free strip to the left of their simple pole at the topological entropy $h_{\mathrm{top}}(\psi_{\mathbf{b}}^{t})$, thus establishing exponential mixing for the measure of maximal entropy. Strictly speaking, this result is only true when $\mathbf{b}$ satisfies the hypotheses of Theorem C.(3),.(3),. ( 3 ) , though in general one may pass to a smaller subset of roots and obtain the same result (see [CS24b, Remark $1.3$]).
Similar results were previously obtained for irreducible projective Anosov subgroups and arbitrary Gibbs measures in [DMS24], where the local non-integrability condition is geometrically manifested by the existence of a contact structure on the ambient Axiom A system. In general, Theorem 3 shows that under the failure of Theorem C.(3).(3). ( 3 ) and therefore the failure of the Axiom A system to be contact, there is a canonical contact reduction (e.g. quotient), which reveals that all refraction flows may be embedded in a contact Axiom A system, and provides a differential-geometric mechanism for the local non-integrability in consonance with the results of [CS24b]. Put simply, the success/failure of local non-integrability is exactly captured by whether the ambient Axiom A system is contact, and the fact that the contact property can be recovered via contact reduction (Theorem 3) is in harmony with [CS24b, Remark $1.3$]. Looking towards the future, exponential mixing for arbitrary Gibbs measure and with no Zariski-density/irreducibility hypothesis on $\Gamma<G$ remains an open problem; an analysis in the style of [DMS24] utilizing the work of Stoyanov [Sto11] is one promising avenue of attack.
1.3. Further questions and outlook
As noted above, the contact property of the Axiom A flows obtained in Theorem C is a strong version of the non-integrability condition promoted by Dolgopyat for the study of rates of mixing. Combined with the real analyticity of all dynamical foliations, these tools should allow the construction of very refined anisotropic Banach spaces in which to study transfer operators. In particular, we have in mind a mixture of the analytic/holomorphic approach of Rugh [Rug96] and Fried [Fri95], approaches making key use of a contact structure (e.g. [Liv04], [Sto25]), and microlocal approaches for non-compact systems as in [DG18a]. One would ultimately like to extend these results to allow for non-smooth Hölder continuous potentials, at least those that appear as reparametrizations in the work of Sambarino (see [Sam24]).
By applying the full technical machinery of microlocal analysis from [DG18b, DG18a, GMT21], Theorem E can be accompanied by much more comprehensive, and in parts more technical, results on the existence of discrete spectra of Ruelle-Pollicott resonances with associated (co-)resonant states (c.f. [DMS24, Thm. F]). This will be the subject of future works. We are aware of forthcoming independent work by Guedes Bonthonneau-Lefeuvre-Weich [BLW25] establishing a spectral theory for the multiflows of Corollary B.
Regarding non-Anosov subgroups $\Gamma<G,$ the larger class of $\Theta$-transverse subgroups (see [BCZZ24, CZZ24, KOW24, KOW25] and the survey [CZZ25]) includes all relatively Anosov subgroups and all discrete subgroups in real rank one. They are essentially defined as the largest class of subgroups for which there exists a pair of $\Gamma$-invariant transverse $\Theta$-limit sets $\Lambda_{\Gamma}^{\pm}\subset G/P_{\Theta}^{\pm}.$ This class of groups was initially identified by Kapovich-Leeb-Porti [KLP17] where they were called $\tau_{mod}$-regular, antipodal subgroups. It would be interesting to know whether every $\Theta$-transverse subgroup $\Gamma<G$ acts properly discontinuously on a non-empty open subset $\widetilde{\mathcal{N}}\subset\mathbb{W}_{\Theta}$ in the $\Theta$-multiflow space $\mathbb{W}_{\Theta}$, in some hyperbolic flow space ${\mathbb{L}}_{\mathbf{b}}$, or even possibly in some intermediate quotients of rank $1<k<|\Theta|$.
The existence of a $(G,X)$-structure of a given type on a manifold can require specific topological properties, leaving us to wonder if the geometric structures constructed in this article can impose restrictions on which hyperbolic groups admit Anosov representations (for a given $G$ and $\Theta\subset\Delta$). This is a notoriously difficult problem that we will not address in this paper. See [CT20, Tso20, Dey25, DGR24, KT25] for further information.
1.4. Structure of the paper and proof strategy
After a brief Background Section 2, we turn in Section 3 to Lie theoretic concepts such as flag manifolds, transverse flag spaces, and the classification of parabolic and Levi subgroups. Crucial preparations for the main proofs are made in Section 3.8, where we consider linear actions on tangent spaces of flag manifolds. Section 4 is devoted to the general concept of affine line bundles and the language of period functions and cocycles. This is then applied in Section 5 to density and multidensity bundles, the latter forming a graded analog of the usual notion of $s$-densities. While the flow spaces $\mathbb{L}_{\mathbf{b}}$ admit a very simple description in terms of coset spaces as above, our proofs are not algebraic in spirit and do not proceed in terms of this description. Instead, we construct specific differential geometric models of the space $\mathbb{L}_{\mathbf{b}}$ in terms of multidensities, which is the content of the dedicated Section 6. Anosov subgroups enter the picture in Section 7, where we prove our first main Theorem 1 and make the connection to Sambarino’s refraction flows in Section 7.3. Together with Theorem 2 and Lemma 8.3 from the subsequent Section 8 focusing on the dynamics on the $\Gamma$-quotients, this proves Theorems A, C(1), C(2) and D. In Section 9 we study in detail various geometric properties of our flow spaces $\mathbb{L}_{\mathbf{b}}$, with Theorem 3 giving the proof of (a more precise statement than) Theorem C(3). The passage to multiflow spaces is made in Section 10, where Corollary B is proved as Corollary 10.3. We take special linear groups as a family of running examples throughout the paper, which is complemented by a final Examples Section 11 describing our general constructions in concrete situations.
Acknowledgments
This project has received funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Priority Program (SPP) 2026 “Geometry at Infinity” as well as Project-ID 491392403 – TRR 358. B.D. thanks Tobias Weich for countless stimulating discussions. D.M. is thankful for the hospitality of IHES where he stayed during part of this work. A.S. thanks Beatrice Pozzetti and Thi Dang for numerous conversations about Anosov subgroups and Weyl chamber flows which led to the consideration of homogeneous affine line bundles over transverse flag manifolds.
2. Background
This section represents a continuation of the corresponding section [DMS24, Section 2] in Part I of this series. While we do (re-)introduce the required notation and preliminaries used in this paper to make it self-contained, we keep the repetitions minimal.
2.1. Dynamics
Let ${\mathcal{M}}$ be a metrizable topological space and $\phi^{t}:{\mathcal{M}}\to{\mathcal{M}}$ a continuous flow defined for all $t\in\mathbb{R}$ which has no fixed points.
Definition 2.1.
  • The non-wandering set ${\mathcal{NW}}(\phi^{t})$ of the flow $\phi^{t}$ is the set of all points $x\in{\mathcal{M}}$ for which there are sequences $x_{N}\to x$ in ${\mathcal{M}}$ and $t_{N}\to+\infty$ in $\mathbb{R}$ such that $\phi^{t_{N}}(x_{N})\to x$.
  • The trapped set ${\mathcal{T}}(\phi^{t})$ of the flow $\phi^{t}$ is the set of all points $x\in{\mathcal{M}}$ whose $\phi^{t}$-orbits are relatively compact in ${\mathcal{M}}$.
  • The set $\mathrm{Per}(\phi^{t})$ of periodic points of the flow $\phi^{t}$ consists of all points $x\in{\mathcal{M}}$ for which there exists $T>0$ with $\phi^{T}(x)=x$.
Definition 2.2.
[…]
$\left\|\phi^{t}_{E}(v)\right\|_{\phi^{t}(p)}\leq Ce^{-c|t|}\left\|v\right\|_{p}$
for all $t\geq 0$ (resp. $t\leq 0$).
For the remainder of Section 2.1, suppose now that ${\mathcal{M}}$ is a smooth manifold and $\phi^{t}$ is a smooth flow with generating vector field $\mathcal{X}:{\mathcal{M}}\to T{\mathcal{M}}$. It is nowhere vanishing since we assume that $\phi^{t}$ has no fixed points.
Definition 2.3.
A compact $\phi^{t}$-invariant set $\mathcal{K}\subset{\mathcal{M}}$ is called hyperbolic for the flow $\phi^{t}$ if the restriction of the tangent bundle $T{\mathcal{M}}$ to $\mathcal{K}$ admits a Whitney sum decomposition
$T{\mathcal{M}}|_{\mathcal{K}}=E^{0}\oplus E^{\mathrm{s}}\oplus E^{\mathrm{u}}$ (2.1)
where $E^{0}_{p}=\mathbb{R}\mathcal{X}(p)$ for all $p\in\mathcal{K}$ and $E^{\mathrm{s}}$, $E^{\mathrm{u}}$ are $d\phi^{t}$-invariant continuous subbundles such that $d\phi^{t}$ is uniformly contracting (resp. expanding) on $E^{\mathrm{s}}$ (resp. $E^{\mathrm{u}}$).
Definition 2.4 (c.f. [Sma67, §II.5 (5.1)]).
The flow $\phi^{t}$ is an Axiom A flow if the non-wandering set ${\mathcal{NW}}(\phi^{t})$ is compact and hyperbolic and coincides with the closure in ${\mathcal{M}}$ of the set of periodic points $\mathrm{Per}(\phi^{t})$.
Note that the manifold $\mathcal{M}$ itself is not assumed to be compact. A hyperbolic set $\mathcal{K}\subset{\mathcal{M}}$ for the flow $\phi^{t}$ is basic if it is locally maximal for $\phi^{t}$ (i.e., there is a neighborhood $\mathcal{U}\subset{\mathcal{M}}$ of $\mathcal{K}$ such that $\mathcal{K}=\bigcap_{t\in\mathbb{R}}\phi^{t}(\mathcal{U})$), the flow $\phi^{t}|_{\mathcal{K}}$ is topologically transitive (i.e., $\mathcal{K}$ contains a dense $\phi^{t}$-orbit), and $\mathcal{K}$ is the closure in ${\mathcal{M}}$ of the set of periodic points of $\phi^{t}|_{\mathcal{K}}$.
2.2. Hyperbolic groups
Here we repeat only the required notation from [DMS24, Section 2.2] and otherwise we formulate only additional preliminary results that are not contained in that reference. For an overview and comprehensive background information, we refer the reader to [Gro87, Gd90, KB02]. Let $\Gamma$ be a finitely generated group, and fix a a choice of generators to define the word length function $|\cdot|:\Gamma\to[0,\infty)$. The stable length of $\gamma\in\Gamma$ is defined by $|\gamma|_{\infty}:=\lim_{n\to\infty}\frac{|\gamma^{n}|}{n}$. Suppose that $\Gamma$ is hyperbolic (synonyms: word-hyperbolic, Gromov-hyperbolic) and let $\partial_{\infty}\Gamma$ be the Gromov boundary of $\Gamma$. Every infinite order element $\gamma\in\Gamma$ has two distinct fixed points $\gamma^{+}:=\lim_{n\to+\infty}\gamma^{n}\in\partial_{\infty}\Gamma$ and $\gamma^{-}:=\lim_{n\to-\infty}\gamma^{n}\in\partial_{\infty}\Gamma$.
The first part of the following lemma generalizes [DMS24, Lemma 2.10] to sequences that do not consist exclusively of infinite order elements.
Lemma 2.5.
Consider a sequence $(\gamma_{k})$ in $\Gamma$ such that the limits $\gamma_{+}=\lim_{k\to\infty}\gamma_{k}\in\partial_{\infty}\Gamma$ and $\gamma_{-}=\lim_{k\to\infty}\gamma_{k}^{-1}\in\partial_{\infty}\Gamma$ exist and are distinct. Then the following holds:
  1. (1)
    $\gamma_{k}$ has infinite order when $k$ is large enough. Furthermore, $\gamma_{k}^{+}\to\gamma_{+}$ and $\gamma_{k}^{-}\to\gamma_{-}$.
  2. (2)
    $|\gamma_{k}|_{\infty}\to\infty$.
Proof.
(1) Let $U_{+},U_{-}\subset\partial_{\infty}\Gamma$ be open neighbourhoods of $\gamma_{+},\gamma_{-}$ respectively with disjoint closures, such that $U_{-}\cup U_{+}\neq\partial_{\infty}\Gamma$ (this is only possible if $\partial_{\infty}\Gamma$ is infinite, but in other cases $\Gamma$ is either finite or virtually isomorphic to $\mathbb{Z}$, the result is straightforward in these cases). From the convergence property of the action on, $\partial_{\infty}\Gamma$, we have that
$\gamma_{k}\cdot\big{(}\partial_{\infty}\Gamma\setminus U_{-}\big{)}\subset U_{+}\quad\textrm{ and }\quad\gamma_{k}^{-1}\cdot\big{(}\partial_{\infty}\Gamma\setminus U_{+}\big{)}\subset U_{-}$
when $k$ is large enough. Such elements $\gamma_{k}$ must have infinite order: otherwise, any element $x\notin U_{-}$ would satisfy $\gamma_{k}^{r}\cdot x\in U_{+}$ where $r$ is the order of $\gamma_{k}$, thus $x\in U_{+}$. The fact that
$\gamma_{k}\cdot U_{+}\subset U_{+}\quad\textrm{ and }\quad\gamma_{k}^{-1}\cdot U_{-}\subset U_{-}$
also implies that $\gamma_{k}^{+}\in U_{+}$ and $\gamma_{k}^{-}\in U_{-}$. By choosing $U_{+}$ and $U_{-}$ arbitrarily small we find that $\gamma_{k}^{+}\to\gamma_{+}$ and $\gamma_{k}^{-}\to\gamma_{-}$.
(2) We know from (1) that $\gamma_{k}$ has infinite order for large $k$, and that $\gamma_{k}^{+}\to\gamma_{+}$ and $\gamma_{k}^{-}\to\gamma_{-}$. Let $\eta=d_{\infty}(\gamma_{+},\gamma_{-})/2$ where $d_{\infty}$ is a distance on $\partial_{\infty}\Gamma$ inducing its cone topology. Then $\gamma_{k}$ is $\eta$-hyperbolic for $k$ large enough (i.e. $d_{\infty}(\gamma_{k}^{+},\gamma_{k}^{-})\geq\eta$), so by [GMT23, Proposition 2.39], there is $C_{1}>0$ such that $|\gamma_{k}|\leq\ell(\gamma_{k})+C_{1}$ where $\ell(\gamma)$ denotes the translation length of $\gamma$. But according to [CDP90, Ch. 10, Prop. 6.4], there is a constant $C_{2}>0$ such that $\ell(\gamma)\leq|\gamma|_{\infty}+C_{2}$ for all $\gamma\in\Gamma$, hence $|\gamma_{k}|_{\infty}\geq|\gamma_{k}|-C_{1}-C_{2}\to+\infty$. ∎
2.3. Semisimple Lie groups and reductive Lie algebras
Let $G$ be a finite dimensional real Lie group with Lie algebra $\mathfrak{g}.$ Then $G$ admits a canonical real analytic structure such that multiplication, inversion, and the exponential map $\exp:\mathfrak{g}\rightarrow G$ are real analytic maps. The Killing form $\mathrm{B}_{\mathfrak{g}}:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathbb{R}$ is defined by the trace pairing
$\mathrm{B}_{\mathfrak{g}}(X,Y)=\mathrm{tr}(\mathrm{ad}_{X}\circ\mathrm{ad}_{Y}).$
The Lie algebra $\mathfrak{g}$ is semisimple if and only if $\mathrm{B}_{\mathfrak{g}}$ is non-degenerate. In this article, we use the definition that the real Lie group $G$ is semisimple if $\mathfrak{g}$ is semisimple and $G$ has finitely many connected components and finite center.
Recall that the real Lie algebra $\mathfrak{g}$ is reductive if it decomposes as a direct sum
$\mathfrak{g}=\mathfrak{z}(\mathfrak{g})\oplus\mathfrak{g}_{\mathrm{ss}}$
where $\mathfrak{g}_{\mathrm{ss}}$ is semisimple. In this article, the only Lie groups with non-semisimple reductive Lie algebra that appear are Levi subgroups $L<G$ where $G$ is semisimple. While there are various definitions of a real reductive Lie group in the literature (see e.g. [Kna02, VII.2]), for the aforementioned reason we will not need a general theory of reductive Lie groups.
3. Parabolic and Levi subgroups, flag manifolds, transverse flag spaces
Let $G$ be a non-compact real semisimple Lie group. In this paper we will use two common approaches to characterize parabolic and Levi subgroups of $G$: First, via hyperbolic elements and stabilizers (Section 3.1). Second, via restricted roots and Weyl chambers (Section 3.4). In that section we also make the connection between the two approaches.
3.1. Parabolic and Levi subgroups as stabilizers
The Killing form $\mathrm{B}_{\mathfrak{g}}$ on $\mathfrak{g}$ is a non-degenerate indefinite symmetric bilinear form giving a canonical duality $\mathfrak{g}\simeq\mathfrak{g}^{\ast}$ intertwining the adjoint and coadjoint $G$-actions. Via the latter actions, $G$ acts on the Grassmannians $\mathcal{G}_{d}(\mathfrak{g})$, $\mathcal{G}_{d}(\mathfrak{g}^{\ast})$ for all $0\leq d\leq\dim{\bf\mathfrak{g}}$. An element $X\in\mathfrak{g}$ is called hyperbolic if $\operatorname{ad}_{X}:\mathfrak{g}\to\mathfrak{g}$ is diagonalizable over $\mathbb{R}$. We write $\mathfrak{g}_{\mathrm{hyp}}\subset\mathfrak{g}$ for the subset of hyperbolic elements. For $X\in\mathfrak{g}_{\mathrm{hyp}}$, we consider
$\begin{split}\mathfrak{l}_{X}&:=\ker(\operatorname{ad}_{X}),\\ \mathfrak{p}_{X}&:=\bigoplus_{\lambda\geq 0}\ker(\operatorname{ad}_{X}-\lambda\mathrm{id}\,),\\ L_{X}&:=\mathrm{Stab}_{G}(X)=\left\{g\in G\,\middle|\,\mathrm{Ad}(g)X=X\right\},\\ P_{X}&:=\mathrm{Stab}_{G}({\bf\mathfrak{p}}_{X})=\left\{g\in G\,\middle|\,\mathrm{Ad}(g)\mathfrak{p}_{X}=\mathfrak{p}_{X}\right\}.\end{split}$ (3.1)
The following is well-known (see e.g. [Vog97, Ch. 7],[Kna02, VII.7], [War72, Ch. 1.2]):
Fact 3.1.
  1. (1)
    $\mathfrak{l}_{X},\mathfrak{p}_{X}\subset\mathfrak{g}$ are Lie subalgebras.
  2. (2)
    $L_{X},P_{X}\subset G$ are closed subgroups whose Lie algebras are $\mathfrak{l}_{X}$ and $\mathfrak{p}_{X}$, respectively, and $L_{X}$ is reductive.
  3. (3)
    $L_{X}=P_{X}\cap P_{-X}$.
Definition 3.2.
A Levi subgroup of $G$ is a subgroup $L<G$ for which there is $X\in\mathfrak{g}_{\mathrm{hyp}}$ with $L=L_{X}$.
A parabolic subgroup of $G$ is a subgroup $P<G$ for which there is $X\in\mathfrak{g}_{\mathrm{hyp}}$ with $P=P_{X}$.
A parabolic subalgebra of ${\bf\mathfrak{g}}$ is a subalgebra ${\bf\mathfrak{p}}\subset{\bf\mathfrak{g}}$ for which there is $X\in\mathfrak{g}_{\mathrm{hyp}}$ with ${\bf\mathfrak{p}}={\bf\mathfrak{p}}_{X}$.
Example 3.3.
In the case of $\mathfrak{g}=\mathfrak{sl}(d,\mathrm{k})$, $\mathrm{k}=\mathbb{R}$ or $\mathbb{C}$, an element $X\in\mathfrak{sl}(d,\mathbb{R})$ is hyperbolic if and only if it is diagonalizable with real eigenvalues. For some integer $r\in\mathbb{N}$, consider vectors $\mathbf{d}=(d_{1},...,d_{r})\in\mathbb{N}^{r}$ and $\mathbf{x}=(x_{1},\dots,x_{r})\in\mathbb{R}^{r}$ such that
$d_{1}+\cdots+d_{r}=d,\qquad d_{1}x_{1}+\cdots+d_{r}x_{r}=0,\qquad\textrm{and }x_{1}>\cdots>x_{r},$
and the diagonal matrix
$X=\begin{pmatrix}x_{1}\mathbf{1}_{d_{1}}&&\\ &\ddots&\\ &&x_{r}\mathbf{1}_{d_{r}}\end{pmatrix}\in\mathfrak{sl}(d,\mathrm{k})\,.$
Then $L_{X}$ consists of block-wise diagonal matrices, and $P_{X}$ (resp. $P_{-X}$) consists of upper (resp. lower) block-wise triangular matrices:
$L_{X}=\left\{\begin{pmatrix}\ast&&0\\ &\ddots&\\ 0&&\ast\end{pmatrix}\right\}\,,\quad P_{X}=\left\{\begin{pmatrix}\ast&&\ast\\ &\ddots&\\ 0&&\ast\end{pmatrix}\right\}\,,\quad P_{-X}=\left\{\begin{pmatrix}\ast&&0\\ &\ddots&\\ \ast&&\ast\end{pmatrix}\right\}\,.$
3.1.1. The split part of the center of a Levi subgroup
Consider a Levi subgroup $L<G$, and denote by $\mathfrak{l}$ its Lie algebra. By definition there is $X\in{\bf\mathfrak{g}}_{\mathrm{hyp}}$ such that $L=L_{X}$, but this element is far from unique. Two such elements must however commute, and a good understanding of the center of $L$ will be important later. Its identity component $Z_{\circ}(L)$ is a connected abelian Lie group, so it splits as a product $Z_{\circ}(L)=R\times T$ where $R$ is isomorphic to some $\mathbb{R}^{n}$ and $T$ is a compact torus. While $R$ is not unique, the maximal compact torus $T$ is, as it can be defined as the set of elements $g\in Z_{\circ}(L)$ such that the closure of $\left\{g^{n}\,\middle|\,n\in\mathbb{Z}\right\}$ is compact. The Killing form of ${\bf\mathfrak{g}}$ and the Lie algebra $\mathfrak{t}$ of the maximal compact torus $T<Z_{\circ}(L)$ allow us to define a canonical $\mathbb{R}$-split torus $Z_{\mathrm{split}}(L)$ in the center of $L$.
Definition 3.4.
Let $L<G$ be a Levi subgroup. The split part $Z_{\mathrm{split}}(L)$ of its center is the group $Z_{\mathrm{split}}(L)=\exp\left(\mathfrak{z}_{\mathrm{split}}(\mathfrak{l})\right)$ where $\mathfrak{z}_{\mathrm{split}}(\mathfrak{l})\subset\mathfrak{z}(\mathfrak{l})$ is the orthogonal (with respect to the restriction of the Killing form $\mathrm{B}_{\bf\mathfrak{g}}$ of ${\bf\mathfrak{g}}$) of the Lie algebra $\mathfrak{t}\subset\mathfrak{z}(\mathfrak{l})$ of the unique maximal compact torus $T<Z_{\circ}(L)$.
Note that for any $X\in{\bf\mathfrak{g}}_{\mathrm{hyp}}$, we have that $X\in\mathfrak{z}_{\mathrm{split}}(\mathfrak{l}_{X})$.
Example 3.5.
In the setting of Example 3.3, we find
$Z_{\mathrm{split}}(L_{X})=\left\{\begin{pmatrix}\lambda_{1}\mathbf{1}_{d_{1}}&&0\\ &\ddots&\\ 0&&\lambda_{r}\mathbf{1}_{d_{r}}\end{pmatrix}\,\middle|\,\lambda_{1},\dots,\lambda_{r}\in\mathbb{R}_{>0}\,,~{}\lambda_{1}^{d_{1}}\cdots\lambda_{r}^{d_{r}}=1\right\}\,.$
3.1.2. The structure of parabolic subgroups
For $X\in\mathfrak{g}_{\mathrm{hyp}}$, consider the the Lie subalgebra
$\mathfrak{n}_{X}:=\bigoplus_{\lambda>0}\ker(\operatorname{ad}_{X}-\lambda\mathrm{id})\subset{\bf\mathfrak{p}}_{X}\,,$ (3.2)
so that we have further decompositions of Lie algebras
$\begin{split}\mathfrak{p}_{X}&=\mathfrak{l}_{X}\oplus\mathfrak{n}_{X},\\ {\bf\mathfrak{g}}&=\mathfrak{n}_{-X}\oplus\mathfrak{l}_{X}\oplus\mathfrak{n}_{X}\,.\end{split}$ (3.3)
Then the subgroup
$N_{X}:=\exp(\mathfrak{n}_{X})=\left\{g\in G\,\middle|\,\lim_{t\to+\infty}\exp(-tX)g\exp(tX)=e\right\}$ (3.4)
is a simply connected closed Ad-unipotent subgroup of $G$ normalizing $L_{X}$ and satisfying $N_{X}\cap L_{X}=\{e\}$ [Vog97, Prop. 7.1]. While the Levi subgroup $L_{X}<P_{X}$ is not uniquely defined as a maximal reductive subgroup of $P_{X},$ the unipotent radical is the unique maximal connected Ad-unipotent normal subgroup $N_{X}\lhd P_{X},$ and there is a short exact sequence
$1\rightarrow N_{X}\rightarrow P_{X}\rightarrow P_{X}/N_{X}\rightarrow 1$
The above short exact sequence is split by $L_{X}\simeq P_{X}/N_{X}$ acting on $N_{X}$ via conjugation and presenting
$P_{X}=N_{X}\rtimes L_{X}=\mathrm{Stab}_{G}(\mathfrak{n}_{X})=\left\{g\in G\,\middle|\,\mathrm{Ad}(g)\mathfrak{n}_{X}=\mathfrak{n}_{X}\right\},$ (3.5)
see [Vog97, Prop. 7.1], [Kna02, VII.7, Prop. 7.82 & 7.83]. Note that we have
$L_{-X}=L_{X},$
while
$N_{-X}\cap N_{X}=\{e\}.$
Given any parabolic subgroup $P$ with Lie algebra ${\bf\mathfrak{p}}$, by definition there is $X\in\mathfrak{g}_{\mathrm{hyp}}$ such that $P=P_{X}$ and ${\bf\mathfrak{p}}={\bf\mathfrak{p}}_{X}$, but as for Levi subgroups this $X$ is far from unique. While the Levi subgroup $L_{X}$ and its Lie algebra $\mathfrak{l}_{X}$ will be different for different choices of $X$, the subgroup $N_{X}$ and (most importantly for us) the subalgebra $\mathfrak{n}_{X}\subset{\bf\mathfrak{p}}_{X}$ do not depend on the choice of $X$, as $\mathfrak{n}_{X}$ is the nilpotent radical of ${\bf\mathfrak{p}}_{X}$.
So denoting by $\mathfrak{n}_{\bf\mathfrak{p}}$ the nilpotent radical of ${\bf\mathfrak{p}}$, we arrive at the equalities
$P=\mathrm{Stab}_{G}({\bf\mathfrak{p}})=\mathrm{Stab}_{G}({\bf\mathfrak{n}}_{\bf\mathfrak{p}})$
which will be the two main characterizations of parabolic subgroups used throughout this paper. Regarding a parabolic Lie algebra ${\bf\mathfrak{p}}$, we will use the tautological property that ${\bf\mathfrak{p}}$ is the Lie algebra of its own stabilizer.
Example 3.6.
In the setting of Example 3.3, we find that $N_{X}$ consists of block-wise upper triangular matrices with identity diagonal blocks:
$N_{X}=\left\{\begin{pmatrix}\mathbf{1}_{d_{1}}&&\ast\\ &\ddots&\\ 0&&\mathbf{1}_{d_{r}}\end{pmatrix}\right\}~{}.$
3.2. Flag manifolds and transverse flag spaces
We now turn to the homogeneous spaces associated to the groups defined in the previous subsection.
Definition 3.7.
A flag manifold is a $G$-homogeneous space $\mathcal{F}$ whose point stabilizers are parabolic subgroups.
The geometric approach to flag manifolds usually consists in interpreting them as orbits in the visual boundary of the Riemannian symmetric space of $G$ (see e.g. [Ebe96, 2.17]), but we will use a different description by embedding them in Grassmannian manifolds of the Lie algebra ${\bf\mathfrak{g}}$. For $X\in{\bf\mathfrak{g}}_{\mathrm{hyp}}$, we consider
$\mathcal{F}_{X}:=G\cdot\mathfrak{p}_{X}\subset\mathcal{G}_{\dim\mathfrak{p}_{X}}(\mathfrak{g}).$
By the definition of $P_{X}$, there is an isomorphism $\mathcal{F}_{X}\simeq G/P_{X}$ as analytic $G$-manifolds. Since flag manifolds are compact [Kna02, VII, Prop. 7.83 (f)], $\mathcal{F}_{X}\subset\mathcal{G}_{\dim\mathfrak{p}_{X}}(\mathfrak{g})$ is closed.
Definition 3.8.
A standard flag manifold is a subset $\mathcal{F}\subset\mathcal{G}_{d}(\mathfrak{g})$ (for some $0\leq d\leq\dim\mathfrak{g}$) for which there is $X\in\mathfrak{g}_{\mathrm{hyp}}$ with $\mathcal{F}=\mathcal{F}_{X}$.
Note that any flag manifold $\mathcal{F}$ is $G$-equivariantly diffeomorphic to the standard flag manifold $\left\{\mathfrak{stab}(\mathbf{x})\,\middle|\,\mathbf{x}\in{\mathcal{F}}\right\}$ where $\mathfrak{stab}(\mathbf{x})$ denotes the Lie algebra of the stabilizer of $\mathbf{x}$.
While flag manifolds will play a central role in our arguments, the most important objects will be the open $G$-orbits formed by the transverse pairs in products of opposite flag manifolds.
Definition 3.9.
Consider two flag manifolds $\mathcal{F}^{+}$ and $\mathcal{F}^{-}$. A pair $(\mathbf{x}^{+},\mathbf{x}^{-})\in\mathcal{F}^{+}\times\mathcal{F}^{-}$ is transverse if there is $X\in\mathfrak{g}_{\mathrm{hyp}}$ such that $\mathrm{Stab}_{G}(\mathbf{x}^{+})=P_{X}$ and $\mathrm{Stab}_{G}(\mathbf{x}^{-})=P_{-X}$. We write $\mathbf{x}^{+}\pitchfork\mathbf{x}^{-}$ if $(\mathbf{x}^{+},\mathbf{x}^{-})$ is transverse and denote the set of all transverse pairs in $\mathcal{F}^{+}\times\mathcal{F}^{-}$ by ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ (or ${{\mathcal{F}}^{\pitchfork}}$ for short, when the factors ${\mathcal{F}}^{+}$, ${\mathcal{F}}^{-}$ are clear from the context).
The flag manifolds $\mathcal{F}^{+},\mathcal{F}^{-}$ are opposite if ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}\neq\emptyset$. In this case, we call ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ a transverse flag space.
In view of (3.3) one then has for every pair $(\mathbf{x}^{+},\mathbf{x}^{-})\in{\mathcal{F}}^{+}\times{\mathcal{F}}^{-}$, denoting by ${\bf\mathfrak{p}}^{+},{\bf\mathfrak{p}}^{-}$ the Lie algebras of their stabilizers and $\mathfrak{n}_{{\bf\mathfrak{p}}^{+}},\mathfrak{n}_{{\bf\mathfrak{p}}^{-}}$ their nilpotent radicals,
$\mathbf{x}^{+}\pitchfork\mathbf{x}^{-}\iff\mathfrak{p}^{+}\oplus\mathfrak{n}_{\mathfrak{p}^{-}}=\mathfrak{n}_{\mathfrak{p}^{+}}\oplus\mathfrak{p}^{-}={\bf\mathfrak{g}},$ (3.6)
which justifies the terminology “transverse” and shows that the set ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$, which by definition is a single $G$-orbit, is open and dense in ${\mathcal{F}}^{+}\times{\mathcal{F}}^{-}$.
Every transverse flag space ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ comes with a pair of projections
${{\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}}$${{\mathcal{F}}^{+}}$${{\mathcal{F}}^{-}}$$\scriptstyle{p_{{\mathcal{F}}^{+}}}$$\scriptstyle{p_{{\mathcal{F}}^{-}}}$ (3.7)
given by the restriction to ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ of the canonical projections ${\mathcal{F}}^{+}\times{\mathcal{F}}^{-}\to{\mathcal{F}}^{\pm}$.
Example 3.10.
Consider a vector $\mathbf{d}=(d_{1},...,d_{r})\in\mathbb{N}^{r}$ with $d_{1}+\cdots+d_{r}=d$. The space $\mathcal{F}_{\mathbf{d}}$ of flags of type $\mathbf{d}$ is defined as
$\mathcal{F}_{\mathbf{d}}(\mathrm{k}^{d})=\left\{V_{\bullet}\in\prod_{i=0}^{r}{\mathcal{G}}_{d_{1}+\cdots+d_{i}}(\mathrm{k}^{d})\,\middle|\,\{0\}=V_{0}\subset V_{1}\subset\cdots\subset V_{r-1}\subset V_{r}=\mathrm{k}^{d}\right\}\,.$
Just as in Example 3.3, consider a real diagonal matrix $X=\mathrm{Diag}\left(x_{1}\mathbf{1}_{d_{1}},\dots,x_{r}\mathbf{1}_{d_{r}}\right)\in\mathfrak{sl}(d,\mathrm{k})$, with $x_{1}>\cdots>x_{r}$. The map
$\left\{\begin{array}[]{ccc}{\mathcal{F}}_{\mathbf{d}}(\mathrm{k}^{d})&\to&{\mathcal{F}}_{X}\\ V_{\bullet}&\mapsto&\left\{Y\in\mathfrak{sl}(d,\mathrm{k})\,\middle|\,Y\cdot V_{i}\subset V_{i}~{}\forall i\right\}\end{array}\right.$
is an $\operatorname{SL}(d,\mathrm{k})$-equivariant diffeomorphism. The same formula gives an $\operatorname{SL}(d,\mathrm{k})$-equivariant diffeomorphism between ${\mathcal{F}}_{\iota(\mathbf{d})}(\mathrm{k}^{d})$ and ${\mathcal{F}}_{-X}$, where $\iota(\mathbf{d})=(d_{r},\dots,d_{1})$, and the transverse flag space
${\mathcal{F}}_{\mathbf{d}}(\mathrm{k}^{d})\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}_{\iota(\mathbf{d})}(\mathrm{k}^{d})=\left\{\big{(}V_{\bullet},W_{\bullet}\big{)}\,\middle|\,V_{i}\oplus W_{r-i}=\mathrm{k}^{d}~{}\forall i\right\}$
can be identified with the space of gradings of type $\mathbf{d}$
${\mathcal{F}}_{\mathbf{d}}^{\pitchfork}(\mathrm{k}^{d}):=\left\{E_{\bullet}\in\prod_{i=1}^{r}\mathcal{G}_{d_{i}}(\mathrm{k}^{d})\,\middle|\,E_{1}\oplus\cdots\oplus E_{r}=\mathrm{k}^{d}\right\}$
through the map
$\left\{\begin{array}[]{ccc}{\mathcal{F}}_{\mathbf{d}}(\mathrm{k}^{d})\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}_{\iota(\mathbf{d})}(\mathrm{k}^{d})&\to&{\mathcal{F}}_{\mathbf{d}}^{\pitchfork}(\mathrm{k}^{d})\\ \big{(}V_{\bullet},W_{\bullet}\big{)}&\mapsto&(V_{i}\cap W_{r+1-i})_{1\leq i\leq r}\end{array}\right.$
whose inverse maps $E_{\bullet}\in{\mathcal{F}}_{\mathbf{d}}^{\pitchfork}(\mathrm{k}^{d})$ to the pair $(V_{\bullet},W_{\bullet})$ where $V_{i}=E_{1}\oplus\cdots\oplus E_{i}$ and $W_{i}=E_{r}\oplus\cdots\oplus E_{r+1-i}$.
3.3. Pseudo-Riemannian structure of transverse flag spaces
The tangent space of a transverse flag space ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ at a point $(\mathbf{x}^{+},\mathbf{x}^{-})$ has a nice description: denoting by ${\bf\mathfrak{p}}^{+}$ and ${\bf\mathfrak{p}}^{-}$ the Lie algebras of their respective stabilizers and $\mathfrak{l}={\bf\mathfrak{p}}^{+}\cap{\bf\mathfrak{p}}^{-}$, the projection of $\mathfrak{n}_{\mathfrak{p}^{+}}\oplus\mathfrak{n}_{\mathfrak{p}^{-}}\subset{\bf\mathfrak{g}}$ onto ${\bf\mathfrak{g}}/\mathfrak{l}$ is an isomorphism, and the latter can be identified with $T_{(\mathbf{x}^{+},\mathbf{x}^{-})}{{\mathcal{F}}^{\pitchfork}}$ via the derivative of the $G$-action.
This description allows for the definition of a $G$-invariant pseudo-Riemannian metric induced by the Killing form $\mathrm{B}_{\bf\mathfrak{g}}$ because the subspace $\mathfrak{n}_{\mathfrak{p}^{+}}\oplus\mathfrak{n}_{\mathfrak{p}^{-}}\subset{\bf\mathfrak{g}}$ is non-degenerate. More details concerning this pseudo-Riemannian metric and other geometric structures on transverse flag spaces will be discussed in Section 9.
3.4. Standard parabolic subgroups and flag manifolds
Let $\theta:{\bf\mathfrak{g}}\to{\bf\mathfrak{g}}$ be a Cartan involution defining a Cartan decomposition ${\bf\mathfrak{g}}={\bf\mathfrak{k}}\oplus{\bf\mathfrak{k}}^{\perp}$ into its $+1$ and $-1$ eigenspaces, which are orthogonal with respect to the Killing form $\mathrm{B}_{\bf\mathfrak{g}}$, and let $K<G$ be the maximal compact subgroup with Lie algebra ${\bf\mathfrak{k}}$. It is the fixed point set of the global Cartan involution, which by abuse of notation we again denote by $\theta:G\to G$. The Cartan decomposition ${\bf\mathfrak{g}}={\bf\mathfrak{k}}\oplus{\bf\mathfrak{k}}^{\perp}$ is $\mathrm{Ad}(K)$-invariant.
Choose a maximal abelian subalgebra ${\bf\mathfrak{a}}$ of ${\bf\mathfrak{k}}^{\perp}$. Then the Killing form $\mathrm{B}_{\bf\mathfrak{g}}$ restricts to an inner product on ${\bf\mathfrak{a}}$ and we get the restricted root decomposition
$\mathfrak{g}=\mathfrak{m}\oplus\mathfrak{a}\oplus\bigoplus_{\alpha\in\Sigma}\mathfrak{g}_{\alpha},$
where $\mathfrak{m}=\mathfrak{z}_{\bf\mathfrak{k}}({\bf\mathfrak{a}})$. Now, we choose a simple system
$\Delta\subset\Sigma,$
which defines a closed Weyl chamber $\mathfrak{a}^{+}=\{Y\in{\bf\mathfrak{a}}^{\ast}\,|\,\alpha(Y)\geq 0\;\forall\;\alpha\in\Delta\}$. The latter is a fundamental domain for the action of the Weyl group $W=N_{K}({\bf\mathfrak{a}})/Z_{K}({\bf\mathfrak{a}})$. We also obtain the Iwasawa decomposition
$G=KAN,$
where $A=\exp({\bf\mathfrak{a}})$, $N=\exp({\bf\mathfrak{n}})$ with ${\bf\mathfrak{n}}:=\bigoplus_{\alpha\in\Sigma^{+}}\mathfrak{g}_{\alpha}$ and $\Sigma^{+}\subset\Sigma$ consists of all roots that are positive linear combinations of roots in $\Delta$. Let $w_{0}\in W$ be the longest Weyl group element and $\iota_{\mathrm{o}}:{\bf\mathfrak{a}}\to{\bf\mathfrak{a}}$, $X\mapsto-w_{0}(X)$, the opposition involution. The latter has the key property that it maps $\mathfrak{a}^{+}$ into itself, and it also acts on $\Delta$. For more details on all this, we refer the reader to [Kna02, Ch. VII], [War72, Ch. 2] and [Wal88, Ch. 1 & 2].
Fact 3.11 ([Vog97, Thm. 6.2]).
The inclusions ${\bf\mathfrak{a}}\hookrightarrow{\bf\mathfrak{k}}^{\perp}\hookrightarrow{\bf\mathfrak{g}}$ induce bijections between the $W$-orbits in ${\bf\mathfrak{a}}$, the $\mathrm{Ad}(K)$-orbits in ${\bf\mathfrak{k}}^{\perp}$, and the hyperbolic $\mathrm{Ad}(G)$-orbits in ${\bf\mathfrak{g}}$. In particular, every hyperbolic adjoint $G$-orbit intersects ${\bf\mathfrak{a}}^{+}$ in exactly one element, and one has $\mathrm{Ad}(G){\bf\mathfrak{a}}^{+}={\bf\mathfrak{g}}_{\mathrm{hyp}}$, $\mathrm{Ad}(K){\bf\mathfrak{a}}^{+}={\bf\mathfrak{k}}^{\perp}$.
Fact 3.11 allows us to use the simple system $\Delta$ for classifying parabolic and Levi subgroups as well as the associated flag manifolds and transverse flag spaces. The so-called standard parabolic subgroups associated to $\Delta$ can be described in many equivalent ways (see e.g. [Kna02, VII.7], [Wal88, Ch. 2]). For us, the following rather quick description suffices (c.f. [Kna02, VII, Prop. 7.82, 7.83], [GW12, Sec. 3.2]): Let ${\Theta}\subset\Delta$ be a subset and define
${\bf\mathfrak{a}}_{\Theta}:=\bigcap_{\alpha\in\Delta\setminus{\Theta}}\ker\alpha\subset{\bf\mathfrak{a}},\qquad P_{\Theta}:=Z_{K}({\bf\mathfrak{a}}_{\Theta})AN.$ (3.8)
$P_{\Theta}$ is the standard parabolic subgroup associated to $\Theta$. Its opposite is $\overline{P}_{\Theta}:=\theta(P_{\Theta})$, which is conjugate to $P_{\iota_{o}(\Theta)}$. The standard Levi subgroup $L_{\Theta}:=P_{\Theta}\cap\overline{P}_{\Theta}$ and its split component are given by $L_{\Theta}=Z_{G}({\bf\mathfrak{a}}_{\Theta})$,
$Z_{\mathrm{split}}(L_{\Theta})=\exp({\bf\mathfrak{a}}_{\Theta})\subset A,$ (3.9)
see [Kna02, VII, Prop. 7.82]. Every parabolic subgroup of $G$ containing the minimal parabolic subgroup $P_{\Delta}=Z_{K}({\bf\mathfrak{a}})AN$ is of the form $P_{\Theta}$ for a unique subset ${\Theta}\subset\Delta$ [Kna02, VII, Prop. 7.76].
Let us now make the connection to Section 3.1: Given $X\in\mathfrak{a}^{+}$, consider the set
$\Theta_{X}:=\left\{\alpha\in\Delta\,\middle|\,\alpha(X)\neq 0\right\}=\left\{\alpha\in\Delta\,\middle|\,\alpha(X)>0\right\}\subset\Delta.$ (3.10)
Then one finds that $P_{X}=P_{\Theta_{X}}$ and $L_{X}=L_{\Theta_{X}}$. As for the Lie algebras, one gets
$\begin{split}\mathfrak{p}_{X}&=\mathfrak{m}\oplus\mathfrak{a}\oplus\bigoplus_{\begin{subarray}{c}\alpha\in\Sigma\\ \alpha(X)\geq 0\end{subarray}}\mathfrak{g}_{\alpha},\\ \mathfrak{l}_{X}&=\mathfrak{m}\oplus\mathfrak{a}\oplus\bigoplus_{\begin{subarray}{c}\alpha\in\Sigma\\ \alpha(X)=0\end{subarray}}\mathfrak{g}_{\alpha},\\ \mathfrak{n}_{X}&=\bigoplus_{\begin{subarray}{c}\alpha\in\Sigma\\ \alpha(X)>0\end{subarray}}\mathfrak{g}_{\alpha}.\end{split}$ (3.11)
It follows that $P_{X}$ depends only indirectly on $X$ via the set $\Theta_{X}$. More precisely, it follows from the one-to-one correspondence betweeen standard parabolic subgroups and subsets of $\Delta$ that for all $X,X^{\prime}\in\mathfrak{a}^{+}$ we have
$\Theta_{X}\subset\Theta_{X^{\prime}}\;\iff\;P_{X^{\prime}}<P_{X}.$ (3.12)
Definition 3.12.
For $\Theta\subset\Delta$, the standard flag manifold associated to $\Theta$ is $\mathcal{F}_{\Theta}:=G\cdot\mathfrak{p}_{\Theta}$.
Note that ${\mathcal{F}}_{\Theta}$ is in fact a standard flag manifold according to Definition 3.8, since ${\mathcal{F}}_{\Theta}={\mathcal{F}}_{X}$ for any $X\in{\bf\mathfrak{a}}^{+}$ such that $\Theta_{X}=\Theta$. From (3.11) and Fact 3.11, we get the following:
Proposition 3.13.
Every parabolic subgroup is conjugate to a standard parabolic subgroup $P_{\Theta}$ for a unique subset $\Theta\subset\Delta$, and every flag manifold is $G$-equivariantly diffeomorphic to a unique standard flag manifold.
Choosing a different root space decomposition provides standard parabolic subgroups that are conjugate to each other. However, the standard flag manifolds actually equal, and not just isomorphic to each other.
Proposition 3.14.
For every standard flag manifold $\mathcal{F}$, there is a unique subset $\Theta\subset\Delta$ such that $\mathcal{F}=\mathcal{F}_{\Theta}$.
Proof.
By Fact 3.11 every $X\in\mathfrak{g}_{\mathrm{hyp}}$ is transported by $\mathrm{Ad}(G)$ to a unique element of $X^{\prime}\in\mathfrak{a}^{+}$, and ${\mathcal{F}}_{X}={\mathcal{F}}_{X^{\prime}}={\mathcal{F}}_{\Theta_{X^{\prime}}}$ by (3.11). The uniqueness follows from Proposition 3.13. ∎
Example 3.15.
In the study of $\operatorname{SL}(d,\mathrm{k})$, $\mathrm{k}=\mathbb{R}$ or $\mathbb{C}$, the usual description (see [Kna02, p. 509]) leads to choosing for ${\bf\mathfrak{a}}\subset\mathfrak{sl}(d,\mathrm{k})$ the space of real diagonal traceless matrices, and choosing as a closed Weyl chamber
${\bf\mathfrak{a}}^{+}=\left\{\mathrm{Diag}(x_{1},\ldots,x_{d})\in{\bf\mathfrak{a}}\,\middle|\,x_{1}\geq\cdots\geq x_{d}\right\}\,,$
corresponding to simple restricted roots $\Delta=\{\alpha_{1},\dots,\alpha_{d-1}\}$ where
$\alpha_{j}\big{(}\mathrm{Diag}(x_{1},\ldots,x_{d})\big{)}:=x_{j}-x_{j+1}\,.$
For a matrix $X=\mathrm{Diag}\left(x_{1}\mathbf{1}_{d_{1}},\dots,x_{r}\mathbf{1}_{d_{r}}\right)\in{\bf\mathfrak{a}}^{+}$ as in Example 3.3, we find
$\Theta_{X}=\{\alpha_{d_{1}},\alpha_{d_{1}+d_{2}},\dots,\alpha_{d_{1}+\cdots+d_{r-1}}\}\,.$
The corresponding standard parabolic subgroups are the block-wise upper triangular groups described in Example 3.3, and the opposition involution acts on $\Delta$ as $\iota_{o}(\alpha_{j})=\alpha_{d-j}$.
3.5. The hierarchy of flag manifolds
The correspondence between standard flag manifolds and subsets of simple restricted roots endows the set of standard flag manifolds with a partial ordering, corresponding to inclusion of subsets of $\Delta$. We shall see in this section how to describe this partial ordering without fixing a restricted root decomposition.
Lemma 3.16.
Let ${\mathcal{F}},{\mathcal{F}}^{\prime}$ be two flag manifolds. If there exists a $G$-equivariant map ${\mathcal{F}}\to{\mathcal{F}}^{\prime}$, then it is unique and it is a fiber bundle projection.
Proof.
Consider two $G$-equivariant maps $\varphi,\psi:{\mathcal{F}}\to{\mathcal{F}}^{\prime}$. Fix a restricted root space decomposition, as well as an element $\mathbf{x}_{0}\in{\mathcal{F}}$ whose stabilizer $\mathrm{Stab}_{G}(\mathbf{x}_{0})$ is a standard parabolic subgroup $P_{\Theta}$ as defined in section 3.4. Then the stabilizers of $\varphi(\mathbf{x}_{0})$ and $\psi(\mathbf{x}_{0})$ are both parabolic subgroups of $G$ containing $P_{\Theta}$, so according to [Kna02, VII, Prop. 7.76], they are also standard parabolic subgroups. But they are conjugate to each other, so $\varphi(\mathbf{x}_{0})$ and $\psi(\mathbf{x}_{0})$ have the same stabilizer $P_{\Theta^{\prime}}$ for some $\Theta^{\prime}\subset\Theta$, thus $\varphi(\mathbf{x}_{0})=\psi(\mathbf{x}_{0})$ (because parabolic subgroups are their own normalizers), and the equality $\varphi=\psi$ follows from the $G$-equivariance. In the coset description $\varphi$ corresponds to the canonical projection $G/P_{\Theta}\to G/P_{\Theta^{\prime}}$. ∎
In the case of isomorphic (i.e. $G$-equivariantly diffeomorphic) flag manifolds, Lemma 3.16 has a strong implication: the classification up to isomorphisms of Proposition 3.13 is actually a classification up to unique isomorphism. Practically, this means that choosing between working with abstract flag manifolds or standard flag manifolds is merely a matter of taste. Lemma 3.16 is also a motivation for the following partial ordering of flag manifolds:
Definition 3.17.
Let ${\mathcal{F}},{\mathcal{F}}^{\prime}$ be two flag manifolds. We say that ${\mathcal{F}}$ dominates ${\mathcal{F}}^{\prime}$, and denote it by ${\mathcal{F}}^{\prime}\prec{\mathcal{F}}$, if there is a $G$-equivariant map $\Pi^{\mathcal{F}}_{\mathcal{F}^{\prime}}:{\mathcal{F}}\to{\mathcal{F}}^{\prime}$.
The hierarchy defined by this partial ordering corresponds to inclusion for subsets of simple restricted roots.
Proposition 3.18.
Fix a restricted root space decomposition of ${\bf\mathfrak{g}}$. For any two subsets $\Theta,\Theta^{\prime}\subset\Delta$ we have $\mathcal{F}_{\Theta}\prec\mathcal{F}_{\Theta^{\prime}}$ iff $\Theta\subset\Theta^{\prime}$.
Proof.
It follows from (3.12). ∎
So far there is an important operation on simple restricted roots that we have not described in terms of simple flag manifolds: the opposition involution. It happens to coincide with the notion of opposite flag manifolds.
Proposition 3.19.
Every standard flag manifold has a unique opposite standard flag manifold. Given a restricted root space decomposition and a subset $\Theta\subset\Delta$, the standard flag manifold opposite to $\mathcal{F}_{\Theta}$ is $\mathcal{F}_{\iota_{\mathrm{o}}(\Theta)}$.
Proof.
Let ${\mathcal{F}}$ be a standard flag manifold. Proposition 3.14 tells us that for a fixed choice of restricted root space decomposition with closed Weyl chamber ${\bf\mathfrak{a}}^{+}$ there is $X\in{\bf\mathfrak{a}}^{+}$ such that ${\mathcal{F}}={\mathcal{F}}_{X}={\mathcal{F}}_{\Theta_{X}}$ in the notation of Definition 3.12, and the set ${\Theta_{X}}$ determines ${\mathcal{F}}$ uniquely. Now ${\mathcal{F}}_{-X}$ is opposite to ${\mathcal{F}}$, and by Fact 3.11 the adjoint $G$-orbit of $-X$ intersects $\mathfrak{a}^{+}$ in a unique element $X^{\prime}$, so that ${\mathcal{F}}_{-X}={\mathcal{F}}_{\Theta_{X^{\prime}}}$ is uniquely determined by $\Theta_{X^{\prime}}$. Concretely $X^{\prime}$ is given by $X^{\prime}=\iota_{\mathrm{o}}(X)$ and we get $\Theta_{X^{\prime}}=\iota_{\mathrm{o}}(\Theta_{X})$. In particular, $\Theta_{X^{\prime}}$ is uniquely determined by $\Theta_{X}$ If $Y\in{\bf\mathfrak{a}}^{+}$ is another element such that ${\mathcal{F}}={\mathcal{F}}_{Y}$, then by Proposition 3.14 we have $\Theta_{Y}=\Theta_{X}$ and thus ${\mathcal{F}}_{-Y}={\mathcal{F}}_{\iota_{\mathrm{o}}(\Theta_{X})}={\mathcal{F}}_{-X}$. Thus, ${\mathcal{F}}_{\iota_{\mathrm{o}}(\Theta_{X})}$ is the unique opposite of ${\mathcal{F}}_{\Theta_{X}}$ and the proof is finished. ∎
Given a standard flag manifold ${\mathcal{F}}$, we write $\overline{{\mathcal{F}}}$ for its opposite standard flag manifold. The notions of “opposition” for standard flag manifolds and parabolic subgroups are compatible in the sense that the opposite of ${\mathcal{F}}_{\Theta}=G\cdot{\bf\mathfrak{p}}_{\Theta}$ is ¯$\overline{\mathcal{F}}_{\Theta}={\mathcal{F}}_{\iota_{o}(\Theta)}=G\cdot\overline{\bf\mathfrak{p}}_{\Theta}$, where $\overline{\bf\mathfrak{p}}_{\Theta}=\theta({\bf\mathfrak{p}}_{\Theta})$ is the image under the Cartan involution of the Lie algebra of $P_{\Theta}$.
Corollary 3.20.
Consider two transverse flag spaces ${\mathcal{F}}_{1}^{\pitchfork}={\mathcal{F}}^{+}_{1}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}_{1}^{-}$ and ${\mathcal{F}}_{2}^{\pitchfork}={\mathcal{F}}^{+}_{2}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}_{2}^{-}$. Then ${\mathcal{F}}_{1}^{+}\prec{\mathcal{F}}_{2}^{+}$ if and only if ${\mathcal{F}}_{1}^{-}\prec{\mathcal{F}}_{2}^{-}$. Furthermore, if these dominations occur, then
$\Big{(}\Pi^{{\mathcal{F}}^{+}_{2}}_{{\mathcal{F}}_{1}^{+}}(\mathbf{x}^{+}),\Pi^{{\mathcal{F}}_{2}^{-}}_{{\mathcal{F}}_{1}^{-}}(\mathbf{x}^{-})\Big{)}\in{\mathcal{F}}_{1}^{\pitchfork}\qquad\forall(\mathbf{x}^{+},\mathbf{x}^{-})\in{\mathcal{F}}_{2}^{\pitchfork}\,.$
Proof.
The equivalence follows from Proposition 3.14 and Proposition 3.18. For the second statement, we may fix a restricted root space decomposition and assume that $\mathrm{Stab}_{G}(\mathbf{x}^{+})=P_{\Theta_{2}}$ and $\mathrm{Stab}_{G}(\mathbf{x}^{-})=\overline{P}_{\Theta_{2}}$ for some subset $\Theta_{2}\subset\Delta$. Simplify the notations to
$\pi^{+}:=\Pi^{{\mathcal{F}}^{+}_{2}}_{{\mathcal{F}}_{1}^{+}}~{},\qquad\pi^{-}:=\Pi^{{\mathcal{F}}_{2}^{-}}_{{\mathcal{F}}_{1}^{-}}\,.$
Now $\mathrm{Stab}_{G}(\pi^{+}(\mathbf{x}^{+}))$ is a parabolic subgroup of $G$ containing $P_{\Theta_{2}}$, so by [Kna02, VII, Prop. 7.76] there is a subset $\Theta_{1}\subset\Delta$ such that $\mathrm{Stab}_{G}(\pi^{+}(\mathbf{x}^{+}))=P_{\Theta_{1}}$. Similarly there is a subset $\Theta^{\prime}_{1}\subset\Delta$ such that $\mathrm{Stab}_{G}(\pi^{-}(\mathbf{x}^{-}))=\overline{P}_{\Theta^{\prime}_{1}}$. But the fact that the flag manifolds ${\mathcal{F}}_{1}^{+}={\mathcal{F}}_{\Theta_{1}}$ and ${\mathcal{F}}_{1}^{-}={\mathcal{F}}_{\iota_{o}(\Theta^{\prime}_{1})}$ are opposite implies that $\Theta^{\prime}_{1}=\Theta_{1}$, therefore $(\pi^{+}(\mathbf{x}^{+}),\pi^{-}(\mathbf{x}^{-}))\in{\mathcal{F}}_{1}^{\pitchfork}$. ∎
3.6. Simple flag manifolds
The partial ordering introduced in Definition 3.17 has a minimal element $\{\mathfrak{g}\}\in\mathcal{G}_{\dim\mathfrak{g}}(\mathfrak{g})$ given by a one point space (associated to the empty set $\emptyset\subset\Delta$ in a restricted root space decomposition). We shall call a flag manifold non-trivial if it does not consist of a single point.
Definition 3.21.
A simple flag manifold is a non-trivial standard flag manifold that does not dominate any other non-trivial flag manifold.
Once a restricted root space decomposition has been fixed, Proposition 3.18 implies that the simple flag manifolds are exactly the ${\mathcal{F}}_{\alpha}:={\mathcal{F}}_{\{\alpha\}}$ for $\alpha\in\Delta$, thus providing a bijection
$\begin{split}\Delta&\simeq\{\text{simple flag manifolds}\}\\ \alpha&\mapsto{\mathcal{F}}_{\alpha}.\end{split}$ (3.13)
The opposition involution $\iota_{\mathrm{o}}$ on $\Delta$ can be transported to the set of simple flag manifolds by the bijection (3.13), but thanks to Proposition 3.19 it has an intrinsic meaning: it consists of the map sending a flag manifold to its unique opposite simple flag manifold.
In many occurences, all that will be used from a root space decomposition is the possibility of indexing flag manifolds by subsets of $\Delta$ and the fact that indexing opposite flag manifolds corresponds to the opposition involution. In this case, (3.13) allows us to forget about root space decompositions and just consider $\Delta$ as any finite set indexing simple flag manifolds (e.g. the set of simple flag manifolds itself), and define the opposition involution by the relation ${\mathcal{F}}_{\iota_{o}(\alpha)}=\overline{\mathcal{F}}_{\alpha}$ for $\alpha\in\Delta$. Then any flag manifold ${\mathcal{F}}$ corresponds to a subset $\Theta_{\mathcal{F}}\subset\Delta$ which is the set of simple flag manifolds dominated by ${\mathcal{F}}$:
$\begin{split}\Theta_{{\mathcal{F}}}&:=\left\{\alpha\in\Delta\,\middle|\,{\mathcal{F}}_{\alpha}\prec{\mathcal{F}}\right\}.\end{split}$ (3.14)
Concretely, this means that we can either start our discussion with a subset $\Theta\subset\Delta$ (in a fixed restricted root space decomposition) and consider the associated standard flag manifold ${\mathcal{F}}_{\Theta}$, or start with a flag manifold ${\mathcal{F}}$ and consider the set $\Theta_{\mathcal{F}}$ from (3.14). By using the notations ${\mathcal{F}}$ for the flag manifold and $\Theta\subset\Delta$ for the subset, both point of views lead to the same subsequent discussion.
A good example of a property that can be expressed with either point of view is the definition of a $G$-equivariant fiber bundle projection
$\Pi^{\mathcal{F}}_{\alpha}:{\mathcal{F}}\to{\mathcal{F}}_{\alpha}$ (3.15)
for each $\alpha\in\Theta$. From the point of view of a fixed restricted root decomposition, it comes from the inclusion $P_{\Theta}<P_{\{\alpha\}}$ and the identifications ${\mathcal{F}}\simeq G/P_{\Theta}$ and ${\mathcal{F}}_{\alpha}\simeq G/P_{\{\alpha\}}$. From the intrinsic point of view on simple flag manifolds, the existence of a $G$-equivariant map ${\mathcal{F}}\to{\mathcal{F}}_{\alpha}$ for $\alpha\in\Theta$ is the definition of $\Theta$, and Lemma 3.16 guarantees its uniqueness. It also has a simple meaning: $\Pi_{\alpha}^{\mathcal{F}}(\mathbf{x})$ is the unique element of ${\mathcal{F}}_{\alpha}\subset\mathcal{G}_{d_{\alpha}}({\bf\mathfrak{g}})$ containing the Lie algebra of the stabilizer of $\mathbf{x}\in{\mathcal{F}}$.
In the following sections, a restricted root space decomposition will only be introduced when further algebraic considerations are necessary.
Example 3.22.
Carrying on with the notations of Example 3.15, given a dimension vector $\mathbf{d}=(d_{1},\dots,d_{r})\in\mathbb{N}^{r}$ with $d_{1}+\cdots+d_{r}=d$, the space ${\mathcal{F}}_{\mathbf{d}}(\mathrm{k}^{d})$ of flags of type $\mathbf{d}$ introduced in Example 3.10 is $\operatorname{SL}(d,\mathrm{k})$-equivariantly diffeomorphic to the standard flag manifold corresponding to the subset
$\Theta_{\mathbf{d}}:=\{\alpha_{d_{1}},\alpha_{d_{1}+d_{2}},\dots,\alpha_{d_{1}+\cdots+d_{r-1}}\}\,.$
Simple flag manifolds correspond to Grassmannians ${\mathcal{G}}_{n}(\mathrm{k}^{d})=\mathcal{F}_{\{n,d-n\}}(\mathrm{k}^{d})$, and the equivariant projections $\Pi^{\mathbf{d}}_{j}:{\mathcal{F}}_{\mathbf{d}}(\mathrm{k}^{d})\to{\mathcal{G}}_{d_{1}+\cdots+d_{j}}(\mathrm{k}^{d})$ are defined by
$\Pi^{\mathbf{d}}_{j}\big{(}V_{\bullet}\big{)}=V_{j}\,.$
3.7. Flag manifolds and the Jordan decomposition
The Jordan decomposition tells us that every group element $g\in G$ can be uniquely written as a product $g=g_{e}g_{h}g_{u}$, where $g_{e}$, $g_{h}$, and $g_{u}$ are elliptic, hyperbolic, and unipotent, respectively and commute with each other (see [Kos73, Prop. 2.1]). Here $g_{e}$ being elliptic means that $\mathrm{Ad}(g_{e})$ is diagonalizable over $\mathbb{C}$ with eigenvalues of absolute value $1$, $g_{h}$ being hyperbolic means that $g_{h}=\exp(X_{h})$ with $X_{h}\in{\bf\mathfrak{g}}$ hyperbolic (i.e., $\operatorname{ad}_{X}$ is diagonalizable over $\mathbb{R}$, as in Section 3.1) and $g_{u}$ being unipotent means that $g_{u}=\exp(X_{n})$ with $X_{n}\in{\bf\mathfrak{g}}$ nilpotent.
3.7.1. Jordan projections in $G$
Let us fix a restricted root space decomposition as in Section 3.4, in particular we fix an Iwasawa decomposition $G=KAN$ which, in contrast to the Jordan decomposition, involves making choices. Then a group element being elliptic (resp. hyperbolic, unipotent) means that it is conjugate to an element in $K$ (resp. $A$, $N$).
Definition 3.23.
Let $g\in G$. The Jordan projection [^1] The Jordan projection is sometimes also called Lyapunov projection. of $g$ with respect to the Weyl chamber $\mathfrak{a}^{+}$ is the element $\lambda(g)\in\mathfrak{a}^{+}$ such that $g_{h}$ is conjugate to $\exp(\lambda(g))$.
Fact 3.11 tells us that the Jordan projection is well-defined and has the property
$\lambda(g^{-1})=\iota_{o}(\lambda(g))\qquad\forall\;g\in G.$ (3.16)
3.7.2. Fundamental weights
Fixing a restricted root space decomposition provides a vector $X_{\alpha}\in{\bf\mathfrak{g}}_{\mathrm{hyp}}$ such that ${\mathcal{F}}_{\alpha}={\mathcal{F}}_{X_{\alpha}}$ for each $\alpha\in\Delta$. Indeed, consider the fundamental weight $w_{\alpha}\in\mathfrak{a}^{*}$ defined by
$\mathrm{B}_{\mathfrak{g}}(w_{\alpha},\alpha)=\frac{1}{2}\mathrm{B}_{\mathfrak{g}}(\alpha,\alpha),\qquad\mathrm{B}_{\mathfrak{g}}(w_{\alpha},\alpha^{\prime})=0\quad\forall\;\alpha^{\prime}\in\Delta\setminus\{\alpha\}.$ (3.17)
Let $X_{\alpha}\in{\bf\mathfrak{a}}$ be the $\mathrm{B}_{\bf\mathfrak{g}}$-dual of $w_{\alpha}$ (i.e. the co-root of $\alpha$). Then $X_{\alpha}\in{\bf\mathfrak{a}}^{+}\cap{\bf\mathfrak{a}}_{\Theta}$ and using the notation of (3.10) we find $\Theta_{X_{\alpha}}=\{\alpha\}$, therefore
${\mathcal{F}}_{\alpha}={\mathcal{F}}_{X_{\alpha}}~{}.$ (3.18)
Given a subset $\Theta\subset\Delta$, the vectors $X_{\alpha}\in{\bf\mathfrak{a}}^{+}$ for $\alpha\in\Theta$ form a basis of ${\bf\mathfrak{a}}_{\Theta}=\cap_{\alpha\in\Delta\setminus\Theta}\ker\alpha$, and from the definition of the fundamental weights $w_{\alpha}$ one easily infers that their restricttions to ${\bf\mathfrak{a}}_{\Theta}$ form a basis of the dual space ${\bf\mathfrak{a}}_{\Theta}^{*}$:
Lemma 3.24.
For all $\alpha\in\Theta$ and $X\in{\bf\mathfrak{a}}$ one has $w_{\alpha}(X)=w_{\alpha}(p_{\Theta}(X))$, where $p_{\Theta}:{\bf\mathfrak{a}}\to{\bf\mathfrak{a}}_{\Theta}$ is the orthogonal projection with respect to the Killing form $\mathrm{B}_{\bf\mathfrak{g}}$.∎
3.7.3. Class functions on $G$
A restricted root space decomposition is often used in order to define certain class functions (i.e. functions on $G$ that are invariant under conjugation) by evaluating some element of ${\bf\mathfrak{a}}^{\ast}$ on the Jordan projection $\lambda(g)\in{\bf\mathfrak{a}}$ of an element $g\in G$. We will see that they can be defined using a flag manifold as the starting point. The first example consists in applying a simple restricted root.
Definition 3.25.
Given $\alpha\in\Delta$, the $\alpha$-proximality class function is
$\mathcal{P}_{\alpha}:\left\{\begin{array}[]{ccc}G&\to&\mathbb{R}_{\geq 0}\\ g&\mapsto&\alpha(\lambda(g)).\end{array}\right.$
Lemma 3.26.
The class function $\mathcal{P}_{\alpha}$ only depends on the simple flag manifold ${\mathcal{F}}_{\alpha}$, i.e. it is independent of the restricted root space decomposition used to define it.
Proof.
As a consequence of Fact 3.11, any two possible choices of Weyl chambers ${\bf\mathfrak{a}}_{1}^{+},\mathfrak{a}_{2}^{+}\subset{\bf\mathfrak{g}}$ are related by $\mathfrak{a}_{2}^{+}=\mathrm{Ad}(g_{0}){\bf\mathfrak{a}}_{1}^{+}$ for some $g_{0}\in G$. Now, the corresponding Jordan projections $\lambda_{i}(g)$ and restricted roots $\alpha_{i}\in{\bf\mathfrak{a}}_{i}^{*}$ defined using ${\bf\mathfrak{a}}^{+}_{i}$, $i=1,2$, differ by
$\lambda_{2}(g)=\mathrm{Ad}(g_{0})\lambda_{1}(g)~{};\qquad\alpha_{2}=\alpha_{1}\circ\mathrm{Ad}(g_{0}^{-1})~{};$
so that $\alpha_{1}(\lambda_{1}(g))=\alpha_{2}(\lambda_{2}(g))$ as claimed. ∎
This ultimately means that $\mathcal{P}_{\alpha}$ has a geometric meaning, this will be made more precise in Lemma 3.41.
Next, we consider an arbitrary flag manifold ${\mathcal{F}}$ and the associated subset $\Theta\subset\Delta$. Let $Y_{\Theta}\in\mathfrak{a}_{\Theta}$ be the unique vector such that $\alpha(Y_{\Theta})=1$ for all $\alpha\in\Theta$ and define
$\Sigma^{+}_{\Theta}:=\left\{\alpha\in\Sigma\,\middle|\,\alpha(Y_{\Theta})>0\right\},$
so that
$\mathfrak{n}_{\mathfrak{p}_{\Theta}}=\bigoplus_{\alpha\in\Sigma^{+}_{\Theta}}{\bf\mathfrak{g}}_{\alpha}.$
Definition 3.27.
For a flag manifold ${\mathcal{F}}$ associated to a subset $\Theta\subset\Delta$, the Jacobian class function of ${\mathcal{F}}$ is
$\mathcal{J}_{\mathcal{F}}:\left\{\begin{array}[]{ccc}G&\to&\mathbb{R}_{\geq 0}\\ g&\mapsto&2\varrho_{\Theta}\big{(}p_{\Theta}(\lambda(g))\big{)},\end{array}\right.$ (3.19)
where $p_{\Theta}:\mathfrak{a}\to\mathfrak{a}_{\Theta}$ is the Killing-orthogonal projection and
$\varrho_{\Theta}:=\frac{1}{2}\sum_{\alpha\in\Sigma^{+}_{\Theta}}\dim\mathfrak{g}_{\alpha}\alpha\quad\in{\bf\mathfrak{a}}^{\ast}.$ (3.20)
Here again the geometric terminology will be justified later, in Lemma 3.39.
Lemma 3.28.
The class function $\mathcal{J}_{{\mathcal{F}}}$ only depends on the flag manifold ${\mathcal{F}}$, i.e. it is independent of the restricted root space decomposition used to define it.
Proof.
The same proof as for Lemma 3.26 applies here. ∎
The two class functions introduced above satisfy a simple symmetry when passing to opposite flag manifolds.
Lemma 3.29.
The proximality class functions of $\alpha\in\Delta$ and its opposite $\iota_{o}(\alpha)$ are related by
$\mathcal{P}_{\alpha}(g)=\mathcal{P}_{\iota_{o}(\alpha)}(g^{-1})\qquad\forall\;g\in G.$
The Jacobian class functions of a pair of opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ are related by
$\mathcal{J}_{{\mathcal{F}}^{+}}(g)=\mathcal{J}_{{\mathcal{F}}^{-}}(g^{-1})\qquad\forall\;g\in G.$
Proof.
Both follow from the simple observation that
$\iota_{o}(\alpha)\big{(}\lambda(g)\big{)}=\alpha\big{(}\iota_{o}(\lambda(g))\big{)}=\alpha\big{(}\lambda(g^{-1})\big{)},$
where we used (3.16) in the last step. ∎
For a simple flag manifold ${\mathcal{F}}_{\alpha}$ the Jacobian and proximality class functions $\mathcal{J}_{{\mathcal{F}}_{\alpha}}$ and $\mathcal{P}_{\alpha}$ are in general not multiples of each other because the element $\varrho_{\{\alpha\}}$ turns out to be a multiple of the fundamental weight $w_{\alpha}$ rather than the root $\alpha$:
Lemma 3.30.
For each $\alpha\in\Delta$ there is a positive integer $m_{\alpha}\in\mathbb{N}=\{1,2,\ldots\}$ such that
$\varrho_{\{\alpha\}}=\frac{m_{\alpha}}{2}w_{\alpha}.$ (3.21)
Proof.
First we rewrite the index set $\Sigma^{+}_{\{\alpha\}}$ in the sum defining $\varrho_{\{\alpha\}}$. Note that the $\mathrm{B}_{\bf\mathfrak{g}}$-dual $X_{\alpha}$ of $w_{\alpha}$ lies in ${\bf\mathfrak{a}}^{+}$. For all $\beta\in\Sigma$ we have $\beta(X_{\alpha})=\mathrm{B}_{\bf\mathfrak{g}}(\beta,w_{\alpha})$ hence
$2\varrho_{\{\alpha\}}=\sum_{\begin{subarray}{c}\beta\in\Sigma:\\ \mathrm{B}_{\bf\mathfrak{g}}(\beta,w_{\alpha})>0\end{subarray}}\dim\mathfrak{g}_{\beta}\beta.$ (3.22)
Now, given any two simple roots $\alpha^{\prime},\alpha^{\prime\prime}\in\Delta$, the root reflection $r_{\alpha^{\prime}}:{\bf\mathfrak{a}}^{\ast}\to{\bf\mathfrak{a}}^{\ast}$ is an orthogonal transformation with respect to $\mathrm{B}_{\bf\mathfrak{g}}$ acting on the weight $w_{\alpha^{\prime\prime}}$ according to
$r_{\alpha^{\prime}}(w_{\alpha^{\prime\prime}})=w_{\alpha^{\prime\prime}}-\delta_{\alpha^{\prime},\alpha^{\prime\prime}}{\alpha^{\prime\prime}},$ (3.23)
where $\delta_{\alpha^{\prime},\alpha^{\prime\prime}}=1$ if $\alpha^{\prime}=\alpha^{\prime\prime}$ and $\delta_{\alpha^{\prime},\alpha^{\prime\prime}}=0$ otherwise. In particular, when $\alpha^{\prime}\neq\alpha$, every $\beta\in\Sigma$ satisfies
$\mathrm{B}_{\bf\mathfrak{g}}(\beta,w_{\alpha})>0\iff\mathrm{B}_{\bf\mathfrak{g}}(r_{\alpha^{\prime}}(\beta),w_{\alpha})>0.$
Thus (3.22) shows that $\varrho_{{\mathcal{F}}_{\alpha}}$ is invariant under all root reflections besides that of $\alpha$:
$r_{\alpha^{\prime}}(\varrho_{\{\alpha\}})=\varrho_{\{\alpha\}}\qquad\forall\;\alpha^{\prime}\in\Delta\setminus\{\alpha\}.$ (3.24)
This implies that $2\varrho_{\{\alpha\}}$ is an integer multiple of $w_{\alpha}$. Indeed, since the fundamental weights form a basis of ${\bf\mathfrak{a}}^{\ast}$ in which each root has integer coefficients, and $2\varrho_{\{\alpha\}}$ is by definition a linear combination of roots weighted by integers, we can write
$2\varrho_{\{\alpha\}}=\sum_{\alpha^{\prime}\in\Delta}m_{\alpha^{\prime}}w_{\alpha^{\prime}}$
with unique coefficients $m_{\alpha^{\prime}}\in\mathbb{Z}$. If $\alpha^{\prime}\in\Delta\setminus\{\alpha\}$, then (3.24) and (3.23) imply
$m_{\alpha^{\prime}}=m_{\alpha^{\prime}}(1-C_{\alpha^{\prime},\alpha^{\prime}}),$
where $C_{\alpha^{\prime},\alpha^{\prime}}$ is the coefficient in front of $w_{\alpha^{\prime}}$ when writing $\alpha^{\prime}$ as a linear combination of the fundamental weights. Recalling the definition of $w_{\alpha^{\prime}}$ one finds that $C_{\alpha^{\prime},\alpha^{\prime}}=2$, so $m_{\alpha^{\prime}}=0$.
It remains to show that $m_{\alpha}>0$. To this end, we apply $\mathrm{B}_{\bf\mathfrak{g}}$ to $2\varrho_{\{\alpha\}}$ and $w_{\alpha}$:
$\mathrm{B}_{\bf\mathfrak{g}}(2\varrho_{\{\alpha\}},w_{\alpha})=m_{\alpha}\underbrace{\mathrm{B}_{\bf\mathfrak{g}}(w_{\alpha},w_{\alpha})}_{>0}=\sum_{\begin{subarray}{c}\beta\in\Sigma:\\ \mathrm{B}_{\bf\mathfrak{g}}(\beta,w_{\alpha})>0\end{subarray}}\underbrace{\mathrm{B}_{\bf\mathfrak{g}}(\beta,w_{\alpha})}_{>0}\underbrace{\dim\mathfrak{g}_{\beta}}_{>0}.$
The sum’s index set is non-empty, it contains $\alpha$. Thus $m_{\alpha}>0$ and the proof is finished. ∎
For general flag manifolds we have the following fact about Jacobian class functions:
Lemma 3.31.
Let ${\mathcal{F}}$ be a flag manifold associated to a subset $\Theta\subset\Delta$. There is a collection of positive rational numbers $\mathbf{r}=(r_{\alpha})_{\alpha\in\Theta}\in\mathbb{Q}_{>0}^{\Theta}$ such that $\mathcal{J}_{{\mathcal{F}}}=\sum_{\alpha\in\Theta}r_{\alpha}\mathcal{J}_{{\mathcal{F}}_{\alpha}}$.
Proof.
The linear form $\varrho_{\Theta}\circ p_{\Theta}\in\mathfrak{a}^{\ast}$ vanishes on $\ker p_{\Theta}={\bf\mathfrak{a}}_{\Theta}^{\perp}=\cap_{\alpha\in\Theta}\ker w_{\alpha}$, so from Lemma 3.30 and simple linear algebra we get the existence of a tuple of real numbers $\mathbf{r}=(r_{\alpha})_{\alpha\in\Theta}\in\mathbb{R}^{\Theta}$ such that $\mathcal{J}_{{\mathcal{F}}}=\sum_{\alpha\in\Theta}r_{\alpha}\mathcal{J}_{{\mathcal{F}}_{\alpha}}$. In fact these numbers must be rational because weights and roots are rational linear combinations of each other and the ratio obtained in (3.21) is rational. Their positivity comes from the fact that $X_{w_{\beta}}\in{\bf\mathfrak{a}}^{+}\setminus\{0\}$ for any simple restricted root $\beta$. ∎
Example 3.32.
Still using the notations of Example 3.15 for $\operatorname{SL}(d,\mathrm{k})$, the fundamental weight $w_{\alpha_{j}}$ corresponding to the simple restricted root $\alpha_{j}$ is given by
$w_{\alpha_{j}}\Big{(}\mathrm{Diag}(x_{1},\dots,x_{d})\Big{)}=x_{1}+\cdots+x_{j}\,,$
and we find that $m_{\alpha_{j}}=d$ for $\mathrm{k}=\mathbb{R}$ and $m_{\alpha_{j}}=2d$ for $\mathrm{k}=\mathbb{C}$.
3.7.4. Jordan projections within a Levi subgroup
The Jordan decomposition also holds for reductive groups [Vog00, Prop. 2.10], [Mil17, Ch. 9]. Considering a Levi subgroup $L<G$, this means that the Jordan decomposition $g=g_{e}g_{h}g_{u}$ in $G$ of $g\in L$ produces $g_{e},g_{h},g_{u}\in L$.
Since we still consider a fixed restricted root space decomposition, let us consider the standard Levi subgroup $L_{\Theta}$ associated to a set $\Theta\subset\Delta$ of simple restricted roots, and investigate the difference between the Jordan projection $\lambda_{L_{\Theta}}$ within the reductive group $L_{\Theta}$ and the restriction to $L_{\Theta}$ of the Jordan projection $\lambda:G\to{\bf\mathfrak{a}}^{+}$ of $G$. The first is that they do not have the same target space. Indeed, ${\bf\mathfrak{a}}\subset\mathfrak{l}_{\Theta}$ is still a maximal abelian subspace of $\mathfrak{l}_{\Theta}\cap\mathfrak{k}^{\perp}$, but ${\bf\mathfrak{a}}^{+}$ is not a fundamental domain of the corresponding Weyl group $W_{\Theta}$. Instead we consider
${\bf\mathfrak{a}}^{+}_{L_{\Theta}}:={\bf\mathfrak{a}}_{\Theta}\oplus{\bf\mathfrak{a}}_{\Delta\setminus\Theta}^{+}=\left\{X\in{\bf\mathfrak{a}}\,\middle|\,\forall\alpha\in\Delta\setminus\Theta~{}\alpha(X)\geq 0\right\}\subset{\bf\mathfrak{a}}~{},$ (3.25)
where ${\bf\mathfrak{a}}_{\Theta}=\cap_{\alpha\in\Delta\setminus\Theta}\ker\alpha$ as before, and ${\bf\mathfrak{a}}_{\Delta\setminus\Theta}^{+}:={\bf\mathfrak{a}}_{\Delta\setminus\Theta}\cap{\bf\mathfrak{a}}^{+}$. The Jordan projection
$\lambda_{L_{\Theta}}:L_{\Theta}\to{\bf\mathfrak{a}}_{L_{\Theta}}^{+}$
is the map sending $g\in L_{\Theta}$ to the unique $X\in{\bf\mathfrak{a}}_{L_{\Theta}}^{+}$ such that $\exp(X)$ is conjugate within $L_{\Theta}$ to the hyperbolic component $g_{h}\in L_{\Theta}$ of $g$. Note that ${\bf\mathfrak{a}}^{+}_{L_{\Theta}}$ contains the full linear subspace ${\bf\mathfrak{a}}_{\Theta}$. Since this space lies in the center of $\mathfrak{l}_{\Theta}$, we cannot conjugate its elements into ${\bf\mathfrak{a}}_{\Theta}^{+}={\bf\mathfrak{a}}_{\Theta}\cap{\bf\mathfrak{a}}^{+}$ within $L_{\Theta}$. In particular, for $X\in{\bf\mathfrak{a}}_{L_{\Theta}}^{+}$, the vectors $X=\lambda_{L_{\Theta}}(\exp(X))$ and $\lambda(\exp(X))\in{\bf\mathfrak{a}}^{+}$ are usually distinct, and we have the equivalence
$\lambda(\exp(X))=X\iff X\in{\bf\mathfrak{a}}^{+}~{}.$ (3.26)
The Jordan projection $\lambda_{L_{\Theta}}:L_{\Theta}\to{\bf\mathfrak{a}}_{L_{\Theta}}^{+}$ can be composed with the Killing-orthogonal projection $p_{\Theta}:\mathfrak{a}\to\mathfrak{a}_{\Theta}$, thus defining a projection into the split part $\mathfrak{a}_{\Theta}=\mathfrak{z}_{\mathrm{split}}(\mathfrak{l}_{\Theta})$ of the center of the Lie algebra $\mathfrak{l}_{\Theta}$ of $L_{\Theta}$.
Definition 3.33.
The split central projection of $L_{\Theta}$ is the map
$\lambda_{L_{\Theta}}^{Z}=p_{\Theta}\circ\lambda_{L_{\Theta}}:L_{\Theta}\to{\bf\mathfrak{a}}_{\Theta}=\mathfrak{z}_{\mathrm{split}}(\mathfrak{l}_{\Theta})\,.$ (3.27)
A noticeable difference with Jordan projections considered so far is that $\lambda_{L_{\Theta}}^{Z}$ is a group homomorphism onto the additive group ${\bf\mathfrak{a}}_{\Theta}$. Indeed, the Langlands decomposition [Kna02, VII.7, Prop. 7.82] tells us there is a reductive subgroup $M_{\Theta}<L_{\Theta}$ with compact center such that $L_{\Theta}$ decomposes into the direct product
$L_{\Theta}=Z_{\mathrm{split}}(L_{\Theta})M_{\Theta}\,,$ (3.28)
and $\exp\circ\lambda_{L_{\Theta}}^{Z}$ is the projection onto the first factor in this direct product. The translation of Definition 3.27 in the framing of $L_{\Theta}$ therefore gives not just a class function, but an additive character of $L_{\Theta}$.
Definition 3.34.
Let $L_{\Theta}<G$ be a standard Levi subgroup. The Jacobian character of $L_{\Theta}$ is the function
$\mathbf{J}_{L_{\Theta}}:\left\{\begin{array}[]{ccc}L_{\Theta}&\to&\mathbb{R}\\ g&\mapsto&2\varrho_{\Theta}\big{(}\lambda^{Z}_{L_{\Theta}}(g)\big{)}.\end{array}\right.$
Since any element $X\in{\bf\mathfrak{g}}_{\mathrm{hyp}}$ is contained in a Weyl chamber ${\bf\mathfrak{a}}^{+}$ for an appropriate root space decomposition, any Levi subgroup $L<G$ can be considered as a standard Levi subgroup, thus defining a split central projection $\lambda^{Z}_{L}\in\mathrm{Hom}(L,\mathfrak{z}_{\mathrm{split}}(\mathfrak{l}))$ and a Jacobian character $\mathbf{J}_{L}\in\mathrm{Hom}(L,\mathbb{R})$.
Lemma 3.35.
The split central projection $\lambda^{Z}_{L}\in\mathrm{Hom}(L,\mathfrak{z}_{\mathrm{split}}(\mathfrak{l}))$ and the Jacobian character $\mathbf{J}_{L}\in\mathrm{Hom}(L,\mathbb{R})$ of a Levi subgroup $L<G$ does not depend on the restricted root space decomposition used to define it.
Proof.
Repeat the proof of Lemma 3.26 once again. ∎
3.8. Linear actions on tangent spaces of flag manifolds
[…]
$\mathfrak{p}^{+}\oplus\mathfrak{n}_{\mathfrak{p}^{-}}=\mathfrak{n}_{\mathfrak{p}^{+}}\oplus\mathfrak{p}^{-}={\bf\mathfrak{g}}.$ (3.29)
Every element $X\in\mathfrak{g}$ induces fundamental vector fields on ${\mathcal{F}}^{\pm}$ and ${{\mathcal{F}}^{\pitchfork}}$, which we shall all denote by $\overline{X}$, via the formula $\overline{X}(x)=\left.\frac{d\,}{dt}\right|_{t=0}\exp(tX)\cdot x$. The kernel of the evaluation map
$\begin{split}{\bf\mathfrak{g}}&\to T_{\mathbf{x}^{\pm}}\mathcal{F}^{\pm}\\ X&\mapsto\overline{X}(\mathbf{x}^{\pm})\end{split}$ (3.30)
is the Lie algebra of the stabilizer of $\mathbf{x}^{\pm}$, i.e. $\mathfrak{p}^{\pm}$. Therefore, the decomposition (3.29) provides us with linear isomorphisms
$T_{\mathbf{x}^{+}}{\mathcal{F}}^{+}\simeq\mathfrak{n}_{\mathfrak{p}^{-}},\qquad T_{\mathbf{x}^{-}}{\mathcal{F}}^{-}\simeq\mathfrak{n}_{\mathfrak{p}^{+}},\qquad T_{\mathbf{x}}{{\mathcal{F}}^{\pitchfork}}\simeq\mathfrak{n}_{\mathfrak{p}^{-}}\oplus\mathfrak{n}_{\mathfrak{p}^{+}}.$ (3.31)
Remark 3.1.
Note that these isomorphisms swap the $+$ and -- signs in the upper index, which will be important to keep track of when we apply them.
Moreover, consider the parabolic subgroups $P_{\mathbf{x}^{\pm}}:=\mathrm{Stab}_{G}(\mathbf{x}^{\pm})=\mathrm{Stab}_{G}({\bf\mathfrak{n}}_{\mathfrak{p}^{\pm}})$. Then for every $g\in P_{\mathbf{x}^{+}}$ the fact that $g$ fixes ${\bf\mathfrak{n}}_{\mathfrak{p}^{+}}$ allows us to identify both $T_{\mathbf{x}^{-}}{\mathcal{F}}^{-}$ and $T_{g\cdot\mathbf{x}^{-}}{\mathcal{F}}^{-}$ with ${\bf\mathfrak{n}}_{\mathfrak{p}^{+}}$ via the evaluation map (3.30), and analogously if $g\in P_{\mathbf{x}^{-}}$. In both cases, the derivative $dg:T_{\mathbf{x}^{\mp}}{\mathcal{F}}^{\mp}\to T_{g\cdot\mathbf{x}^{\mp}}{\mathcal{F}}^{\mp}$ of the $g$-action then becomes a linear map ${\bf\mathfrak{n}}_{\mathfrak{p}^{\pm}}\to{\bf\mathfrak{n}}_{\mathfrak{p}^{\pm}}$ given by the restricted adjoint action
$\begin{split}\mathfrak{n}_{\mathfrak{p}^{\pm}}&\to\mathfrak{n}_{\mathfrak{p}^{\pm}}\\ X&\mapsto\mathrm{Ad}(g)X.\end{split}$ (3.32)
This allows us to give a geometric interpretation of the Jacobian characters of Levi subgroups:
Lemma 3.36.
[…]
$e^{\mathbf{J}_{L}(g)}=|\mathrm{det}\,(d_{\mathbf{x}^{-}}g)|=|\mathrm{det}\,(\mathrm{Ad}(g)|_{\mathfrak{n}})|,$
where $d_{\mathbf{x}^{-}}g\in\mathrm{GL}\left(T_{\mathbf{x}^{-}}{\mathcal{F}}^{-}\right)$ denotes the differential of the action of $g$ at its fixed point $\mathbf{x}^{-}\in{\mathcal{F}}^{-}$.
Proof.
The second equality comes from the discussion above. Considering an element $Y\in{\bf\mathfrak{g}}_{\mathrm{hyp}}$ such that $\mathrm{Stab}_{G}(\mathbf{x}^{+})=P_{Y}$ and $\mathrm{Stab}_{G}(\mathbf{x}^{-})=P_{-Y}$, and a restricted root space decomposition such that $Y\in{\bf\mathfrak{a}}^{+}$, we may write $L=L_{\Theta}$ for some subset $\Theta\subset\Delta$. Now consider the Jordan decomposition $g=g_{e}g_{h}g_{u}$ in the reductive Lie group $L_{\Theta}$. Since the elements $g_{e},g_{h},g_{u}\in L_{\Theta}<P_{\Theta}$ fix ${\bf\mathfrak{n}}$, we immediately have
$|\mathrm{det}\,(\mathrm{Ad}(g_{e})|_{\mathfrak{n}})|=\mathrm{det}\,(\mathrm{Ad}(g_{u})|_{\mathfrak{n}})=1~{},$
thus
$|\mathrm{det}\,(\mathrm{Ad}(g)|_{\mathfrak{n}})|=|\mathrm{det}\,(\mathrm{Ad}(g_{h})|_{\mathfrak{n}})|.$ (3.33)
Since we also have $\mathbf{J}_{L}(g)=\mathbf{J}_{L}(g_{h})$ by definition of the Jordan projection $\lambda_{L_{\Theta}}$, we may now assume that $g$ is hyperbolic, i.e. $g=g_{h}$. Furthermore conjugating in $L_{\Theta}$, we may assume that $g=\exp(X)$ for some $X\in{\bf\mathfrak{a}}^{+}_{L_{\Theta}}$, where ${\bf\mathfrak{a}}^{+}_{L_{\Theta}}$ was defined in (3.25).
It now comes from the very definition of $\varrho_{\Theta}=\frac{1}{2}\sum_{\alpha\in\Sigma^{+}_{\Theta}}\dim\mathfrak{g}_{\alpha}\alpha$ that the determinant of the action of $\mathrm{Ad}(\exp(X))=\exp(\operatorname{ad}_{X})$ on $\mathfrak{n}=\bigoplus_{\alpha\in\Sigma_{\Theta}^{+}}\mathfrak{g}_{\alpha}$ is
$\mathrm{det}\,\left(\mathrm{Ad}(\exp(X))|_{\mathfrak{n}}\right)=e^{2\varrho_{\Theta}\circ p_{\Theta}(X)}~{},$
and $2\varrho_{\Theta}\circ p_{\Theta}(X)$ is the definition of $\mathbf{J}_{L_{\Theta}}(\exp(X))$. ∎
In order to give a similar geometric interpretation of the class functions $\mathcal{P}_{\alpha}$ and $\mathcal{J}_{\mathcal{F}}$ defined on the whole group $G$, we must first guarantee the existence of fixed points in flag manifolds, then select a preferred fixed point as there can be several. This is done through the notion of proximality.
Definition 3.37.
Let ${\mathcal{F}}$ be a flag manifold, $\mathbf{x}\in{\mathcal{F}}$, and $g\in G$. We say that $\mathbf{x}$ is an attracting fixed point of $g$ if $g\in\mathrm{Stab}_{G}(\mathbf{x})$ and the linear action of $g$ on $T_{\mathbf{x}}{\mathcal{F}}$ given by the derivative of the $G$-action has spectral radius $<1$.
An element $g\in G$ is called ${\mathcal{F}}$-proximal if $g$ has an attracting fixed point in ${\mathcal{F}}$.
Numerous useful properties of ${\mathcal{F}}$-proximal elements are proved in [GGKW17, Lemma 2.26, Prop. 3.3 (c)]. For instance, it is shown there that every ${\mathcal{F}}$-proximal element $g\in G$ has a unique attracting fixed point in ${\mathcal{F}}$, and that
$\begin{split}g\text{ is ${\mathcal{F}}$-proximal}&\iff g^{-1}\text{ is $\overline{\mathcal{F}}$-proximal}\\ &\iff\mathcal{P}_{\alpha}(g)>0~{}\forall\alpha\in\Theta\,.\end{split}$ (3.34)
Definition 3.38.
Consider a pair of opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ and the transverse flag space ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$. An element $g\in G$ is ${{\mathcal{F}}^{\pitchfork}}$-proximal if $g$ is ${\mathcal{F}}^{+}$-proximal (which by (3.34) is equivalent to $g^{-1}$ being ${\mathcal{F}}^{-}$-proximal) and the attracting fixed points $g^{+}\in{\mathcal{F}}^{+}$, $g^{-}\in{\mathcal{F}}^{-}$ of $g$ and $g^{-1}$ satisfy $(g^{+},g^{-})\in{{\mathcal{F}}^{\pitchfork}}$. In this case, $g^{-}$ is called the repelling fixed point of $g$.
With this terminology, we can provide a geometric interpretation of the Jacobian class functions of flag manifolds:
Lemma 3.39.
[…]
$e^{\mathcal{J}_{{\mathcal{F}}^{+}}(g)}=\big{|}\mathrm{det}\,\big{(}\left.\mathrm{Ad}(g)\right|_{\mathfrak{n}}\big{)}\big{|}.$ (3.35)
Proof.
Let $L=\mathrm{Stab}_{G}(\mathbf{x})$, so that according to Lemma 3.36 we only need to show that $\mathcal{J}_{{\mathcal{F}}^{+}}(g)=\mathbf{J}_{L}(g)$. Repeating the process used in the proof of Lemma 3.36, we can consider a fixed restricted root space decomposition such that $L=L_{\Theta}$ is a standard Levi subgroup and reduce the study to the case of $g=\exp(X)$ with $X\in{\bf\mathfrak{a}}_{L_{\Theta}^{+}}$.
Assuming that $\mathbf{x}^{-}$ is the repelling fixed point of $g$ means that all the eigenvalues of the action of $\exp(X)$ on the tangent space $T_{\mathbf{x}^{-}}\mathcal{F}^{-}$ are greater than one (in absolute value). But the fact that $\exp(X)\in L_{\Theta}$ means that this derivative is conjugate to the action of $\mathrm{Ad}(\exp(X))=\exp(\operatorname{ad}_{X})$ on $\mathfrak{n}$. The eigenvalues of the action of $\operatorname{ad}_{X}$ on $\mathfrak{n}=\bigoplus_{\alpha\in\Sigma_{\Theta}^{+}}\mathfrak{g}_{\alpha}$ are therefore positive, hence $\alpha(X)>0$ for any $X\in\Theta$. But we already had that $X\in{\bf\mathfrak{a}}_{L_{\Theta}}^{+}$, i.e. that $\alpha(X)\geq 0$ for any $\alpha\in\Delta\setminus\Theta$, thus $X\in{\bf\mathfrak{a}}^{+}$ and $\lambda(\exp(X))=X=\lambda_{L_{\Theta}}(\exp(X))$ by (3.26). It follows that $\mathcal{J}_{{\mathcal{F}}^{+}}(g)=\mathbf{J}_{L}(g)$. ∎
Remark 3.2.
Note that from Lemma 3.36 we also obtain $e^{\mathcal{J}_{{\mathcal{F}}^{+}}(g)}=\big{|}\mathrm{det}\,\big{(}d_{\mathbf{x}^{-}}g\big{)}\big{|}$. This geometric interpretation of the Jacobian class function does not involve the attracting fixed point, but the proof relies on its existence and transversality to the repelling fixed point.
For later use, we record the following elementary naturality property of proximality.
Lemma 3.40.
Let ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ be a pair of opposite flag manifolds corresponding to subsets $\Theta\subset\Delta$ and $\iota_{o}(\Theta)$ respectively. Denote the transverse flag spaces by ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ and ${\mathcal{F}}_{\alpha}^{\pitchfork}={\mathcal{F}}_{\alpha}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}_{\iota_{o}(\alpha)}$ for $\alpha\in\Theta$. Then any ${{\mathcal{F}}^{\pitchfork}}$-proximal element $g\in G$ with attracting/repelling fixed points $(\mathbf{x}^{+},\mathbf{x}^{-})\in{{\mathcal{F}}^{\pitchfork}}$ is also ${\mathcal{F}}^{\pitchfork}_{\alpha}$-proximal with attracting/repelling fixed points $(\mathbf{x}^{+}_{\alpha},\mathbf{x}^{-}_{\alpha})\in{\mathcal{F}}^{\pitchfork}_{\alpha}$ for all $\alpha\in\Theta$, where $\mathbf{x}^{+}_{\alpha}=\Pi^{{\mathcal{F}}^{+}}_{\alpha}(\mathbf{x}^{+})$ and $\mathbf{x}^{-}_{\alpha}=\Pi^{{\mathcal{F}}^{-}}_{\iota_{o}(\alpha)}(\mathbf{x}^{-})$ are the images under the unique projections introduced in (3.15).
Proof.
The fact that $(\mathbf{x}^{+}_{\alpha},\mathbf{x}^{-}_{\alpha})\in{\mathcal{F}}^{\pitchfork}_{\alpha}$ comes from Corollary 3.20, and the contraction/dilation of the derivatives at these fixed points follows from the same property on ${\mathcal{F}}^{\pm}$ by simple linear algebra. ∎
Let us conclude this section with the geometric interpretation of the proximality class functions.
Lemma 3.41.
Let $\mathcal{F}_{\alpha}$ be a simple flag manifold and ${\mathcal{F}}^{\pitchfork}_{\alpha}={\mathcal{F}}_{\alpha}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}_{\iota_{o}(\alpha)}$ the corresponding transverse flag space. For any $\mathcal{F}^{\pitchfork}_{\alpha}$-proximal element $g\in G$, we have
$\mathcal{P}_{\alpha}(g)=-\lambda_{1}\big{(}d_{\mathbf{x}^{-}}g^{-1}\big{)},$
where $d_{\mathbf{x}^{-}}g^{-1}\in\mathrm{GL}\left(T_{\mathbf{x}^{-}}{\mathcal{F}}_{\iota_{o}(\alpha)}\right)$ denotes the differential of the action of $g^{-1}$ at the repelling fixed point $\mathbf{x}^{-}\in{\mathcal{F}}_{\iota_{o}(\alpha)}$ of $g$ and $\lambda_{1}(d_{\mathbf{x}^{-}}g^{-1})$ is the logarithm of its spectral radius.
Remark 3.3.
This means that if we rank the logarithms of moduli of eigenvalues
$\lambda_{1}(d_{\mathbf{x}^{-}}g)\geq\cdots\geq\lambda_{n}(d_{\mathbf{x}^{-}}g)~{},$
then $\mathcal{P}_{\alpha}(g)=\lambda_{n}(d_{\mathbf{x}^{-}}g)$, i.e. $\mathcal{P}_{\alpha}(g)$ measures the dilation of $g$ near its repelling fixed point.
Proof.
Just as in the proof of Lemma 3.39, $\mathcal{F}^{\pitchfork}_{\alpha}$-proximality means that we can work in the context of a fixed restricted root space decomposition and assume that $g=\exp(X)$ for some $X\in{\bf\mathfrak{a}}^{+}$, in which case $\mathcal{P}_{\alpha}(g)=\alpha(X)$ and the derivative $d_{\mathbf{x}^{-}}g$ is conjugate to the action of $\mathrm{Ad}(\exp(X))=\exp(\operatorname{ad}_{X})$ on the nilpotent radical $\mathfrak{n}=\bigoplus_{\beta\in\Sigma_{\{\alpha\}}^{+}}{\bf\mathfrak{g}}_{\beta}$, so we find
$\lambda_{n}(d_{\mathbf{x}^{-}}g)=\min\left\{\beta(X)\,\middle|\,\beta\in\Sigma_{\{\alpha\}}^{+}\right\}\,.$
Now every $\beta\in\Sigma_{\{\alpha\}}^{+}\subset\Sigma^{+}$ can be expressed as $\beta=\alpha+\beta^{\prime}$ where $\beta^{\prime}$ is a combination of elements of $\Delta$ with non-negative coefficients, thus $\lambda_{n}(d_{\mathbf{x}^{-}}g)=\alpha(X)$. ∎
4. Affine line bundles and period functions
[…]
4.1. $G$-equivariant affine line bundles
We will apply the following general definition with ${\mathbb{F}}$ either a flag manifold or a transverse flag space:
Definition 4.1.
Let ${\mathbb{F}}$ be a homogeneous $G$-manifold.
  • A $G$-equivariant affine line bundle ${\mathbb{L}}\to{\mathbb{F}}$ is a principal $\mathbb{R}$-bundle over ${\mathbb{F}}$ such that:
    • ${\mathbb{L}}$ is a $G$-manifold.
    • The projection ${\mathbb{L}}\to{\mathbb{F}}$ is $G$-equivariant.
    • The $G$-action and the $\mathbb{R}$-action on ${\mathbb{L}}$ commute.
  • A morphism ${\mathbb{L}}\to{\mathbb{L}}^{\prime}$ of $G$-equivariant affine line bundles ${\mathbb{L}},{\mathbb{L}}^{\prime}$ over ${\mathbb{F}}$ is a $G$-equivariant map ${\mathbb{L}}\to{\mathbb{L}}^{\prime}$ that forms a morphism of principal $\mathbb{R}$-bundles over ${\mathbb{F}}$. We write ${\mathbb{L}}\simeq{\mathbb{L}}^{\prime}$ for an isomorphism of $G$-equivariant affine line bundles.
  • A $G$-equivariant affine line bundle ${\mathbb{L}}\to{\mathbb{F}}$ is trivial if ${\mathbb{L}}\simeq{\mathbb{F}}\times\mathbb{R}$, where on the right $G$ acts by the product of the $G$-action on ${\mathbb{F}}$ and the trivial action on $\mathbb{R}$.
  • The time-reversal $\overline{\mathbb{L}}\to{\mathbb{F}}$ of a $G$-equivariant affine line bundle ${\mathbb{L}}\to{\mathbb{F}}$ is obtained from ${\mathbb{L}}$ by flipping the sign in the $\mathbb{R}$-action, i.e., composing with the map $\mathbb{R}\owns t\mapsto-t\in\mathbb{R}$.
  • A $G$-equivariant affine line bundle ${\mathbb{L}}$ over ${\mathbb{F}}$ is homogeneous if the $G$-manifold ${\mathbb{L}}$ is homogeneous, i.e., the $G$-action on ${\mathbb{L}}$ is transitive. In this case, ${\mathbb{L}}\to{\mathbb{F}}$ is called a homogeneous affine line bundle over ${\mathbb{F}}$.
Note that a trivial $G$-equivariant affine line bundle is never homogeneous. Conversely, any non-trivial $G$-equivariant affine line bundle ${\mathbb{L}}\to{\mathbb{F}}$ is homogeneous.
4.2. Affine line bundles and oriented vector line bundles
Principal $\mathbb{R}$-bundles correspond to oriented vector line bundles [^2] We use the term vector line bundle for what is usually just called line bundle, i.e. a vector bundle of rank one, in order to draw a clear distinction from an affine line bundle., since the group $\mathbb{R}$ is isomorphic to the identity component $\mathrm{GL}^{+}(1,\mathbb{R})$. Note that an orientation on a vector line bundle $\mathcal{L}\to{\mathbb{F}}$ is just a smooth choice of a connected component $\mathcal{L}_{>0}$ of the complement of the zero section in each fiber. Then $\mathcal{L}_{>0}\to{\mathbb{F}}$ can be given the structure of a principal $\mathbb{R}$-bundle by defining the action of $t\in\mathbb{R}$ on $v\in\mathcal{L}_{>0}$ to be $v\cdot t=e^{t}v$.
Starting with a principal $\mathbb{R}$-bundle ${\mathbb{L}}\to{\mathbb{F}}$, the usual construction of an associated bundle
$\mathcal{L}=({\mathbb{L}}\times\mathbb{R})/\mathbb{R}\to{\mathbb{F}},$
where the action of $t\in\mathbb{R}$ on $(\hat{\mathbf{x}},s)\in{\mathbb{L}}\times\mathbb{R}$ is $(\hat{\mathbf{x}},s)\cdot t=(\hat{\mathbf{x}}\cdot t,e^{-t}s)$, has a natural orientation defined as
$\mathcal{L}_{>0}:=({\mathbb{L}}\times\mathbb{R}_{>0})/\mathbb{R}.$
In the case of $G$-equivariant bundles over a homogeneous $G$-manifold ${\mathbb{F}}$, both directions of this correspondence are compatible with $G$-actions. Note that the time reversal $\overline{\mathbb{L}}$ corresponds to switching the orientation of the vector line bundle $\mathcal{L}$ from $\mathcal{L}_{>0}$ to $\mathcal{L}_{<0}$.
While affine line bundles are more fitted to our dynamical intentions, since the $\mathbb{R}$-action can be interpreted as a flow, vector line bundles are more practical when it comes to algebraic constructions such as tensor products. Most of our focus will be on the affine point of view (including the current section), but the constructions in Section 5.1 will be achieved using oriented vector line bundles.
4.3. Homogeneous affine line bundles and additive characters
Consider the case of a $G$-homogeneous space ${\mathbb{F}}$, and fix a base point $\mathbf{x}\in{\mathbb{F}}$ so that we can identify ${\mathbb{F}}=G/H$ where $H<G$ is the stabilizer of $\mathbf{x}$. The following basic correspondence is well-known.
Lemma 4.2.
There is a bijective correspondence between Lie group homomorphisms $b:H\to\mathbb{R}$ (considered up to precomposition by an isomorphism) and (equivalence classes of) $G$-equivariant affine line bundles over $G/H$.∎
A classification of homogeneous affine line bundles over a given homogeneous space $G/H$ therefore reduces to a classification of Lie group homomorphisms $H\to\mathbb{R}$. In the case of a transverse flag space, the isotropy group is a Levi subgroup $L<G$. Recall from (3.28) that $L$ splits into a direct product $L=Z_{\mathrm{split}}(L)M_{L}$ where $M_{L}$ is a reductive group with compact center, so the map
$\left\{\begin{array}[]{ccc}\mathfrak{z}_{\mathrm{split}}(\mathfrak{l})^{*}&\to&\mathrm{Hom}(L,\mathbb{R})\\ \beta&\mapsto&\beta\circ\lambda^{Z}_{L}\end{array}\right.$ (4.1)
is an isomorphism, where $\lambda^{Z}_{L}:L\to\mathfrak{z}_{\mathrm{split}}(\mathfrak{l})$ is the split central projection.
For a standard Levi subgroup $L_{\Theta}$, a natural basis of the dual space of $\mathfrak{z}_{\mathrm{split}}(\mathfrak{l}_{\Theta})={\bf\mathfrak{a}}_{\Theta}$ is given by the fundamental weights $w_{\alpha}$, $\alpha\in\Theta$.
Corollary 4.3.
Consider a restricted root space decomposition, a subset $\Theta\subset\Delta$ and the corresponding standard Levi subgroup $L_{\Theta}$. There is a bijective correspondence between isomorphism classes of $G$-equivariant affine line bundles over the transverse flag space ${\mathcal{F}}_{\Theta}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}_{\iota_{o}(\Theta)}\simeq G/L_{\Theta}$ and linear combinations of the fundamental weights $w_{\alpha}$, $\alpha\in\Theta$.
4.4. The period function of a $G$-equivariant affine line bundle
We now give a second description of $G$-equivariant affine line bundles over a homogeneous space ${\mathbb{F}}$, where in contrast to Section 4.3 we do not need to fix a base point. To this end, we consider a real additive character
$\ell_{\mathbb{L}}(\mathbf{x},-)\in\mathrm{Hom}(\mathrm{Stab}_{G}(\mathbf{x}),\mathbb{R})$
for each $\mathbf{x}\in{\mathbb{F}}$, defined through the action
$g\cdot\hat{\mathbf{x}}=\hat{\mathbf{x}}\cdot\ell_{\mathbb{L}}(\mathbf{x},g)\,,\quad\forall\hat{\mathbf{x}}\in{\mathbb{L}}_{\mathbf{x}}\,\forall g\in\mathrm{Stab}_{G}(\mathbf{x})\,.$
If we consider for some $h\in G$ the point $h\cdot\mathbf{x}\in{\mathbb{F}}$, then we get $\mathrm{Stab}_{G}(h\cdot\mathbf{x})=h\mathrm{Stab}_{G}(\mathbf{x})h^{-1}$ and for $g\in\mathrm{Stab}_{G}(\mathbf{x})$ we find
$\ell_{\mathbb{L}}(h\cdot\mathbf{x},hgh^{-1})=\ell_{\mathbb{L}}(\mathbf{x},g)$
because the $G$- and $\mathbb{R}$-actions on ${\mathbb{L}}$ commute. We equip the stabilizer bundle
$\mathrm{Stab}_{\mathbb{F}}:=\{(\mathbf{x},g)\in{\mathbb{F}}\times G\,|\,g\in\mathrm{Stab}_{G}(\mathbf{x})\}$
with the $G$-action $h\cdot(\mathbf{x},g):=(h\cdot\mathbf{x},hgh^{-1})$. The projection onto the first factor makes $\mathrm{Stab}_{{\mathbb{F}}}$ a $G$-equivariant fiber bundle over ${\mathbb{F}}$.
Definition 4.4.
The $G$-invariant function
$\ell_{\mathbb{L}}:\left\{\begin{array}[]{ccc}\mathrm{Stab}_{{\mathbb{F}}}&\to&\mathbb{R}\\ (\mathbf{x},g)&\mapsto&\ell_{\mathbb{L}}(\mathbf{x},g)\end{array}\right.$
is called the period function of the $G$-equivariant affine line bundle ${\mathbb{L}}$.
Period functions classify $G$-equivariant affine line bundles up to isomorphism:
Lemma 4.5.
For any two $G$-equivariant affine line bundles ${\mathbb{L}}$, ${\mathbb{L}}^{\prime}$ over a $G$-homogeneous space ${\mathbb{F}}$, the following statements are equivalent:
  1. (1)
    ${\mathbb{L}}\simeq{\mathbb{L}}^{\prime}$.
  2. (2)
    $\ell_{{\mathbb{L}}}=\ell_{{\mathbb{L}}^{\prime}}$.
  3. (3)
    There is a point $\mathbf{x}\in{\mathbb{F}}$ such that $\ell_{{\mathbb{L}}}(\mathbf{x},-)=\ell_{{\mathbb{L}}^{\prime}}(\mathbf{x},-):\mathrm{Stab}_{G}(\mathbf{x})\to\mathbb{R}$.
Proof.
The equivalence between (1) and (3) is Lemma 4.2, and it is straightforward to see that (1$\Rightarrow$(2$\Rightarrow$(3). ∎
A base-point free version of Lemma 4.2 therefore goes through the following definition:
Definition 4.6.
The space $\mathcal{P}({\mathbb{F}})$ of period functions of ${\mathbb{F}}$ is the space of $G$-invariant functions $\mathrm{Stab}_{\mathbb{F}}\to\mathbb{R}$ that restrict to a Lie homomorphism on each fiber.
Note that $\mathcal{P}({\mathbb{F}})$ is a finite dimensional real vector space.
Corollary 4.7.
Let ${\mathbb{F}}$ be a homogeneous $G$-manifold. The map ${\mathbb{L}}\mapsto\ell_{\mathbb{L}}$ establishes a bijective correspondence between isomorphism classes of $G$-equivariant affine line bundles over ${\mathbb{F}}$ and $\mathcal{P}({\mathbb{F}})$.∎
In particular, the vector space structure on the set of isomorphism classes of $G$-equivariant affine line bundles over ${\mathbb{F}}$ given by identifying it with additive characters in Lemma 4.2 does not depend on the choice of a base point.
Consider opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ and the transverse flag space ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$. The Jacobian characters of Levi subgroups can be combined into the period function
$\left\{\begin{array}[]{ccc}\mathrm{Stab}_{{\mathcal{F}}^{\pitchfork}}&\to&\mathbb{R}\\ (\mathbf{x},g)&\mapsto&\mathbf{J}_{\mathrm{Stab}_{G}(\mathbf{x})}(g).\end{array}\right.$
If we now consider the corresponding subsets $\Theta,\iota_{o}(\Theta)\subset\Delta$, then for each $\alpha\in\Theta$ there is a unique $G$-equivariant projection $\pi_{\alpha}:{{\mathcal{F}}^{\pitchfork}}\to{\mathcal{F}}^{\pitchfork}_{\alpha}={\mathcal{F}}_{\alpha}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}_{\iota_{o}(\alpha)}$ given by Corollary 3.20. We can use it to define
$\mathbf{J}_{{{\mathcal{F}}^{\pitchfork}}}^{\alpha}:\left\{\begin{array}[]{ccc}\mathrm{Stab}_{{\mathcal{F}}^{\pitchfork}}&\to&\mathbb{R}\\ (\mathbf{x},g)&\mapsto&\mathbf{J}_{\mathrm{Stab}_{G}(\pi_{\alpha}(\mathbf{x}))}(g).\end{array}\right.$
Proposition 4.8.
Consider opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ and their corresponding subsets $\Theta,\iota_{o}(\Theta)\subset\Delta$, and let ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$. The functions $(\mathbf{J}_{{{\mathcal{F}}^{\pitchfork}}}^{\alpha})_{\alpha\in\Theta}$ form a basis of $\mathcal{P}({{\mathcal{F}}^{\pitchfork}})$.
Proof.
It follows from Lemma 3.30 ($\varrho_{\{\alpha\}}$ and $w_{\alpha}$ are proportional) and Corollary 4.3. ∎
4.5. Sections and cocycles
An principal $\mathbb{R}$-bundle over any manifold is trivializable. However, a homogeneous affine line bundle cannot be equivariantly trivializable because the trivial bundle is not homogeneous. The failure of equivariance of a given trivialization (i.e. a section) is encoded in a cocycle.
Definition 4.9.
Let $\sigma$ be a section of a $G$-equivariant affine line bundle ${\mathbb{L}}\to{\mathbb{F}}$. The cocycle associated to $\sigma$ is the function $\mathcal{C}_{\sigma}:{\mathbb{F}}\times G\to\mathbb{R}$ defined by the relation
$g\cdot\sigma(\mathbf{x})=\sigma(g\cdot\mathbf{x})\cdot\mathcal{C}_{\sigma}(\mathbf{x},g)\quad\forall(\mathbf{x},g)\in{\mathbb{F}}\times G~{}.$
Note that if $g\cdot\mathbf{x}=\mathbf{x}$, the value of $\mathcal{C}_{\sigma}(\mathbf{x},\nu)$ does not depend on $\sigma$ since $\mathcal{C}_{\sigma}(\mathbf{x},\nu)=\ell_{\mathbb{L}}(\mathbf{x},g)$ in this case. So one can see $\mathcal{C}_{\sigma}$ as an extension of the period function $\ell_{\mathbb{L}}$ to ${\mathbb{F}}\times G$ that satisfies the cocycle relation
$\mathcal{C}_{\sigma}(\mathbf{x},g_{2}g_{1})=\mathcal{C}_{\sigma}(g_{1}\cdot\mathbf{x},g_{2})+\mathcal{C}_{\sigma}(\mathbf{x},g_{1})\quad\forall(\mathbf{x},g_{1},g_{2})\in{\mathbb{F}}\times G\times G~{}.$
Although the cocycle depends on the choice of a section, its asymptotic behavior when $g\to\infty$ in $G$ will not depend on this choice provided that ${\mathbb{F}}$ is compact (which is the case for flag manifolds, but not for transverse flag spaces).
Lemma 4.10.
Consider two sections $\sigma,\sigma^{\prime}$ of a $G$-equivariant affine line bundle ${\mathbb{L}}\to{\mathbb{F}}$. If the space ${\mathbb{F}}$ is compact, then the difference between the two cocycles
$\mathcal{C}_{\sigma}-\mathcal{C}_{\sigma^{\prime}}:{\mathbb{F}}\times G\to\mathbb{R}$
is uniformly bounded.
Proof.
Consider the function $\tau:{\mathbb{F}}\to\mathbb{R}$ such that $\sigma^{\prime}(\mathbf{x})=\sigma(\mathbf{x})\cdot\tau(\mathbf{x})$ for all $\mathbf{x}\in{\mathbb{F}}$. The difference between the two cocycles is $\mathcal{C}_{\sigma}(\mathbf{x},g)-\mathcal{C}_{\sigma^{\prime}}(\mathbf{x},g)=\tau(\mathbf{x})-\tau(g\cdot\mathbf{x})$. ∎
5. Densities and multidensities
We will apply the construction from Section 4.2 to density and multidensity bundles, which are natural families of oriented vector line bundles on any smooth manifold. We begin by quickly recalling their general definitions and relevant properties.
5.1. General definitions and properties
Let $E$ be a finite dimensional real vector space of dimension $d>0.$ Let $\mathcal{B}_{E}$ be the set of ordered bases of $E$, which is the unique open $\mathrm{GL}(E)$-orbit in the product of $d$ copies of $E$
Definition 5.1.
For $s\in\mathbb{R},$ an $s$-density on $E$ is a map $\nu:\mathcal{B}_{E}\rightarrow\mathbb{R}$ such that for every $g\in\mathrm{GL}(E):$
$\nu(g\cdot e_{1},\ldots,g\cdot e_{d})=\lvert\mathrm{det}(g)\rvert^{s}\nu(e_{1},\ldots,e_{d})$
for all $\mathbf{e}=(e_{1},\ldots,e_{d})\in\mathcal{B}_{E}.$
We write $\mathscr{D}^{s}(E)$ for the real vector space of $s$-densities on $E$, and $\mathscr{D}^{s}_{>0}(E)$ for the subset of positive $\mathbb{R}_{>0}$-valued $s$-densities.
Note that $\mathscr{D}^{s}(E)$ is a $1$-dimensional space (because $\mathcal{B}_{E}$ is a $\mathrm{GL}(E)$-torsor), oriented by $\mathscr{D}^{s}_{>0}(E)$.
Example 5.2.
Any volume form $\omega\in\Lambda^{d}E^{*}$ produces an $s$-density defined by $\lvert\omega\rvert^{s}$, which is a positively oriented basis of $\mathscr{D}^{s}(E)$.
Example 5.3.
A non-degenerate symmetric bilinear form $\varphi:E\times E\to\mathbb{R}$ produces a positive $s$-density $\mathrm{vol}_{\varphi}^{s}\in\mathscr{D}^{s}_{>0}(E)$ defined by $\mathrm{vol}_{\varphi}^{s}(\mathbf{e})=|\mathrm{det}\,[\varphi]_{\mathbf{e}}|^{s/2}$, where the $(d\times d)$-Gram matrix $[\varphi]_{\mathbf{e}}$ has coefficients $\varphi(e_{i},e_{j})$ for $\mathbf{e}=(e_{1},\dots,e_{d})$.
Next, consider a grading $E_{\bullet}$ of $E$, that is a collection $E_{\bullet}=(E_{i})_{i\in I}\in\prod_{i\in I}\mathcal{G}_{d_{i}}(E)$ of vector subspaces of dimensions $d_{i}>0$ such that
$E=\bigoplus_{i\in I}E_{i}~{}.$
Denote by $\mathrm{GL}(E_{\bullet})<\mathrm{GL}(E)$ the subgroup of grading-preserving transformations (i.e. preserving each subspace $E_{i}$). We now adapt the notion of density to account for the grading.
Definition 5.4.
For a tuple $\mathbf{s}=(s_{i})_{i\in I}\in\mathbb{R}^{I}$, which we will occasionally call a weight vector, an $\mathbf{s}$-multidensity on $E$ with respect to the grading $E_{\bullet}$ is a map
$\nu:\prod_{i\in I}\mathcal{B}_{E_{i}}\rightarrow\mathbb{R}$
such that for every $g\in\mathrm{GL}(E_{\bullet}):$
$\nu\big{(}g\cdot(\mathbf{e}_{i})_{i\in I}\big{)}=\prod_{i\in I}\lvert\mathrm{det}(g_{i})\rvert^{s_{i}}\nu\big{(}(\mathbf{e}_{i})_{i\in I}\big{)}$
for every ordered basis $\mathbf{e}_{i}\in\mathcal{B}_{E_{i}}$ and every $i\in I.$ Above, the linear map $g_{i}\in\mathrm{GL}(E_{i})$ is the restriction of $g\in\mathrm{GL}(E_{\bullet})$ to $E_{i}$, and $g\cdot(\mathbf{e}_{i})_{i\in I}=(g_{i}\cdot\mathbf{e}_{i})_{i\in I}$. We write $\mathscr{D}^{\mathbf{s}}(E_{\bullet})$ for the real vector space of $\mathbf{s}$-multidensities on $E_{\bullet}$, and $\mathscr{D}^{\mathbf{s}}_{>0}(E_{\bullet})$ for the subset of positive $\mathbf{s}$-multidensities.
Again appealing to the free and transitive action of $\mathrm{GL}(E_{\bullet})$ on $\prod_{i\in I}\mathcal{B}_{E_{i}}$, the space $\mathscr{D}^{\mathbf{s}}(E_{\bullet})$ is $1$-dimensional and oriented.
Example 5.5.
A collection $(\nu_{i})_{i\in I}$ of $s_{i}$-densities on each $E_{i}$ yields an $\mathbf{s}$-multidensity $\prod_{i\in I}\nu_{i}\in\mathscr{D}^{\mathbf{s}}(E_{\bullet})$ defined by
$\left(\prod_{i\in I}\nu_{i}\right)\Big{(}(\mathbf{e}_{i})_{i\in I}\Big{)}:=\prod_{i\in I}\nu_{i}(\mathbf{e}_{i})~{}.$ (5.1)
defining an isomorphism $\bigotimes_{i\in I}\mathscr{D}^{s_{i}}(E_{i})\rightarrow\mathscr{D}^{\mathbf{s}}(E_{\bullet})$ preserving positivity.
Example 5.6.
This product operation can be combined with any concrete construction of $s_{i}$-densities on each $E_{i}$, such as starting with a collection of volume forms $\omega_{i}\in\Lambda^{d_{i}}E_{i}$ for each $i\in I$ or a symmetric bilinear form $\varphi:E\times E\to\mathbb{R}$ whose restriction to each subspace $E_{i}$ is non-degenerate.
Given a vector bundle $E\rightarrow M$ over a smooth manifold $M$, by a grading of $E\to M$ we mean a collection $E_{\bullet}=(E_{i})_{i\in I}$ of vector subbundles of $E$ such that $E=\bigoplus_{i\in I}E_{i}$. Given such a grading, we call the pair $(E,E_{\bullet})$ a graded vector bundle (often written just $E_{\bullet}$). The fiberwise application sending a graded vector space to the space of $\mathbf{s}$-multidensities defines the real vector line bundle $\mathscr{D}^{\mathbf{s}}(E_{\bullet})\rightarrow M$ whose sections are multidensities in the graded fibers of $E$.
Next, consider $\mathbb{F}$ a homogeneous $G$-manifold, and $(E,E_{\bullet})$ a $G$-equivariant graded vector bundle over ${\mathbb{F}}$ (i.e., $E$ and each of its subbundles $E_{i}$ in the grading are $G$-equivariant). Then $\mathscr{D}^{\mathbf{s}}(E_{\bullet})$ is a $G$-equivariant real vector line bundle over $\mathbb{F}.$ A section $\nu$ of $\mathscr{D}^{\mathbf{s}}(E_{\bullet})$ is called an $\mathbf{s}$-multidensity on $E_{\bullet}$ over $\mathbb{F}$. If $E=T\mathbb{F}$ and $\mathbf{s}=s$ is just a scalar parameter (identified with a $1$-tuple), it is convention to refer to sections of $\mathscr{D}^{s}(T\mathbb{F})$ as simply the $s$-densities on $\mathbb{F}$.
5.2. (Multi-)density bundles on flag manifolds
We now apply the constructions from Sections 5.1 and 4.2 to concrete (graded) vector bundles over flag manifolds.
5.2.1. Density bundles
Let ${\mathbb{F}}$ be a homogeneous $G$-manifold and $s\in\mathbb{R}$. We write
$\mathscr{D}^{s}{\mathbb{F}}:=\mathscr{D}^{s}T{\mathbb{F}}$ (5.2)
for the oriented line bundle of $s$-densities on ${\mathbb{F}}$. The total space of this bundle carries a $G$-action obtained from the $G$-action on $T{\mathbb{F}}$ through multilinear extension, making $\mathscr{D}^{s}{\mathbb{F}}$ a $G$-equivariant oriented vector line bundle. Section 4.2 provides the $G$-equivariant affine line bundle
$\mathscr{D}^{s}_{>0}{{\mathbb{F}}}:=\mathscr{D}^{s}_{>0}T{\mathbb{F}}$
of positive $s$-densities on ${\mathbb{F}}$. When ${\mathbb{F}}$ carries a $G$-invariant pseudo-Riemannian metric (such is the case of a transverse flag space ${\mathbb{F}}={\mathcal{F}}^{\pitchfork}$, as seen in Section 3.3), the $G$-equivariant affine line bundle $\mathscr{D}^{s}_{>0}{{\mathbb{F}}}$ has a $G$-equivariant section, so it is $G$-equivariantly trivializable.
When ${\mathbb{F}}={\mathcal{F}}$ is a non-trivial flag manifold and $s\neq 0$, then $\mathscr{D}^{s}_{>0}{{\mathcal{F}}}$ is a homogeneous affine line bundle. Given a pair of opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$, the period functions of the density bundles $\mathscr{D}^{s}_{>0}{\mathcal{F}}^{\pm}$ are related to Jacobian characters of Levi subgroups:
Corollary 5.7.
Consider a pair of opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$, suppose that $g\in G$ has a pair of transverse fixed points $(\mathbf{x}^{+},\mathbf{x}^{-})\in{\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ and write $L=\mathrm{Stab}_{G}(\mathbf{x})$ for the corresponding Levi subgroup. Then one has
$\ell_{\mathscr{D}^{s}_{>0}{\mathcal{F}}^{+}}(\mathbf{x}^{+},g)$ $=s\,\mathbf{J}_{L}(g^{-1}),$
$\ell_{\mathscr{D}^{s}_{>0}{\mathcal{F}}^{-}}(\mathbf{x}^{-},g)$ $=s\,\mathbf{J}_{L}(g).$
Proof.
Apply Lemma 3.36. ∎
Here the presence of $g^{-1}$ in the first equation reflects the sign change pointed out in Remark 3.1. From Lemma 3.39, we also get a way of computing the cocycle associated to a section of $\mathscr{D}^{s}_{>0}{\mathcal{F}}$ (Definition 4.9).
Corollary 5.8.
Consider a pair of opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$, let $\sigma$ be a section of $\mathscr{D}^{s}_{>0}{\mathcal{F}}^{-}$ for some $s\in\mathbb{R}$, $g\in G$ an ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$-proximal element with attracting fixed point ${\mathbf{x}}^{+}\in{\mathcal{F}}^{+}$ and fix a basis $(X_{1},\dots,X_{n})$ of the nilpotent radical $\mathfrak{n}$ of the Lie algebra of its stabilizer. Then for any $\mathbf{x}\in{\mathcal{F}}^{-}$ transverse to ${\mathbf{x}}^{+}$, one has
$\mathcal{C}_{\sigma}(\mathbf{x},g)+s\mathcal{J}_{{\mathcal{F}}^{+}}(g)=\log\frac{\sigma_{\mathbf{x}}\left(\overline{X_{1}}(\mathbf{x}),\dots,\overline{X_{n}}(\mathbf{x})\right)}{\sigma_{g\cdot\mathbf{x}}\left(\overline{X_{1}}(g\cdot\mathbf{x}),\dots,\overline{X_{n}}(g\cdot\mathbf{x})\right)}\,.$
Proof.
The formula $d_{\mathbf{x}}g\cdot\overline{X}(\mathbf{x})=\overline{\mathrm{Ad}(g)X}(g\cdot\mathbf{x})$ and the basis $(d_{\mathbf{x}}g\cdot\overline{X_{1}}(\mathbf{x}),\dots,d_{\mathbf{x}}g\cdot\overline{X_{n}}(\mathbf{x}))$ of $T_{g\cdot\mathbf{x}}{\mathcal{F}}^{-}$ allow us to express
$\mathcal{C}_{\sigma}(g,\mathbf{x})=\log\frac{\sigma_{\mathbf{x}}\left(\overline{X_{1}}(\mathbf{x}),\dots,\overline{X_{n}}(\mathbf{x})\right)}{\sigma_{g\cdot\mathbf{x}}\left(d_{\mathbf{x}}g\cdot\overline{X_{1}}(\mathbf{x}),\dots,d_{\mathbf{x}}g\cdot\overline{X_{n}}(\mathbf{x})\right)}\,,$
while Lemma 3.39 and the definition of a density yield
$s\mathcal{J}_{{\mathcal{F}}^{+}}(g)$ $=s\log\left|\mathrm{det}\,\big{(}\mathrm{Ad}(g)|_{\mathfrak{n}}\big{)}\right|$
$=\log\frac{\sigma_{g\cdot\mathbf{x}}\left(\overline{\mathrm{Ad}(g)X_{1}}(g\cdot\mathbf{x}),\dots,\overline{\mathrm{Ad}(g)X_{n}}(g\cdot\mathbf{x})\right)}{\sigma_{g\cdot\mathbf{x}}\left(\overline{X_{1}}(g\cdot\mathbf{x}),\dots,\overline{X_{n}}(g\cdot\mathbf{x})\right)}$
$=\log\frac{\sigma_{g\cdot\mathbf{x}}\left(d_{\mathbf{x}}g\cdot\overline{X_{1}}(\mathbf{x}),\dots,d_{\mathbf{x}}g\cdot\overline{X_{n}}(\mathbf{x})\right)}{\sigma_{g\cdot\mathbf{x}}\left(\overline{X_{1}}(g\cdot\mathbf{x}),\dots,\overline{X_{n}}(g\cdot\mathbf{x})\right)}\,.$
5.2.2. Multidensity bundles
The graded vector bundles required to introduce multidensity bundles on flag manifolds will be defined using the hierarchy introduced in Section 3.6. Consider a flag manifold ${\mathcal{F}}$ and the associated subset $\Theta\subset\Delta$.
Recall from (3.15) and Lemma 3.16 that for each $\alpha\in\Theta$ there is a unique $G$-equivariant projection $\Pi_{\alpha}^{\mathcal{F}}:{\mathcal{F}}\to{\mathcal{F}}_{\alpha}$. We can use these projections to define the vector bundle
$E^{\mathcal{F}}:=\bigoplus_{\alpha\in\Theta}\big{(}\Pi^{\mathcal{F}}_{\alpha}\big{)}^{*}\big{(}T{\mathcal{F}}_{\alpha}\big{)}$
over ${\mathcal{F}}$ with the obvious grading $E^{\mathcal{F}}_{\bullet}=\big{(}E^{\mathcal{F}}_{\alpha}\big{)}_{\alpha\in\Theta}$ defined over $\mathbf{x}\in{\mathcal{F}}$ by
$\big{(}E^{\mathcal{F}}_{\alpha}\big{)}_{\mathbf{x}}:=\left\{(v_{\alpha^{\prime}})_{\alpha^{\prime}\in\Theta}\in\bigoplus_{\alpha^{\prime}\in\Theta}T_{\Pi^{\mathcal{F}}_{\alpha^{\prime}}(\mathbf{x})}{\mathcal{F}}_{\alpha^{\prime}}\,\middle|\,v_{\alpha^{\prime}}=0~{}\forall\alpha^{\prime}\neq\alpha\right\}\simeq T_{\Pi^{\mathcal{F}}_{\alpha}(\mathbf{x})}{\mathcal{F}}_{\alpha}.$
The following diagram depicts the maps involved in the construction of the grading $E^{\mathcal{F}}_{\bullet}$.
[…]
Definition 5.9.
Let ${\mathcal{F}}$ be a flag manifold corresponding to a subset $\Theta\subset\Delta$, and $\mathbf{s}=(s_{\alpha})_{\alpha\in\Theta}\in\mathbb{R}^{\Theta}$. As in Section 5.1, we define the $\mathbf{s}$-multidensity bundle over ${\mathcal{F}}$ to be
$\mathscr{D}^{\mathbf{s}}{\mathcal{F}}_{\bullet}:=\mathscr{D}^{\mathbf{s}}E^{\mathcal{F}}_{\bullet}$
and corresponding positive $\mathbf{s}$-multidensity bundle
$\mathscr{D}^{\mathbf{s}}_{>0}{\mathcal{F}}_{\bullet}:=\mathscr{D}^{\mathbf{s}}_{>0}E^{\mathcal{F}}_{\bullet}~{}.$
The positive $\mathbf{s}$-multidensity bundle $\mathscr{D}^{\mathbf{s}}_{>0}{\mathcal{F}}_{\bullet}$ is a $G$-equivariant affine line bundle over ${\mathcal{F}}$.
Lemma 5.10.
[…]
$\ell_{\mathscr{D}^{\mathbf{s}}_{>0}{{\mathcal{F}}^{+}_{\bullet}}}(\mathbf{x}^{+},g)$ $=\sum_{\alpha\in\Theta}s_{\alpha}\mathbf{J}_{L_{\alpha}}(g^{-1}),$
$\ell_{\mathscr{D}^{\mathbf{s}}_{>0}{{\mathcal{F}}^{-}_{\bullet}}}(\mathbf{x}^{-},g)$ $=\sum_{\alpha\in\Theta}s_{\alpha}\mathbf{J}_{L_{\alpha}}(g).$
Proof.
Write $\mathbf{x}^{+}_{\alpha}=\Pi^{{\mathcal{F}}^{+}}_{\alpha}(\mathbf{x}^{+})$ and $\mathbf{x}^{-}_{\alpha}=\Pi^{{\mathcal{F}}^{-}}_{\alpha}(\mathbf{x}^{-})$, so that we have
$\ell_{\mathscr{D}^{\mathbf{s}}_{>0}{{\mathcal{F}}^{+}_{\bullet}}}(\mathbf{x}^{+},g)$ $=\sum_{\alpha\in\Theta}\ell_{\mathscr{D}^{s_{\alpha}}_{>0}{{\mathcal{F}}^{+}_{\alpha}}}(\mathbf{x}^{+}_{\alpha},g),$
$\ell_{\mathscr{D}^{\mathbf{s}}_{>0}{{\mathcal{F}}^{-}_{\bullet}}}(\mathbf{x}^{-},g)$ $=\sum_{\alpha\in\Theta}\ell_{\mathscr{D}^{s_{\alpha}}_{>0}{{\mathcal{F}}^{+}_{\alpha}}}(\mathbf{x}^{+}_{\alpha},g).$
By Lemma 3.40, each pair $(\mathbf{x}^{+}_{\alpha},\mathbf{x}^{-}_{\alpha})$ is transverse, so the claim follows from Corollary 5.7. Note that the change of signs present in Corollary 5.7, which results from Remark 3.1, is reflected here in the appearance of $g^{-1}$ in the $+$ case and $g$ in the -- case. ∎
Remark 5.1.
It is important to note that, while the ordinary density bundles were defined in (5.2) using the tangent bundle as the underlying vector bundle, the multidensity bundle $\mathscr{D}^{\mathbf{s}}{\mathcal{F}}_{\bullet}$ is not constructed from the tangent bundle of ${\mathcal{F}}$ (unless ${\mathcal{F}}$ is a simple flag manifold).
6. Flow spaces
With the preparations from Sections 4 and 5 at hand, it is now time to introduce the desired flow spaces. The idea is to create a multidensity analog of the hyperbola branch $\{(x^{+},x^{-})\in\mathbb{R}_{>0}^{2}\,|\,x^{+}\cdot x^{-}=1\}$. To this end, we need to pullback both $\mathbf{s}$-multidensity bundles $\mathscr{D}^{\mathbf{s}}{\mathcal{F}}^{+}_{\bullet}$ and $\mathscr{D}^{\mathbf{s}}{\mathcal{F}}^{-}_{\bullet}$ to the transverse flag space ${\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ where a pairing can be defined thanks to the pseudo-Riemannian structure.
6.1. Pairings of multidensities on a transverse flag space
Consider a pair of opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ with corresponding subsets $\Theta\subset\Delta$ and $\iota_{o}(\Theta)$ respectively, and a weight vector $\mathbf{s}\in\mathbb{R}^{\Theta}$. By using the projections
${{\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}}$${{\mathcal{F}}^{+}}$${{\mathcal{F}}^{-}}$$\scriptstyle{p_{{\mathcal{F}}^{+}}}$$\scriptstyle{p_{{\mathcal{F}}^{-}}}$
we can consider the real rank-two vector bundle
$\mathscr{D}^{\mathbf{s}}\mathcal{F}^{+}_{\bullet}\boxplus\mathscr{D}^{\mathbf{s}}\mathcal{F}^{-}_{\bullet}:=p_{{\mathcal{F}}^{+}}^{*}\big{(}\mathscr{D}^{\mathbf{s}}\mathcal{F}^{+}_{\bullet}\big{)}\oplus p_{{\mathcal{F}}^{-}}^{*}\big{(}\mathscr{D}^{\mathbf{s}}\mathcal{F}^{-}_{\bullet}\big{)}$ (6.1)
[…]
$\mu^{+}:\prod_{\alpha\in\Theta}\mathcal{B}_{T_{\mathbf{x}^{+}_{\alpha}}{\mathcal{F}}^{\alpha}}\to\mathbb{R}~{},\quad\mu^{-}:\prod_{\alpha\in\Theta}\mathcal{B}_{T_{\mathbf{x}^{-}_{\alpha}}{\mathcal{F}}_{\iota_{o}(\alpha)}}\to\mathbb{R}~{}$
obeying a given transformation rule under grading preserving linear transformations, where $\mathcal{B}_{E}$ stands for the set of ordered bases of a real vector space $E$, and we use the notations $\mathbf{x}^{+}_{\alpha}=\Pi^{{\mathcal{F}}^{+}}_{\alpha}(\mathbf{x}^{+})\in{\mathcal{F}}_{\alpha}$, $\mathbf{x}^{-}_{\alpha}=\Pi^{{\mathcal{F}}^{-}}_{\iota_{o}(\alpha)}(\mathbf{x}^{-})\in{\mathcal{F}}_{\iota_{o}(\alpha)}$ for the projections to simple flag manifolds.
An ordered basis $\mathbf{e}_{\alpha}^{+}=(e_{1}^{+},\dots,e_{n}^{+})$ of the tangent space $T_{\mathbf{x}^{+}_{\alpha}}{\mathcal{F}}_{\alpha}$ together with an ordered basis $\mathbf{e}_{\alpha}^{-}=(e_{1}^{-},\dots,e_{n}^{-})$ of the tangent space $T_{\mathbf{x}^{-}_{\alpha}}{\mathcal{F}}_{\iota_{o}(\alpha)}$ can be concatenated in order to form an ordered basis $(\mathbf{e}^{+}_{\alpha},\mathbf{e}^{-}_{\alpha}):=\big{(}(e_{1}^{+},0),\dots,(e^{+}_{n},0),(0,e^{-}_{1}),\dots,(0,e_{n}^{-})\big{)}$ of $T_{\mathbf{x}^{+}_{\alpha}}{\mathcal{F}}_{\alpha}\oplus T_{\mathbf{x}^{-}_{\alpha}}{\mathcal{F}}_{\iota_{o}(\alpha)}=T_{(\mathbf{x}^{+}_{\alpha},\mathbf{x}^{-}_{\alpha})}{\mathcal{F}}_{\alpha}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}_{\iota_{o}(\alpha)}$. So by using the pseudo-Riemannian metric on the transverse flag manifold ${\mathcal{F}}_{\alpha}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}_{\iota_{o}(\alpha)}$ defined in Section 3.3 and the construction of an associated $s_{\alpha}$-density $\mathrm{vol}^{s_{\alpha}}_{{\mathcal{F}}_{\alpha}^{\pitchfork}}$ from Example 5.3 we can define the number
$\mathrm{vol}^{s_{\alpha}}_{{\mathcal{F}}_{\alpha}^{\pitchfork}}(\mathbf{e}^{+}_{\alpha},\mathbf{e}^{-}_{\alpha})>0~{}.$
We now see from the transformation rule in the definition of a multidensity that the ratio
$\mu^{+}\cdot\mu^{-}:=\frac{\mu^{+}\big{(}(\mathbf{e}^{+}_{\alpha})_{\alpha\in\Theta}\big{)}\mu^{-}\big{(}(\mathbf{e}^{-}_{\alpha})_{\alpha\in\Theta}\big{)}}{\prod_{\alpha\in\Theta}\mathrm{vol}^{s_{\alpha}}_{{\mathcal{F}}_{\alpha}^{\pitchfork}}(\mathbf{e}^{+}_{\alpha},\mathbf{e}^{-}_{\alpha})}$ (6.2)
does not depend on the choice of a family of ordered bases $(\mathbf{e}^{+}_{\alpha})_{\alpha\in\Theta}\in\prod_{\alpha\in\Theta}\mathcal{B}_{T_{\mathbf{x}^{+}_{\alpha}}{\mathcal{F}}^{\alpha}}$ and $(\mathbf{e}^{-}_{\alpha})_{\alpha\in\Theta}\in\prod_{\alpha\in\Theta}\mathcal{B}_{T_{\mathbf{x}^{-}_{\alpha}}{\mathcal{F}}_{\iota_{o}(\alpha)}}$, thus defining an analytic $G$-invariant function
$\begin{split}\mathscr{D}^{\mathbf{s}}\mathcal{F}^{+}_{\bullet}\boxplus\mathscr{D}^{\mathbf{s}}\mathcal{F}^{-}_{\bullet}&\rightarrow\mathbb{R}\\ (\mu^{+},\mu^{-})&\mapsto\mu^{+}\cdot\mu^{-}\end{split}$
that satisfies the homogeneity relation
$\big{(}t^{+}\mu^{+}\big{)}\cdot\big{(}t^{-}\mu^{-}\big{)}=t^{+}t^{-}\mu^{+}\cdot\mu^{-}\qquad\forall\;(t^{+},t^{-})\in\mathbb{R}^{2}~{},$ (6.3)
and that is positive on the subset $\mathscr{D}^{\mathbf{s}}_{>0}\mathcal{F}^{+}_{\bullet}\boxplus\mathscr{D}^{\mathbf{s}}_{>0}\mathcal{F}^{-}_{\bullet}\subset\mathscr{D}^{\mathbf{s}}\mathcal{F}^{+}_{\bullet}\boxplus\mathscr{D}^{\mathbf{s}}\mathcal{F}^{-}_{\bullet}$.
6.2. The multidensity flow
Definition 6.1.
Consider a pair of opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ with corresponding subsets $\Theta\subset\Delta$ and $\iota_{o}(\Theta)$ respectively, and a weight vector $\mathbf{s}\in\mathbb{R}^{\Theta}$.
The flow space $\mathbb{L}^{\mathbf{s}}_{\Theta}$ of ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ with parameter $\mathbf{s}$ is the $G$-equivariant affine line bundle over ${{\mathcal{F}}^{\pitchfork}}$ given by
${\mathbb{L}}^{\mathbf{s}}_{\Theta}:=\left\{\left(\mu^{+},\mu^{-}\right)\in\mathscr{D}^{\mathbf{s}}_{>0}\mathcal{F}^{+}_{\bullet}\boxplus\mathscr{D}^{\mathbf{s}}_{>0}\mathcal{F}^{-}_{\bullet}\,\middle|\,\mu^{+}\cdot\mu^{-}=1\right\},$
on which we let $t\in\mathbb{R}$ act by the flow $\phi^{t}$ defined as
$\phi^{t}\left(\mu^{+},\mu^{-}\right):=\left(e^{-t}\mu^{+},e^{t}\mu^{-}\right).$
Note that this flow is well-defined because of the homogeneity of the pairing (6.3).
More important than the flow spaces themselves are their isomorphism classes, where, as before, “isomorphism” means “isomorphism of $G$-equivariant affine line bundles”, (Definition 4.1). Projecting onto the first and second factors of the flow space ${\mathbb{L}}^{\mathbf{s}}_{\Theta}\subset\mathscr{D}^{\mathbf{s}}_{>0}\mathcal{F}^{+}_{\bullet}\boxplus\mathscr{D}^{\mathbf{s}}_{>0}\mathcal{F}^{-}_{\bullet}$ reveals that ${\mathbb{L}}^{\mathbf{s}}_{\Theta}$ is nothing but a particularly useful (or “symmetric”) representative of the isomorphism class
$[{\mathbb{L}}^{\mathbf{s}}_{\Theta}]=\big{[}p_{{\mathcal{F}}^{-}}^{\ast}\big{(}\mathscr{D}^{\mathbf{s}}_{>0}\mathcal{F}^{-}_{\bullet}\big{)}\big{]}=\big{[}p_{{\mathcal{F}}^{+}}^{\ast}\big{(}\overline{\mathscr{D}^{\mathbf{s}}_{>0}\mathcal{F}^{+}_{\bullet}}\big{)}\big{]}.$ (6.4)
From (6.4) and Lemma 5.10 we learn that the period function of ${\mathbb{L}}^{\mathbf{s}}_{\Theta}$ is given by
$\ell_{{\mathbb{L}}^{\mathbf{s}}_{\Theta}}=\sum_{\alpha\in\Theta}s_{\alpha}\mathbf{J}_{{{\mathcal{F}}^{\pitchfork}}}^{\alpha}\,.$ (6.5)
Corollary 6.2.
Every $G$-equivariant affine line bundle ${\mathbb{L}}\to{{\mathcal{F}}^{\pitchfork}}$ is isomorphic to ${\mathbb{L}}^{\mathbf{s}}_{{{\mathcal{F}}^{\pitchfork}}}$ for a unique $\mathbf{s}\in\mathbb{R}^{\Theta}$.
Proof.
It follows from (6.5), Corollary 4.7 and Proposition 4.8. ∎
7. Anosov subgroups
The proximality class functions $\mathcal{P}_{\alpha}:G\to\mathbb{R}_{\geq 0}$ introduced in Definition 3.25 for the simple flag manifolds ${\mathcal{F}}_{\alpha}$, $\alpha\in\Delta$ provide an intrinsic way of stating Kassel-Potrie’s definition of Anosov subgroups from [KP22], in which we can flexibly switch between the different interpretations of the set $\Delta$ explained in Section 3.6.
Definition 7.1.
Given a subset $\Theta\subset\Delta$, a word-hyperbolic discrete subgroup $\Gamma<G$ is called $\Theta$-Anosov if there are constants $C,c>0$ such that
$\mathcal{P}_{\alpha}(\gamma)\geq c|\gamma|_{\infty}-C$
holds for all $\alpha\in\Theta$ and $\gamma\in\Gamma$.
Here $|\cdot|_{\infty}$ is the stable length with respect to some finite generating set of $\Gamma$, the choice of which does not matter in the definition (see Section 2.2 for more details and references).
Let ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ be a pair of opposite flag manifolds associated to subsets $\Theta,\iota_{o}(\Theta)\subset\Delta$ respectively, ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ their transverse flag space, and let $\Gamma<G$ be a $\Theta$-Anosov subgroup. We denote by $\partial_{\infty}\Gamma$ its Gromov-boundary and write
$\partial_{\infty}\Gamma^{(2)}:=\{(x,y)\in\partial_{\infty}\Gamma\times\partial_{\infty}\Gamma\,|\,x\neq y\}.$
The following consequences of the $\Theta$-Anosov property of $\Gamma$ are proved in [KP22, GGKW17, KLP17, GW12]:
  1. (A1)
    There are continuous $\Gamma$-equivariant maps $\xi^{\pm}:\partial_{\infty}\Gamma\to\mathcal{F}^{\pm}$, called boundary maps.
  2. (A2)
    The boundary maps are transverse in the sense that
    $\forall(x,y)\in\partial_{\infty}\Gamma^{(2)}\quad(\xi^{+}(x),\xi^{-}(y))\in{{\mathcal{F}}^{\pitchfork}}.$
  3. (A3)
    The boundary maps preserve the convergence properties of the $\Gamma$-action on $\partial_{\infty}\Gamma$: For any unbounded sequence $\gamma_{n}\in\Gamma$ with boundary limit points $\gamma_{+}=\lim_{n\to+\infty}\gamma_{n}\in\partial_{\infty}\Gamma$ and $\gamma_{-}=\lim_{n\to+\infty}\gamma_{n}^{-1}\in\partial_{\infty}\Gamma$, there is a subsequence $\gamma_{n_{k}}$ for which the $\Gamma$-actions on ${\mathcal{F}}^{+}$ and ${\mathcal{F}}^{-}$ obey the following dynamics as $k\to+\infty$:
    1. (a)
      $\gamma_{n_{k}}\cdot\mathbf{x}^{+}\rightarrow\xi^{+}(\gamma_{+})$ for all $\mathbf{x}^{+}\in{\mathcal{F}}^{+}$ with $\mathbf{x}^{+}\pitchfork\xi^{-}(\gamma_{-})$;
    2. (b)
      $\gamma_{n_{k}}^{-1}\cdot\mathbf{x}^{+}\rightarrow\xi^{+}(\gamma_{-})$ for all $\mathbf{x}^{+}\in{\mathcal{F}}^{+}$ with $\mathbf{x}^{+}\pitchfork\xi^{-}(\gamma_{+})$;
    3. (c)
      $\gamma_{n_{k}}\cdot\mathbf{x}^{-}\rightarrow\xi^{-}(\gamma_{+})$ for all $\mathbf{x}^{-}\in{\mathcal{F}}^{-}$ with $\xi^{+}(\gamma_{-})\pitchfork\mathbf{x}^{-}$;
    4. (d)
      $\gamma_{n_{k}}^{-1}\cdot\mathbf{x}^{-}\rightarrow\xi^{-}(\gamma_{-})$ for all $\mathbf{x}^{-}\in{\mathcal{F}}^{-}$ with $\xi^{+}(\gamma_{+})\pitchfork\mathbf{x}^{-}$;
    and all these convergences are uniform on compact subsets.
We combine the boundary maps $\xi^{+},\xi^{-}$ to the map
$\xi^{\pitchfork}:\left\{\begin{array}[]{ccc}\partial_{\infty}\Gamma^{(2)}&\to&{{\mathcal{F}}^{\pitchfork}}\\ (z^{+},z^{-})&\mapsto&(\xi^{+}(z^{+}),\xi^{-}(z^{-})).\end{array}\right.$ (7.1)
The images of $\xi^{\pm}$ and $\xi^{\pitchfork}$ are the limit sets in ${\mathcal{F}}^{\pm}$ and ${{\mathcal{F}}^{\pitchfork}}$, respectively, denoted by
$\Lambda^{\pm}_{\Gamma}:=\xi^{\pm}(\partial_{\infty}\Gamma)\subset{\mathcal{F}}^{\pm},\qquad\Lambda^{\pitchfork}_{\Gamma}:=\xi^{\pitchfork}(\partial_{\infty}\Gamma^{(2)})\subset{{\mathcal{F}}^{\pitchfork}}.$
For a given word-hyperbolic subgroup $\Gamma<G$ one has (see [GW12, Lemma 3.18])
[…]
If $\Theta=\emptyset$, then only finite subgroups of $G$ are $\Theta$-Anosov and the claims are trivially true. Thus, in the following we assume that $\Theta\neq\emptyset$.
7.1. Invariant subsets of the flow space
As before, let ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ be a pair of opposite flag manifolds associated to subsets $\Theta,\iota_{o}(\Theta)\subset\Delta$ respectively, with transverse flag space ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$, and let $\Gamma<G$ be $\Theta$-Anosov. Consider a flow space ${\mathbb{L}}^{\mathbf{s}}_{\Theta}$ as defined in Section 6. Then it turns out that $\Gamma$ possesses a non-empty flow-invariant domain of proper discontinuity in ${\mathbb{L}}^{\mathbf{s}}_{\Theta}$ for suitable parameters $\mathbf{s}\in\mathbb{R}^{\Theta}$. The heart of the construction is the following
Definition 7.2.
Consider the following subset of ${{\mathcal{F}}^{\pitchfork}}$:
$\Omega_{\Gamma}:=\left\{(\mathbf{x}^{+},\mathbf{x}^{-})\in{{\mathcal{F}}^{\pitchfork}}\,\middle|\,\forall\,z\in\partial_{\infty}\Gamma:\mathbf{x}^{+}\pitchfork\xi^{-}(z)\textrm{ or }\xi^{+}(z)\pitchfork\mathbf{x}^{-}\right\},$
as well as the subsets of $\mathbb{L}^{\mathbf{s}}_{\Theta}$
~$\widetilde{\mathcal{M}}_{\Gamma,\Theta}^{\mathbf{s}}$ $:=p^{-1}(\Omega_{\Gamma})=\left\{\mu\in\mathbb{L}^{\mathbf{s}}_{\Theta}\,\middle|\,p(\mu)\in\Omega_{\Gamma}\right\},$
$\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$ $:=p^{-1}(\Lambda_{\Gamma}^{\pitchfork})=\left\{\mu\in\mathbb{L}^{\mathbf{s}}_{\Theta}\,\middle|\,p(\mu)\in\Lambda^{\pitchfork}_{\Gamma}\right\},$
where $p:{\mathbb{L}}_{\Theta}^{\mathbf{s}}\to{{\mathcal{F}}^{\pitchfork}}$ is the bundle projection.
Note that ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ (resp. $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$) is the restriction of the bundle $\mathbb{L}^{\mathbf{s}}_{\Theta}$ to the subset $\Omega_{\Gamma}$ (resp. $\Lambda^{\pitchfork}_{\Gamma}$) of its base space ${{\mathcal{F}}^{\pitchfork}}$. In particular, since the flow $\phi^{t}$ preserves the fibers of $\mathbb{L}^{\mathbf{s}}_{\Theta}$, we see that ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ and $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$ are $\phi^{t}$-invariant subsets of $\mathbb{L}^{\mathbf{s}}_{\Theta}$. Further elementary properties can be shown exactly as in the projective Anosov case, see [DMS24, Lemmas 3.4, 3.7]:
Lemma 7.3.
The sets $\Omega_{\Gamma}\subset{{\mathcal{F}}^{\pitchfork}}$ and ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}\subset\mathbb{L}^{\mathbf{s}}_{\Theta}$ are open and $\Gamma$-invariant. The set $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}\subset\mathbb{L}^{\mathbf{s}}_{\Theta}$ is closed, $\Gamma$-invariant, and contained in ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$.
In particular, since $\Lambda^{\pitchfork}_{\Gamma}$, and consequently $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$, is non-empty unless $\Gamma$ is finite, in which case $\Omega_{\Gamma}$ is all of ${{\mathcal{F}}^{\pitchfork}}$, we conclude that the sets $\Omega_{\Gamma}$ and ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ are always non-empty.
7.2. Admissible parameters
The dynamics of the action of $\Gamma$ on $\mathbb{L}_{\Theta}^{\mathbf{s}}$ depends on the parameter $\mathbf{s}\in\Theta$, in order to understand which parameters give a sufficiently well behaved action we need to associate them to linear forms on the Benoist limit cone. Consider a restricted root decomposition of ${\bf\mathfrak{g}}$, and recall the Jordan projection $\lambda:G\to{\bf\mathfrak{a}}^{+}$ of $G$ defined in Section 3.7.
Definition 7.4.
The Benoist limit cone $\mathcal{L}(\Gamma)\subset{\bf\mathfrak{a}}^{+}$ is the smallest closed cone in ${\bf\mathfrak{a}}^{+}$ containing all Jordan projections $\lambda(\gamma)$, where $\gamma\in\Gamma$.
For $\mathbf{s}\in\mathbb{R}^{\Theta}$, consider the element
$w^{\mathbf{s}}_{\Theta}:=\sum_{\alpha\in\Theta}s_{\alpha}m_{\alpha}w_{\alpha}\in{\bf\mathfrak{a}}^{\ast},$ (7.3)
where the $w_{\alpha}$ are the fundamental weights defined in (3.17) and the positive integers $m_{\alpha}\in\mathbb{N}$ were defined in (3.21).
Definition 7.5.
An element $\mathbf{s}\in\mathbb{R}^{\Theta}$ is $(\Gamma,\Theta)$-admissible if $w^{\mathbf{s}}_{\Theta}(X)>0\;\forall\,X\in\mathcal{L}(\Gamma)\setminus\{0\}$.
Even though this definition involves a root space decomposition, we can also consider the associated class function
$\mathcal{J}_{\Theta}^{\mathbf{s}}:=w^{\mathbf{s}}_{\Theta}\circ\lambda=\sum_{\alpha\in\Theta}s_{\alpha}\mathcal{J}_{{\mathcal{F}}_{\alpha}},$ (7.4)
with $\mathcal{J}_{{\mathcal{F}}_{\alpha}}$ the Jacobian class function (Definition 3.27) of the simple flag manifold ${\mathcal{F}}_{\alpha}$.
Remark 7.1.
Let $\iota:\mathbb{R}^{\Theta}\to\mathbb{R}^{\Theta}$ be the involution defined by $\iota(\mathbf{s})=(s_{\iota_{o}(\alpha)})_{\alpha\in\Theta}$. Then it is clear from the definitions that for every $\mathbf{s}\in\mathbb{R}^{\Theta}$ we have:
[…]
Remark 7.2.
Admissible parameters can be interpreted as linear forms in the interior of some cone. Given a restricted root space decomposition, we consider ${\bf\mathfrak{a}}_{\Theta}=\bigcap_{\alpha\in\Delta\setminus\Theta}\ker\alpha$ and the orthogonal projection $p_{\Theta}:{\bf\mathfrak{a}}\to{\bf\mathfrak{a}}_{\Theta}$ with respect to the Killing form, and define the $\Theta$-limit cone $\mathcal{L}_{\Theta}(\Gamma)=p_{\Theta}\big{(}\mathcal{L}(\Gamma)\big{)}\subset{\bf\mathfrak{a}}_{\Theta}$ as well as its dual cone
$\mathcal{L}_{\Theta}(\Gamma)^{\ast}:=\left\{\varphi\in{\bf\mathfrak{a}}_{\Theta}^{*}\,\middle|\,\varphi(X)\geq 0~{}\forall X\in\mathcal{L}_{\Theta}(\Gamma)\right\}\,.$
Then $\mathbf{s}\in\mathbb{R}^{\Theta}$ is admissible if and only if the restriction $w_{\Theta}^{\mathbf{s}}|_{{\bf\mathfrak{a}}_{\Theta}}$ lies in the interior of $\mathcal{L}_{\Theta}(\Gamma)^{\ast}$.
Remark 7.3.
Note that the set of $(\Gamma,\Theta)$-admissible parameters $\mathbf{s}\in\mathbb{R}^{\Theta}$ is always non-empty: it contains $\mathbb{R}_{>0}^{\Theta}$.
7.3. Refraction flows and cocompactness
Recall from Definition 4.9 that fixing a real analytic section $\sigma:{\mathcal{F}}^{\pitchfork}\rightarrow\mathbb{L}_{\Theta}^{\mathbf{s}}$ induces a real analytic cocycle
$\mathcal{C}_{\sigma}:{\mathcal{F}}^{\pitchfork}\times G\rightarrow\mathbb{R}\,.$
It can be pre-composed with the transverse limit map $\xi^{\pitchfork}:\partial_{\infty}\Gamma^{(2)}\rightarrow{\mathcal{F}}^{\pitchfork}$ of the $\Theta$-Anosov subgroup $\Gamma<G$, resulting in the Hölder cocycle
$\mathcal{C}_{\Gamma,\sigma}:\left\{\begin{array}[]{ccc}\partial_{\infty}\Gamma^{(2)}\times\Gamma&\to&\mathbb{R}\\ (\mathbf{x},\gamma)&\mapsto&\mathcal{C}_{\sigma}(\xi^{\pitchfork}(\mathbf{x}),\gamma)\end{array}\right.\,,$
which induces a $\Gamma$-action on $\partial_{\infty}\Gamma^{(2)}\times\mathbb{R}$ via the formula
$\gamma\cdot\big{(}\mathbf{x},s\big{)}:=\big{(}\gamma\cdot\mathbf{x},s-\mathcal{C}_{\Gamma,\sigma}(\mathbf{x},\gamma)\big{)}\,.$ (7.6)
Sambarino proved [Sam24, Corollary 5.3.3] that if the parameter $\mathbf{s}\in\mathbb{R}^{\Theta}$ is $(\Gamma,\Theta)$-admissible, then this action is properly discontinuous and cocompact.
Definition 7.6.
The refraction flow space $\mathfrak{R}_{\sigma}$ associated to the $(\Gamma,\Theta)$-admissible parameter $\mathbf{s}\in\mathbb{R}^{\Theta}$ and the section $\sigma:{\mathcal{F}}^{\pitchfork}\rightarrow{\mathbb{L}}_{\Theta}^{\mathbf{s}}$ is the Hausdorff quotient of $\partial_{\infty}\Gamma^{(2)}\times\mathbb{R}$ by the proper $\Gamma$-action described above. The flow
$\psi_{\sigma}^{t}:\mathfrak{R}_{\sigma}\rightarrow\mathfrak{R}_{\sigma}$
induced by the affine translation action of $\mathbb{R}$ is called the refraction flow.
The name refraction flow is due to Sambarino [Sam24]. Although our use of an arbitrary section of $\sigma:{\mathcal{F}}^{\pitchfork}\rightarrow{\mathbb{L}}_{\Theta}^{\mathbf{s}}$ to define cocycles and $G$-actions may look more general than the explicit Hopf-Busemann-Iwasawa cocycles used in [Sam24], the computation of the period function (6.5) shows that [Sam24, Lemma 5.3.2], and therefore [Sam24, Corollary 5.3.3] or [Sam14, Theorem 3.2], still applies (Remark 7.2 offers a translation between our notations and those used in [Sam24]). The refraction flow does depends on the section $\sigma,$ but for any two sections $\sigma$ and $\sigma^{\prime},$ the resulting refraction flows $\mathfrak{R}_{\sigma}$ and $\mathfrak{R}_{\sigma^{\prime}}$ have the same periods, and are therefore homeomorphically conjugate to each other [Sam14, Rem. 3.1].
Proposition 7.7.
If $\Gamma<G$ is $\Theta$-Anosov and $\mathbf{s}\in\mathbb{R}^{\Theta}$ is $(\Gamma,\Theta)$-admissible, then the action of $\Gamma$ on $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$ is properly discontinuous and cocompact.
Proof.
Consider any section $\sigma:{\mathcal{F}}^{\pitchfork}\rightarrow\mathbb{L}_{\Theta}^{\mathbf{s}}$. The map
$\widetilde{\Phi}_{\sigma}:\left\{\begin{array}[]{ccc}\partial_{\infty}\Gamma^{(2)}\times\mathbb{R}&\to&{\mathbb{L}}_{\Theta}^{\mathbf{s}}\\ (\mathbf{x},s)&\mapsto&\phi^{s}\circ\sigma\circ\xi^{\pitchfork}(\mathbf{x})\end{array}\right.$ (7.7)
is a $\Gamma$-equivariant Hölder-continuous topological embedding with image $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$, so the result follows from [Sam24, Corollary 5.3.3]. ∎
Note that the map $\widetilde{\Phi}_{\sigma}$ of (7.7) satisfies $\widetilde{\Phi}_{\sigma}(\mathbf{x},s+t)=\phi^{t}\circ\widetilde{\Phi}_{\sigma}(\mathbf{x},s)$, so it induces a bi-Hölder homeomorphism
$\Phi_{\sigma}:\mathfrak{R}_{\sigma}\to\mathcal{K}_{\Gamma,\Theta}^{\mathbf{s}}$ (7.8)
conjugating the refraction flow with the quotient flow of $\phi^{t}$ on $\mathcal{K}_{\Gamma,\Theta}^{\mathbf{s}}:=\Gamma\backslash\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$.
7.4. Proof of proper discontinuity
With these preparations we can finally formulate the following theorem, which represents the main result of this section.
Theorem 1.
Let $\Gamma<G$ be $\Theta$-Anosov and consider a $(\Gamma,\Theta)$-admissible parameter $\mathbf{s}\in\mathbb{R}^{\Theta}$. Then $\Gamma$ acts properly discontinuously on ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ and cocompactly on $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$. If $\Gamma$ is torsion-free, then the $\Gamma$-action on ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ is free.
Remark 7.4.
Theorem 1 is a statement about ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ and $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$ as $\Gamma$-spaces, it does not involve the flow. In particular, (6.4) implies that
[…]
But, they are isomorphic as $\Gamma$-flow spaces if and only if $\iota(\mathbf{s})=\mathbf{s}$ and $\iota_{0}(\Theta)=\Theta,$ and therefore the flows are typically irreversible.
Our proof will crucially rely on the following property of Anosov subgroups, where we use the terminology from (7.4) and Definition 7.5.
Proposition 7.8.
Let $\Gamma<G$ be $\Theta$-Anosov, and $\mathbf{s}\in\mathbb{R}^{\Theta}\setminus\{0\}$ be $(\Gamma,\Theta)$-admissible. Then any sequence $(\gamma_{k})$ in $\Gamma$ possessing distinct boundary limits $\gamma_{+}=\lim_{k\to\infty}\gamma_{k}\neq\gamma_{-}=\lim_{k\to\infty}\gamma_{k}^{-1}\in\partial_{\infty}\Gamma$ satisfies
$\lim_{k\to\infty}\mathcal{J}_{\Theta}^{\mathbf{s}}(\gamma_{k})=+\infty\quad\textrm{ and }\quad\lim_{k\to\infty}\mathcal{J}_{\iota_{o}(\Theta)}^{\mathbf{s}}(\gamma_{k})=+\infty~{}.$
Proof.
By Lemma 3.29, (7.2), and (7.5) it suffices to prove the statement on $\mathcal{J}_{\Theta}^{\mathbf{s}}$. Consider a restricted root decomposition and set $\beta=w_{\Theta}^{\mathbf{s}}$. Writing
$\beta(\lambda(\gamma))=\left\|\lambda(\gamma)\right\|\beta\bigg{(}\frac{\lambda(\gamma)}{\left\|\lambda(\gamma)\right\|}\bigg{)},$
for any $\gamma\in\Gamma$ with $\lambda(\gamma)\neq 0$ and any norm on ${\bf\mathfrak{a}}$, the compactness of $\left\{X\in\mathcal{L}(\Gamma)\,\middle|\,\left\|X\right\|=1\right\}$ and the $(\Gamma,\Theta)$-admissibility of $\mathbf{s}$ mean that it suffices to show that $\left\|\lambda(\gamma_{k})\right\|\to\infty$, i.e., $\lambda(\gamma_{k})\to\infty$ in ${\bf\mathfrak{a}}$, for any sequence $(\gamma_{k})$ such as in the statement. To this end, we first apply Lemma 2.5.(2), which says that the stable length $|\gamma_{k}|_{\infty}$ converges to $\infty$. Then by Definition 7.1 we have that $\alpha(\lambda(\gamma_{k}))\to+\infty$, in particular $\lambda(\gamma_{k})\to\infty$ in ${\bf\mathfrak{a}}$. ∎
The following lemma uses cocycles associated to a section (Definition 4.9) and Jacobian class functions (Definition 3.27).
Lemma 7.9.
Suppose that $\Gamma$ is $\Theta$-Anosov. Let $\gamma_{k}\in\Gamma$ be a sequence which possesses distinct boundary limits $\gamma_{+}=\lim\gamma_{k}\neq\lim\gamma_{k}^{-1}=\gamma_{-}\in\partial_{\infty}\Gamma$. Let $s\in\mathbb{R}$ and consider a section $\sigma$ of $\mathscr{D}^{s}_{>0}{\mathcal{F}}^{-}_{\bullet}$. If $(\mathbf{x}_{k})_{k\geq 0}$ is a sequence in ${\mathcal{F}}^{-}$ converging to an element $\mathbf{x}\in\mathcal{F}^{-}$ transverse to $\xi^{+}(\gamma_{-})\in{\mathcal{F}}^{+}$, then the sequence $\mathcal{C}_{\sigma}(\mathbf{x}_{k},\gamma_{k})+s\mathcal{J}_{{\mathcal{F}}^{+}}(\gamma_{k})$ is bounded.
Proof.
Choose a basis $(X_{1,k},\dots,X_{n,k})$ of each nilpotent radical $\mathfrak{n}_{\xi^{+}(\gamma_{k}^{-})}$ converging to a basis $(X_{1},\dots,X_{n})$ of $\mathfrak{n}_{\xi^{+}(\gamma_{-})}$ as $k\to+\infty$. This is possible because we have
$(\xi^{+}(\gamma_{k}^{-}),\xi^{-}(\gamma_{k}^{+}))\stackrel{{\scriptstyle k\to\infty}}{{\longrightarrow}}(\xi^{+}(\gamma_{-}),\xi^{-}(\gamma_{+}))$ (7.10)
by Lemma 2.5 and the continuity of the boundary maps demanded in (A1), and the fact that the spaces $\mathfrak{n}_{\mathbf{x}^{+}}$ over $(\mathbf{x}^{+},\mathbf{x}^{-})\in\mathcal{F}^{\pitchfork}$ form a continuous (even analytic) vector bundle. Since $\mathbf{x}_{k}\to\mathbf{x}$ as $k\to\infty$ and transversality is an open condition, the fundamental vector fields of $X_{1,k},\dots,X_{n,k}$ at $\mathbf{x}_{k}$ form a basis of $T_{\mathbf{x}_{k}}{\mathcal{F}}^{-}$ when $k$ is large enough.
It follows from Corollary 5.8 that
$\mathcal{C}_{\sigma}(\mathbf{x}_{k},\gamma_{k})+s\mathcal{J}_{{\mathcal{F}}^{+}}(\gamma_{k})=\log\frac{\sigma_{\mathbf{x}_{k}}\left(\overline{X_{1,k}}(\mathbf{x}_{k}),\dots,\overline{X_{n,k}}(\mathbf{x}_{k})\right)}{\sigma_{\gamma_{k}\cdot\mathbf{x}_{k}}\left(\overline{X_{1,k}}(\gamma_{k}\cdot\mathbf{x}_{k}),\dots,\overline{X_{n,k}}(\gamma_{k}\cdot\mathbf{x}_{k})\right)}~{}.$ (7.11)
Note that by condition (A3)(d), we have that $\gamma_{k}\cdot\mathbf{x}_{k}\to\xi^{-}(\gamma_{+})$. Since both $\mathbf{x}$ and $\xi^{-}(\gamma_{+})$ are transverse to $\xi^{+}(\gamma_{-})$, the right hand side of (7.11) converges to
$\log\frac{\sigma_{\mathbf{x}}\left(\overline{X_{1}}(\mathbf{x}),\dots,\overline{X_{n}}(\mathbf{x})\right)}{\sigma_{\xi^{-}(\gamma_{+})}\left(\overline{X_{1}}(\xi^{-}(\gamma_{+})),\dots,\overline{X_{n}}(\xi^{-}(\gamma_{+}))\right)}\in\mathbb{R}~{},$
in particular it is bounded. ∎
Corollary 7.10.
Suppose that $\Gamma$ is $\Theta$-Anosov. Let $\gamma_{k}\in\Gamma$ be a sequence which possesses distinct boundary limits $\gamma_{+}=\lim\gamma_{k}\neq\lim\gamma_{k}^{-1}=\gamma_{-}\in\partial_{\infty}\Gamma$. Let $\mathbf{s}\in\mathbb{R}^{\Theta}$ and consider a section $\sigma$ of $\mathscr{D}^{\mathbf{s}}_{>0}{\mathcal{F}}^{-}_{\bullet}$. If $(\mathbf{x}_{k})_{k\geq 0}$ is a sequence in ${\mathcal{F}}^{-}$ converging to an element $\mathbf{x}\in\mathcal{F}^{-}$ transverse to $\xi^{+}(\gamma_{-})\in{\mathcal{F}}^{+}$, then the sequence $\mathcal{C}_{\sigma}(\mathbf{x}_{k},\gamma_{k})+\mathcal{J}^{\mathbf{s}}_{\Theta}(\gamma_{k})$ is bounded.
Proof.
Consider a section $\sigma^{\alpha}$ of $\mathscr{D}^{s_{\alpha}}_{>0}{\mathcal{F}}^{-}_{\alpha}$ for each $\alpha\in\Theta$, and let $\sigma^{\prime}=\prod_{\alpha\in\Theta}(\Pi^{{\mathcal{F}}^{-}}_{\alpha})^{\ast}\sigma^{\alpha}\in\mathscr{D}^{\mathbf{s}}_{>0}{\mathcal{F}}^{-}_{\bullet}$. According to Lemma 4.10, it is sufficient to show that the sequence $\mathcal{C}_{\sigma^{\prime}}(\mathbf{x}_{k},\gamma_{k})+\mathcal{J}^{\mathbf{s}}_{\Theta}(\gamma_{k})$ is bounded. The claim follows from Lemma 7.9, the fact that projections onto simple flag manifolds are continuous and preserve transversality, and the equality
$\mathcal{C}_{\sigma^{\prime}}(\mathbf{y},g)=\sum_{\alpha\in\Theta}\mathcal{C}_{\sigma^{\alpha}}(\Pi^{{\mathcal{F}}^{-}}_{\alpha}(\mathbf{y}),g)\quad\forall\,(\mathbf{y},g)\in{\mathcal{F}}^{-}\times G.\vspace*{-1em}$
The following Lemma is the second crucial ingredient to our proof of Theorem 1 besides Proposition 7.8.
Lemma 7.11.
Suppose that $\Gamma$ is $\Theta$-Anosov. Let $\gamma_{k}\in\Gamma$ be a sequence which possesses distinct boundary limits $\gamma_{+}=\lim\gamma_{k}\neq\lim\gamma_{k}^{-1}=\gamma_{-}\in\partial_{\infty}\Gamma$. Let $\mathbf{s}\in\mathbb{R}^{\Theta}$ and consider a converging sequence $\mathbf{x}_{k}\to\mathbf{x}\in{\mathcal{F}}^{-}$ with $\xi^{+}(\gamma_{-})\pitchfork\mathbf{x}$ as well as a converging sequence of lifts $\nu_{k}\to\nu\in{\mathscr{D}}^{\mathbf{s}}_{>0}\mathcal{F}^{-}_{\bullet}$ in the corresponding fibers.
Then, if $\lim_{k\to\infty}\mathcal{J}_{\Theta}^{\mathbf{s}}(\gamma_{k})=+\infty$, we have $\gamma_{k}\cdot\nu_{k}\to 0$ in ${\mathscr{D}}^{\mathbf{s}}\mathcal{F}^{-}_{\bullet}$ as $k\to\infty$.
Proof.
Let us fix a section $\sigma$ of ${\mathscr{D}}^{\mathbf{s}}_{>0}\mathcal{F}^{-}_{\bullet}$, and write $\nu_{k}=e^{t_{k}}\sigma_{\mathbf{x}_{k}}$ and $\nu=e^{t}\sigma_{\mathbf{x}}$, so that the convergence $\nu_{k}\to\nu\in{\mathscr{D}}^{\mathbf{s}}_{>0}\mathcal{F}^{-}_{\bullet}$ is encoded in $\mathbf{x}_{k}\to\mathbf{x}\in{\mathcal{F}}^{-}$ and $t_{k}\to t\in\mathbb{R}$. We compute
$\gamma_{k}\cdot\nu_{k}$ $=e^{t_{k}}\gamma_{k}\cdot\sigma_{\mathbf{x}_{k}}$
$=e^{t_{k}+\mathcal{C}_{\sigma}(\mathbf{x}_{k},\gamma_{k})}\sigma_{\gamma_{k}\cdot\mathbf{x}_{k}}~{},$
but $\sigma_{\gamma_{k}\cdot\mathbf{x}_{k}}\to\sigma_{\xi^{-}(\gamma_{+})}$ because $\xi^{+}(\gamma_{-})\pitchfork\mathbf{x}$, and $t_{k}+\mathcal{C}_{\sigma}(\mathbf{x}_{k},\gamma_{k})\to-\infty$ because of Corollary 7.10, so $\gamma_{k}\cdot\nu_{k}\to 0\in{\mathscr{D}}^{\mathbf{s}}\mathcal{F}^{-}_{\bullet}$. ∎
Remark 7.5.
The conclusion of Lemma 7.11 implies that $\gamma_{k}\cdot\nu_{k}\to\infty$ in ${\mathscr{D}}^{\mathbf{s}}_{>0}\mathcal{F}^{-}_{\bullet}$ as $k\to\infty$, in the sense that it leaves all compact subsets of the positive multidensity bundle ${\mathscr{D}}^{\mathbf{s}}_{>0}\mathcal{F}^{-}_{\bullet}$.
Finally we are in the position to prove Theorem 1.
Proof of Theorem 1.
The cocompactness of the action on $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}$ comes from Proposition 7.7, and any properly discontinuous action of a torsion-free group is free, so all that is left to prove is the proper discontinuity of the $\Gamma$-action on ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$.
By way of contradiction, suppose there exists $(\nu^{+}_{k},\nu^{-}_{k})\in\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ such that $(\nu^{+}_{k},\nu^{-}_{k})\to(\nu^{+},\nu^{-})\in\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ and $\delta_{k}\rightarrow\infty$ in $\Gamma$ with $\delta_{k}\cdot\nu_{k}\to(m^{+},m^{-})\in\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$. By [DMS24, Lemma 2.11], there exists $r>0$, $f\in\Gamma$ such that by passing to a subsequence we get $d_{\infty}((f\delta_{k})^{+},(f\delta_{k})^{-})\geq r$. Set $\gamma_{k}:=f\delta_{k}$ and $\mu^{\pm}:=f\cdot m^{\pm}$, so that $d_{\infty}(\gamma_{k}^{+},\gamma_{k}^{-})\geq r$ and $\gamma_{k}\cdot(\nu^{+}_{k},\nu^{-}_{k})\to(\mu^{+},\mu^{-})\in\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$.
Now, extract a further subsequence so that $\gamma_{k}$ converges to $\gamma_{+}\in\partial_{\infty}\Gamma$ and $\gamma_{k}^{-1}$ converges to $\gamma_{-}\in\partial_{\infty}\Gamma$ (note that $\gamma_{+}\neq\gamma_{-}$), and denote by $(\mathbf{x}^{+},\mathbf{x}^{-})\in{{\mathcal{F}}^{\pitchfork}}$ the base point of $(\nu^{+},\nu^{-})$. By virtue of $(\nu^{+},\nu^{-})\in\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$, either $\xi^{+}(\gamma_{-})\pitchfork\mathbf{x}^{-}$ or $\mathbf{x}^{+}\pitchfork\xi^{-}(\gamma_{-})$. The first case $\xi^{+}(\gamma_{-})\pitchfork\mathbf{x}^{-}$ implies via Lemma 7.11 and Proposition 7.8 that $\gamma_{k}\cdot\nu^{-}_{k}\rightarrow\infty$ in ${\mathscr{D}}^{\mathbf{s}}_{>0}{\mathcal{F}}^{-}_{\bullet}$, which is absurd because of the assumption $\gamma_{k}\cdot(\nu^{+}_{k},\nu^{-}_{k})\rightarrow(\mu^{+},\mu^{-})\in\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$. In the second case $\mathbf{x}^{+}\pitchfork\xi^{-}(\gamma_{-})$ we can repeat the same argument with $(\Theta,\mathbf{s})$ replaced by $(\iota_{o}(\Theta),\iota(\mathbf{s}))$, as is justified by (7.2), (7.5), and (7.9). This leads to the analogous contradiction, and therefore the $\Gamma$-action on ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$ is properly discontinuous. ∎
Remark 7.6.
Note that our proof of Theorem 1 never explicitly uses the coarsely linear growth defining the $\Theta$-Anosov property in Definition 7.1: the only step in which we use the growth of $\mathcal{P}_{\alpha}(\gamma_{k})$ is at the end of the proof of Proposition 7.8, where we only need that $\mathcal{P}_{\alpha}(\gamma_{k})\to+\infty$ as $k\to\infty.$ Subgroups $\Gamma<G$ satisfying this property are called $\Theta$-transverse, and were discussed in the introduction. In the proof of Lemma 7.9 we use in addition the existence of a pair of continuous, transverse, equivariant, dynamics-preserving boundary maps, and in total this implies the $\Theta$-Anosov property [GGKW17, Thm. 1.7].
8. Dynamics on the quotient manifold
Consider opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ associated to subsets $\Theta,\iota_{o}(\Theta)\subset\Delta$ respectively, ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ the transverse flag space, $\Gamma<G$ a torsion-free $\Theta$-Anosov subgroup, and $\mathbf{s}\in\mathbb{R}^{\Theta}$ a $(\Gamma,\Theta)$-admissible parameter. Then we know from Theorem 1 that $\Gamma$ acts properly discontinuously and freely on the $\phi^{t}$-invariant open set ~$\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}\subset\mathbb{L}^{\mathbf{s}}_{\Theta}$ and cocompactly on the $\phi^{t}$-invariant closed set $\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}\subset\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$. Since the $\Gamma$-actions on these sets commute with the flow $\phi^{t}$, the quotient manifold
${\mathcal{M}}_{\Gamma,\Theta}^{\mathbf{s}}:=\Gamma\backslash\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}}$
inherits a complete analytic flow, which we shall again denote by $\phi^{t}$. It preserves the compact set
$\mathcal{K}_{\Gamma,\Theta}^{\mathbf{s}}:=\Gamma\backslash\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}}\subset{\mathcal{M}}_{\Gamma,\Theta}^{\mathbf{s}},$
which we call the basic set. This terminology will be justified shortly.
8.1. Simplified notations
As before, elements of the flow space $\mathbb{L}^{\mathbf{s}}_{\Theta}$ will be denoted by
$(\mu^{+},\mu^{-})\in\mathbb{L}^{\mathbf{s}}_{\Theta}\subset\mathscr{D}^{\mathbf{s}}_{>0}\mathcal{F}_{\bullet}^{+}\boxplus\mathscr{D}^{\mathbf{s}}_{>0}\mathcal{F}_{\bullet}^{-}\,,$
and their projections on the base space ${{\mathcal{F}}^{\pitchfork}}$ will be written as
$\big{(}[\mu^{+}],[\mu^{-}]\big{)}\in{{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}\subset{\mathcal{F}}^{+}\times{\mathcal{F}}^{-}\,.$
To further ease the notation, we write from now on in this section
$\begin{split}{\mathbb{L}}:={\mathbb{L}}_{\Theta}^{\mathbf{s}},\qquad\widetilde{{\mathcal{M}}}:=\widetilde{{\mathcal{M}}}_{\Gamma,\Theta}^{\mathbf{s}},\qquad\widetilde{\mathcal{K}}:=\widetilde{\mathcal{K}}_{\Gamma,\Theta}^{\mathbf{s}},\qquad{\mathcal{M}}:={\mathcal{M}}_{\Gamma,\Theta}^{\mathbf{s}},\qquad\mathcal{K}:=\mathcal{K}_{\Gamma,\Theta}^{\mathbf{s}}.\end{split}$ (8.1)
8.2. Topological dynamics
Let us now describe the topological dynamics of the flow $\phi^{t}$ on ${\mathcal{M}}$. The following result and its proof are analogous to the projective Anosov case considered in [DMS24, Sec. 3.5]. There is some difference in the notation, so we repeat the arguments here for convenience.
Theorem 2 (C.f. [DMS24, Thm. 2]).
The following subsets of ${\mathcal{M}}$ coincide:
  • The basic set $\mathcal{K}$,
  • The non-wandering set $\mathcal{NW}(\phi^{t})$ of the flow $\phi^{t}$,
  • The closure of the set $\mathrm{Per}(\phi^{t})$ of periodic points of the flow $\phi^{t}$,
  • The trapped set ${\mathcal{T}}(\phi^{t})$ of the flow $\phi^{t}$,
  • The set $\overline{\bigcap_{t\in\mathbb{R}}\phi^{t}(U)}$ for any relatively compact open subset $U\subset{\mathcal{M}}$ containing $\mathcal{K}$.
Moreover, the restriction of the flow $\phi^{t}$ to $\mathcal{K}$ is topologically transitive.
The proof of Theorem 2 is given below after preparing some auxiliary lemmas closely resembling those in [DMS24, Sec. 3.5].
Lemma 8.1 (C.f. [DMS24, Lemma 3.10]).
Consider sequences $x_{k}\in{\mathcal{M}}$ and $t_{k}\to+\infty$ and points $x,y\in{\mathcal{M}}$ such that $x_{k}\to x\in{\mathcal{M}}$ and $\phi^{t_{k}}(x_{k})\to y\in{\mathcal{M}}$. Let $(\nu^{+},\nu^{-})\in\widetilde{\mathcal{M}}$ (resp. $(\mu^{+},\mu^{-})\in\widetilde{\mathcal{M}}$) be a lift of $x$ (resp. of $y\in{\mathcal{M}}$). Then $[\nu^{+}]\in\Lambda^{+}_{\Gamma}$ and $[\mu^{-}]\in\Lambda^{-}_{\Gamma}$.
Proof.
Consider lifts $(\nu^{+}_{k},\nu^{-}_{k})\in\widetilde{\mathcal{M}}$ of $x_{k}$ such that $(\nu^{+}_{k},\nu^{-}_{k})\to(\nu^{+},\nu^{-})$. There is a sequence $\delta_{k}\in\Gamma$ such that $\delta_{k}\cdot(e^{-t_{k}}\nu^{+}_{k},e^{t_{k}}\nu^{-}_{k})\to(\mu^{+},\mu^{-})$, and up to a subsequence, consider (applying [DMS24, Lemma 2.11] to the sequence $(\delta_{k}^{-1})$) some $f\in\Gamma$ such that the sequence $\gamma_{k}=f\delta_{k}^{-1}$ satisfies $\gamma_{k}\to\gamma_{+}\in\partial_{\infty}\Gamma$ and $\gamma_{k}^{-1}\to\gamma_{-}\in\partial_{\infty}\Gamma$ with $\gamma_{+}\neq\gamma_{-}$. Write $(\mu_{k}^{+},\mu^{-}_{k})=\delta_{k}\cdot(e^{-t_{k}}\nu^{+}_{k},e^{t_{k}}\nu^{-}_{k})$.
Assuming that $\xi^{+}(\gamma_{-})$ is transverse to $[\mu^{-}]$ leads to a contradiction: since $\mu_{k}^{-}\to\mu^{-}\neq 0$, Lemma 7.11 would imply that $\gamma_{k}\mu_{k}^{-}\to 0$, yet $\gamma_{k}\mu_{k}^{-}=e^{t_{k}}f\nu_{k}^{-}$ and $\nu_{k}^{-}\to\nu^{-}\neq 0$. Since $(\mu^{+},\mu^{-})\in\widetilde{\mathcal{M}}$, we must have $[\mu^{+}]\pitchfork\xi^{-}(\gamma_{-})$, therefore $\gamma_{k}[\mu^{+}_{k}]\to\xi^{+}(\gamma_{+})$ by (A3). This means that $f[\nu^{+}]=\xi^{+}(\gamma_{+})$, and in particular $[\nu^{+}]\in\Lambda^{+}_{\Gamma}$.
This implies that $\xi^{+}(\gamma_{+})\pitchfork f[\nu^{-}]$, so $\gamma_{k}^{-1}f[\nu^{-}_{k}]\to\xi^{-}(\gamma_{-})$, i.e. $[\mu^{-}]=\xi^{-}(\gamma_{-})\in\Lambda^{-}_{\Gamma}$. ∎
Lemma 8.2 (C.f. [DMS24, Lemma 3.5]).
Let $x\in{\mathcal{M}}$ and consider a lift $(\nu^{+},\nu^{-})\in\widetilde{\mathcal{M}}$. Then $x\in\mathrm{Per}(\phi^{t})$ if and only if there is $\gamma\in\Gamma\setminus\{e\}$ such that $[\nu^{+}]=\xi^{+}(\gamma^{+})$ and $[\nu^{-}]=\xi^{-}(\gamma^{-})$. [^3] Recall that we assume $\Gamma$ to be torsion-free here, so that any $\gamma\in\Gamma\setminus\{e\}$ has infinite order. In this case, if $\gamma$ is primitive, then the period of $x$ is $\mathcal{J}^{\mathbf{s}}_{\Theta}(\gamma)$.
Proof.
Assume that $x\in\mathrm{Per}(\phi^{t})$, and let $T>0$ be such that $\phi^{T}(x)=x$. There is $\gamma\in\Gamma\setminus\{e\}$ such that $(e^{-T}\nu^{+},e^{T}\nu^{-})=\gamma\cdot(\nu^{+},\nu^{-})$. If $[\nu^{-}]$ were to be transverse to $\xi^{+}(\gamma^{-})$, then Lemma 7.11 would imply that $\gamma^{k}\cdot\nu^{-}\to 0$ as $k\to+\infty$, which is a contradiction with $\gamma^{k}\cdot\nu^{-}=e^{kT}\nu^{-}$. As $(\nu^{+},\nu^{-})\in\widetilde{\mathcal{M}}$, we must have $[\nu^{+}]\pitchfork\xi^{-}(\gamma^{-})$, so $\gamma^{k}[\nu^{+}]\to\xi^{+}(\gamma^{+})$ as $k\to+\infty$, i.e. $[\nu^{+}]=\xi^{+}(\gamma^{+})$. Since $[\nu^{-}]\pitchfork[\nu^{+}]=\xi^{+}(\gamma^{+})$, we also find that $\gamma^{k}[\nu^{-}]\to\xi^{-}(\gamma^{-})$ as $k\to-\infty$, therefore $[\nu^{-}]=\xi^{-}(\gamma^{-})$.
If the assumption is now that $[\nu^{+}]=\xi^{+}(\gamma^{+})$ and $[\nu^{-}]=\xi^{-}(\gamma^{-})$ for some element $\gamma\in\Gamma\setminus\{e\}$, then combining (6.5), (3.21) and (3.19) yields $\phi^{w_{{\mathbf{s}},\Theta}(\lambda(\gamma))}(x)=x$. ∎
Proof of Theorem 2.
Any flow satisfies $\overline{\mathrm{Per}(\phi^{t})}\subset\mathcal{NW}(\phi^{t})$. Let $x\in\mathcal{NW}(\phi^{t})$, and consider sequences $x_{k}\to x$ in ${\mathcal{M}}$ and $t_{k}\to+\infty$ in $\mathbb{R}$ such that $\phi^{t_{k}}(x_{k})\to x$. If $(\nu^{+},\nu^{-})\in\widetilde{\mathcal{M}}$ is a lift of $x$, then Lemma 8.1 shows that $[\nu^{+}]\in\Lambda^{+}_{\Gamma}$ and $[\nu^{-}]\in\Lambda^{-}_{\Gamma}$, hence $x\in\mathcal{K}$, i.e. $\mathcal{NW}(\phi^{t})\subset\mathcal{K}$.
The set of poles of $\Gamma$, i.e. pairs $(\gamma^{+},\gamma^{-})$ of fixed points of elements $\gamma\in\Gamma\setminus\{e\}$, is dense in the space $\partial_{\infty}\Gamma^{(2)}$ of pairs of distinct points [Gro87, Corollary 8.2.G], hence $\mathcal{K}\subset\overline{\mathrm{Per}(\phi^{t})}$ thanks to Lemma 8.2.
Let $U\subset{\mathcal{M}}$ be a relatively compact subset containing $\mathcal{K}$, and consider $y\in\overline{\bigcap_{t\in\mathbb{R}}\phi^{t}(U)}$ and a lift $(\mu^{+},\mu^{-})\in\widetilde{\mathcal{M}}$. Consider first sequences $t_{k}\to+\infty$ and $x_{k}\in U$ such that $\phi^{t_{k}}(x_{k})\to y$, and assume without loss of generality that $x_{k}$ has a limit $x\in{\mathcal{M}}$. Lemma 8.1 shows that $[\mu^{-}]\in\Lambda^{-}_{\Gamma}$. Consider now sequences $s_{k}\to-\infty$ and $z_{k}\in U$ such that $y=\lim\phi^{s_{k}}(z_{k})$, and assume that $z_{k}\to z\in{\mathcal{M}}$. Applying Lemma 8.1 to $\phi^{-s_{k}}\left(\phi^{s_{k}}(y)\right)$ shows that $[\mu^{+}]\in\Lambda^{+}_{\Gamma}$, hence $y\in\mathcal{K}$.
The fact that $\mathcal{K}$ is the closure $\overline{\mathrm{Per}(\phi^{t})}$ of periodic points shows that $\mathcal{K}\subset\bigcap_{t\in\mathbb{R}}\phi^{t}(\mathcal{K})$, so finally $\overline{\bigcap_{t\in\mathbb{R}}\phi^{t}(U)}=\mathcal{K}$.
Since $\mathcal{K}$ is compact and $\phi^{t}$-invariant, we automatically get $\mathcal{K}\subset{\mathcal{T}}(\phi^{t})$. Let $x\in{\mathcal{T}}(\phi^{t})$ and consider a lift $(\nu^{+},\nu^{-})\in\widetilde{\mathcal{M}}$. Let $y\in{\mathcal{M}}$ be an $\omega$-limit point of $x$, i.e. $y=\lim_{k\to+\infty}\phi^{t_{k}}(x)$ for some sequence $t_{k}\to+\infty$ (it exists because $x$ is trapped). Then Lemma 8.1 shows that $[\nu^{+}]\in\Lambda^{+}_{\Gamma}$. The same reasoning applied to an $\alpha$-limit point of $x$ (i.e. some $\lim\phi^{s_{k}}(x)$ with $s_{k}\to-\infty$) shows that $[\nu^{-}]\in\Lambda^{-}_{\Gamma}$, hence $x\in\mathcal{K}$.
It remains to prove that the restriction of the flow $\phi^{t}$ to $\mathcal{K}$ is topologically transitive. This is again done by complete analogy to the projective case [DMS24, Lemma 3.12]: By [Gro87, 8.2.H], we may consider $(a,b)\in\partial_{\infty}\Gamma^{(2)}$ whose $\Gamma$-orbit is dense in $\partial_{\infty}\Gamma^{(2)}$. Let $x\in\mathcal{K}$ be an element admitting a lift $(\nu^{+},\nu^{-})\in\widetilde{\mathcal{K}}$ with $[\nu^{+}]=\xi(a)$ and $[\nu^{-}]=\xi^{*}(b)$. Let $y\in\mathcal{K}$ and consider a lift $(\mu^{+},\mu^{-})\in\widetilde{\mathcal{K}}$. Let $U\subset\mathcal{K}$ be an open subset containing $y$, and $\widetilde{U}\subset\widetilde{\mathcal{K}}$ its preimage. As the bundle projection $p:{\mathbb{L}}\to{{\mathcal{F}}^{\pitchfork}}$ is an open map, there is $\gamma\in\Gamma$ such that $\gamma\cdot([\mu^{+}],[\mu^{-}])\in p(\widetilde{U})$, i.e., there is $t\in\mathbb{R}$ such that $\gamma\cdot(e^{-t}\mu^{+},e^{t}\mu^{-})\in\widetilde{U}$, hence $\phi^{t}(x)\in U$. ∎
8.3. Hyperbolicity of the non-wandering set
We now turn to the differentiable dynamics of the flow $\phi^{t}$ on the manifold $\mathcal{M}$, and establish the last missing piece in the definition of Smale’s Axiom A.
Lemma 8.3.
The compact $\phi^{t}$-invariant subset $\mathcal{K}\subset\mathcal{M}$ is a hyperbolic set.
8.3.1. The original definition of Anosov subgroups
The proof of Lemma 8.3 is essentially a rephrasing of Labourie’s original definition just as in [DMS24, Lem. 3.13]. While Labourie worked with an Anosov flow on a closed manifold with a $\Gamma$-cover [Lab06, Sec. 3.1], Guichard and Wienhard used less regular flows [GW12, Def. 2.10], allowing us to work with Sambarino’s refraction flow $\psi^{t}_{\sigma}:\mathfrak{R}_{\sigma}\to\mathfrak{R}_{\sigma}$ [Sam24, Def. 5.3.4] (already discussed in Section 7.3). Consider the $\Gamma$-invariant maps
$f_{+}:\left\{\begin{array}[]{ccc}\partial_{\infty}\Gamma^{(2)}\times\mathbb{R}&\to&{\mathcal{F}}^{+}\\ (z^{+},z^{-},\tau)&\mapsto&\xi^{+}(z^{+})\end{array}\right.,\quad f_{-}:\left\{\begin{array}[]{ccc}\partial_{\infty}\Gamma^{(2)}\times\mathbb{R}&\to&{\mathcal{F}}^{-}\\ (z^{+},z^{-},\tau)&\mapsto&\xi^{-}(z^{-})\end{array}\right..$ (8.2)
The vector bundles $V^{\pm}:=\Gamma\backslash f_{\pm}^{*}T{\mathcal{F}}^{\pm}$ over the refraction flow space $\mathfrak{R}_{\sigma}$ each come equipped with a flow by vector bundle isomorphisms
$\psi^{t}_{\pm}:\left\{\begin{array}[]{ccc}V^{\pm}&\to&V^{\pm}\\ \Gamma\cdot((z^{+},z^{-},\tau),\mathbf{v})&\mapsto&\Gamma\cdot((z^{+},z^{-},t+\tau),\mathbf{v})\end{array}\right.$ (8.3)
lifting the refraction flow $\psi_{\sigma}^{t}$, i.e. such that the following diagrams commute:
[…]
The equivalence of Definition 7.1 with [GW12, Def. 2.10, Rem. 2.11], whose proof goes through [KP22] and [KLP18b, BPS19], gives the following property of $\Theta$-Anosov subgroups:
  1. (A4)
    The flow $\psi^{t}_{+}$ (resp. $\psi^{t}_{-}$) is uniformly dilating (resp. contracting).
8.3.2. The tangent bundle of the flow space
Given a point $\mu=(\mu^{+},\mu^{-})\in{\mathbb{L}}$, denote by $([\mu^{+}],[\mu^{-}])\in{{\mathcal{F}}^{\pitchfork}}$ its projection, and by $\mathfrak{n}^{+}\subset\mathfrak{g}$ (resp. $\mathfrak{n}^{-}\subset\mathfrak{g}$) the nilpotent radical of the parabolic Lie algebra of the stabilizer of $[\mu^{+}]\in{\mathcal{F}}^{+}$ (resp. $[\mu^{-}]\in{\mathcal{F}}^{-}$). Using once again the notation $\overline{X}(\mu)=\left.\frac{d\,}{dt}\right|_{t=0}\exp(tX)\cdot\mu$ for the fundamental vector field of $X\in\mathfrak{g}$, we get a grading of the tangent space $T_{\mu}{\mathbb{L}}$ as
$T_{\mu}{\mathbb{L}}=E^{+}_{\mu}\oplus E^{0}_{\mu}\oplus E^{-}_{\mu}$ (8.4)
where the vector subspaces $E^{+}_{\mu},E^{0}_{\mu},E^{-}_{\mu}\subset T_{\mu}{\mathbb{L}}$
$E^{+}_{\mu}=\left\{\overline{X}(\mu)\,\middle|\,X\in\mathfrak{n}^{-}\right\}~{},\qquad E^{0}_{\mu}=\mathbb{R}\cdot\left.\frac{d\,}{dt}\right|_{t=0}\phi^{t}(\mu)~{},\qquad E^{-}_{\mu}=\left\{\overline{X}(\mu)\,\middle|\,X\in\mathfrak{n}^{+}\right\}~{},$
define $G\times\{d\phi^{t}\}$-equivariant subbundles $E^{+},E^{0},E^{-}\subset T{\mathbb{L}}$.
Restricting this structure to ~$\widetilde{\mathcal{M}}\subset\mathbb{L}$ and taking the quotient by the $\Gamma$-action, the grading (8.4) descends yielding a flow invariant splitting
$T\mathcal{M}=E^{+}\oplus E^{0}\oplus E^{-}\,.$
The sign exchange between $E^{\pm}$ and $\mathfrak{n}^{\mp}$ is chosen so that the projections
$\pi_{\pm}:=p_{{\mathcal{F}}^{\pm}}\circ p:\mathbb{L}\to{\mathcal{F}}^{\pm}$
induce $G$-equivariant isomorphisms
$d\pi_{\pm}:E^{\pm}\to\pi_{\pm}^{*}T{\mathcal{F}}^{\pm}$ (8.5)
as $G$-equivariant vector bundles over $\mathbb{L}.$
8.3.3. Proof of hyperbolicity
We now have all the ingredients to prove Lemma 8.3.
Proof of Lemma 8.3.
Both sides in (8.5) carry a 1-parameter group of vector bundle isomorphisms: the restriction of $d\phi^{t}$ on $E^{\mp}$, and the the flow $\varphi_{\pm}^{t}:\pi_{\pm}^{*}T{\mathcal{F}}^{\pm}\to\pi_{\pm}^{*}T{\mathcal{F}}^{\pm}$ defined by
$\varphi_{\pm}^{t}:\left\{\begin{array}[]{ccc}\pi_{\pm}^{*}T{\mathcal{F}}^{\pm}&\to&\pi_{\pm}^{*}T{\mathcal{F}}^{\pm}\\ (x,\mathbf{v})&\mapsto&(\phi^{t}(x),\mathbf{v})\end{array}\right.$
The flow invariance $\pi_{\pm}\circ\phi^{t}=\pi_{\pm}$, when differentiated, shows that the isomorphism (8.5) intertwines the flows. The $G$-equivariant vector bundles $\pi_{\pm}^{*}T{\mathcal{F}}^{\pm}\to{\mathbb{L}}$ can be restricted to $\Gamma$-invariant bundles over ~$\widetilde{\mathcal{M}}\subset{\mathbb{L}}$ and descended to vector bundles $\mathcal{T}^{\pm}$ over $\mathcal{M}$ each equipped with a 1-parameter group of vector bundle isomorphisms still denoted by $\varphi^{t}_{\pm}$. Upon restriction to $\mathcal{K}\subset\mathcal{M}$, we obtain a commutative diagram
${E^{\pm}|_{\mathcal{K}}}$${\mathcal{T}^{\pm}|_{\mathcal{K}}}$${\mathcal{K}}$${\mathcal{K}}$ (8.6)
of vector bundle isomorphisms intertwining the flows $d\phi^{t}$ and $\varphi^{t}_{\pm}$ (because of the invariance $\pi_{\pm}\circ\phi^{t}=\pi_{\pm}$) given by descending $d\pi_{\pm}$ for the top horizontal arrow and the identity on the bottom.
The discussion of the refraction flow in Section 7.3 can be summarized in the commutative diagram
${\partial_{\infty}\Gamma^{(2)}\times\mathbb{R}}$${\widetilde{\mathcal{K}}}$${\widetilde{\mathcal{M}}}$${\mathfrak{R}_{\sigma}}$${\mathcal{K}}$${\mathcal{M}}$$\scriptstyle{\widetilde{\Phi}_{\sigma}}$$\scriptstyle{\Phi_{\sigma}}$
where the top row consists of $\Gamma$-spaces, vertical arrows are $\Gamma$-covers, unlabeled horizontal arrows are inclusions, and $\Phi_{\sigma}$ is a bi-Hölder conjugacy between the refraction flow on $\mathfrak{R}_{\sigma}$ and the restriction of the flow $\phi^{t}:\mathcal{M}\rightarrow\mathcal{M}$ to the compact invariant set $\mathcal{K}\subset\mathcal{M}.$
Now the $\Gamma$-equivariant map
$\left\{\begin{array}[]{ccc}f^{\ast}_{\pm}T{\mathcal{F}}^{\pm}&\to&\pi_{\pm}^{\ast}T{\mathcal{F}}^{\pm}|_{\widetilde{\mathcal{K}}}\\ \big{(}(z^{+},z^{-},\tau),\mathbf{v}\big{)}&\mapsto&\Big{(}\sigma\big{(}\xi^{\pitchfork}(z^{+},z^{-})\big{)},\mathbf{v}\Big{)}\end{array}\right.$
descends to a vector bundle isomorphism $V^{\pm}\to\mathcal{T}^{\pm}|_{\mathcal{K}}$ that conjugates $\psi^{t}_{\pm}$ to $\varphi^{t}_{\pm}$, so combining with (8.6) we get the following commutative diagram
[…]
of vector bundle isomorphisms conjugating the flows $\psi^{t}_{\pm}$, $\varphi^{t}_{\pm}$ and $d\phi^{t}$, so the hyperbolicity of $\mathcal{K}$ follows from (A4). ∎
9. Geometric structures on homogeneous spaces
In this section, we will explain the existence of various symplectic and contact structures on the homogeneous transverse flag and flow spaces considered in this article.
9.1. Geometry of transverse flag spaces
9.1.1. The universal para-Kähler structure
A para-Kähler structure on a smooth manifold $M$ of dimension $2k$ is a tuple $(\langle-,-\rangle,I,\omega)$ where $\langle$-,$\rangle$ - , - ⟩ is a neutral signature $(k,k)$ pseudo-Riemannian metric on $M$, the bundle homomorphism $I:TM\rightarrow TM$ is parallel for the Levi-Civita connection of $\langle$-,$\rangle$ - , - ⟩ and idempotent (i.e. $I^{2}=\mathrm{Id}$), and $\langle I(-),-\rangle=\omega$ is a symplectic form. It is the analog of a Kähler structure where the complex numbers are replaced by the paracomplex numbers $a+b\varepsilon$ where $a,b\in\mathbb{R}$ and $\varepsilon^{2}=1.$ The ±$\pm 1$-eigenspaces of $I:TM\rightarrow TM$ constitute a pair of transverse, integrable Lagrangian distributions for the symplectic form $\omega$, and the para-Kähler structure is equivalent to a symplectic manifold with a pair of transverse Lagrangian foliations.
[…]
$T_{\mathbf{x}}{{\mathcal{F}}^{\pitchfork}}=\mathcal{L}^{+}_{\mathbf{x}}\oplus\mathcal{L}^{-}_{\mathbf{x}}$ (9.1)
where the vector subspaces $\mathcal{L}^{\pm}_{\mathbf{x}}=\{\overline{X}(\mathbf{x})\,|\,X\in\mathfrak{n}^{\mp}\}$ define $G$-equivariant distributions $\mathcal{L}^{+},\mathcal{L}^{-}\subset T{{\mathcal{F}}^{\pitchfork}}$ tangent to the fibers of the projections $p_{{\mathcal{F}}^{\pm}}:{{\mathcal{F}}^{\pitchfork}}\to{\mathcal{F}}^{\pm}$. The $\mathrm{Stab}_{G}(\mathbf{x})$-equivariant isomorphism
$\left\{\begin{array}[]{ccc}\mathfrak{n}^{-}\oplus\mathfrak{n}^{+}&\to&T_{\mathbf{x}}{{\mathcal{F}}^{\pitchfork}}\\ X&\mapsto&\overline{X}(\mathbf{x})\end{array}\right.$ (9.4)
can be used to push forward the Killing form $\mathrm{B}_{\bf\mathfrak{g}}$ to a $G$-invariant pseudo-Riemannian metric $\langle-,-\rangle_{{{\mathcal{F}}^{\pitchfork}}}$ of signature $(n,n)$ on ${\mathcal{F}}^{\pitchfork}$ (where $n=\dim{\mathcal{F}}^{+}=\dim{\mathcal{F}}^{-}$) [^4] This is the same structure as defined in Section 3.3.
$\langle\overline{X}_{1}(\mathbf{x}),\overline{X}_{2}(\mathbf{x})\rangle_{{\mathcal{F}}^{\pitchfork}}:=\mathrm{B}_{\bf\mathfrak{g}}(X_{1},X_{2})\,,\quad X_{1},X_{2}\in\mathfrak{n}^{-}\oplus\mathfrak{n}^{+}$ (9.5)
for which the subbundles ±$\mathcal{L}^{\pm}\subset T{{\mathcal{F}}^{\pitchfork}}$ are totally isotropic. Note that since the Lie subalgebras $\mathfrak{n}^{+},\mathfrak{n}^{-}$ satisfy $[\mathfrak{n}^{+},\mathfrak{n}^{-}]\subset\mathfrak{l}$, where $\mathfrak{l}$ denotes the Lie algebra of the Levi stabilizer of $\mathbf{x}$, we have that
$\langle\overline{[X_{1},X_{2}]}(\mathbf{x}),\overline{X}_{3}(\mathbf{x})\rangle_{{\mathcal{F}}^{\pitchfork}}=\mathrm{B}_{\bf\mathfrak{g}}([X_{1},X_{2}],X_{3}),\qquad\forall\,X_{1},X_{2},X_{3}\in\mathfrak{n}^{-}\oplus\mathfrak{n}^{+}.$ (9.6)
Using the splitting (9.1), define the endomorphism of the tangent bundle $T{{\mathcal{F}}^{\pitchfork}}$
$I_{{\mathcal{F}}^{\pitchfork}}:\left\{\begin{array}[]{ccc}\mathcal{L}^{+}\oplus\mathcal{L}^{-}&\to&\mathcal{L}^{+}\oplus\mathcal{L}^{-}\\ \mathbf{v}^{+}+\mathbf{v}^{-}&\mapsto&\mathbf{v}^{+}-\mathbf{v}^{-}\end{array}\right.\,.$
Defining $\omega_{{\mathcal{F}}^{\pitchfork}}:=\langle I_{{\mathcal{F}}^{\pitchfork}}(-),\mathbf{-}\rangle_{{\mathcal{F}}^{\pitchfork}}$, we compute
$\omega_{{\mathcal{F}}^{\pitchfork}}\left(\mathbf{v}^{+}_{1}+\mathbf{v}^{-}_{1},\mathbf{v}^{+}_{2}+\mathbf{v}^{-}_{2}\right)=\langle\mathbf{v}^{+}_{1},\mathbf{v}^{-}_{2}\rangle_{{\mathcal{F}}^{\pitchfork}}-\langle\mathbf{v}^{-}_{1},\mathbf{v}^{+}_{2}\rangle_{{\mathcal{F}}^{\pitchfork}}\,,\quad\mathbf{v}^{+}_{1},\mathbf{v}^{+}_{2}\in\mathcal{L}^{+}\,,~{}\mathbf{v}^{-}_{1},\mathbf{v}^{-}_{2}\in\mathcal{L}^{-}\,,$
and find that $\omega_{{\mathcal{F}}^{\pitchfork}}$ is skew-symmetric and non-degenerate (because the Killing form induces a non-degenerate pairing $\mathrm{B}_{\bf\mathfrak{g}}:\mathfrak{n}^{+}\times\mathfrak{n}^{-}\to\mathbb{R}$). Cartan’s magic formula applied to (9.6) shows that $d\omega_{{\mathcal{F}}^{\pitchfork}}=0$, i.e. that $\omega_{{\mathcal{F}}^{\pitchfork}}$ is a symplectic form, so we have defined a $G$-invariant para-Kähler manifold $({\mathcal{F}}^{\pitchfork},\langle-,-\rangle_{{\mathcal{F}}^{\pitchfork}},I_{{\mathcal{F}}^{\pitchfork}},\omega_{{\mathcal{F}}^{\pitchfork}})$.
9.1.2. Kostant-Kirillov-Souriau symplectic structures
Given smooth functions $f,g:\mathfrak{g}^{*}\rightarrow\mathbb{R},$ the differentials at $\eta\in\mathfrak{g}^{*},$
$df_{\eta},dg_{\eta}:\mathfrak{g}^{*}\rightarrow\mathbb{R}$
may be understood as elements $df_{\eta},dg_{\eta}\in\mathfrak{g}$ through the canonical parallelism $T\mathfrak{g}^{*}\simeq\mathfrak{g}^{*}\times\mathfrak{g}^{*}$ and isomorphism $\left(\mathfrak{g}^{*}\right)^{*}\simeq\mathfrak{g}.$ The Poisson bracket
$\{f,g\}(\eta)=-\eta([df_{\eta},dg_{\eta}])\in\mathbb{R}$
defines the Kostant-Kirillov-Souriau (KKS) Poisson structure on $\mathfrak{g}^{*}$ which is well-defined for an arbitrary finite dimensional real Lie algebra $\mathfrak{g}.$ The coadjoint orbits $G\cdot\eta\subset\mathfrak{g}^{*}$ of elements $\eta\in{\bf\mathfrak{g}}^{\ast}$ are equal to the symplectic leaves of this Poisson structure [Kir04], and we denote by $\omega_{\eta}^{\mathrm{KKS}}$ the corresponding symplectic form on the coadjoint orbit $\mathcal{O}_{\eta}=G\cdot\eta\subset\mathfrak{g}^{*}.$ Evaluating the KKS symplectic form at $T_{\eta}\mathcal{O}_{\eta}\simeq{\bf\mathfrak{g}}/\mathfrak{l}_{\eta}$, where $\mathfrak{l}_{\eta}$ is the Lie algebra of $\mathrm{Stab}_{G}(\eta)$, it is given by
$\omega^{\mathrm{KKS}}_{\eta}(\overline{X_{1}}(\eta),\overline{X_{2}}(\eta))=-\eta([X_{1},X_{2}]),$ (9.7)
where on the left-hand side we consider fundamental vector fields for the coadjoint action.
Now, consider $X\in{\bf\mathfrak{g}}_{\mathrm{hyp}}$ associated to the transverse flag space ${\mathcal{F}}_{X}^{\pitchfork}:={\mathcal{F}}_{X}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}_{-X}$, and let $\eta:=\mathrm{B}_{\bf\mathfrak{g}}(X,-)\in{\bf\mathfrak{g}}^{\ast}$ be its Killing dual. Non-degeneracy of $\mathrm{B}_{\bf\mathfrak{g}}$ means that $\mathrm{Stab}_{G}(X)=\mathrm{Stab}_{G}(\eta)=L_{X}$, resulting in a $G$-equivariant diffeomorphism
$\left\{\begin{array}[]{ccc}\mathcal{O}_{\eta}=G\cdot\eta&\to&{\mathcal{F}}_{X}^{\pitchfork}\\ g\cdot\eta&\mapsto&g\cdot(\mathfrak{p}_{X},\mathfrak{p}_{-X})\end{array}\right.\,.$
Apart from some very specific cases, the push-forward of $\omega^{\mathrm{KKS}}$ on ${\mathcal{F}}_{X}$ is different from the symplectic form $\omega_{{\mathcal{F}}_{X}^{\pitchfork}}$ defined in the previous section, and choosing a different $X^{\prime}\in{\bf\mathfrak{g}}_{\mathrm{hyp}}$ with ${\mathcal{F}}_{X^{\prime}}^{\pitchfork}={\mathcal{F}}_{X}^{\pitchfork}$ results in different KKS symplectic forms. They however share the property that the distributions $\mathcal{L}^{\pm}$ of $T{{\mathcal{F}}^{\pitchfork}}$ are Lagrangian.
9.1.3. Period functions and closed 2-forms
Consider a transverse flag manifold ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$, and recall from Definition 4.6 its space $\mathcal{P}({{\mathcal{F}}^{\pitchfork}})$ of period functions of ${{\mathcal{F}}^{\pitchfork}}$, i.e. real valued $G$-invariant functions on $\mathrm{Stab}_{{\mathcal{F}}^{\pitchfork}}=\left\{(\mathbf{x},g)\in{{\mathcal{F}}^{\pitchfork}}\times G\,\middle|\,g\cdot\mathbf{x}=\mathbf{x}\right\}$ that restrict to a group homomorphism on each fiber.
Definition 9.1.
Let $\ell\in\mathcal{P}({{\mathcal{F}}^{\pitchfork}})$. The infinitesimal period function of $\ell$ is the function $d\ell:{{\mathcal{F}}^{\pitchfork}}\to{\bf\mathfrak{g}}^{*}$ defined in the following way: for $\mathbf{x}\in{{\mathcal{F}}^{\pitchfork}}$, denote by $\mathfrak{l}$ the Lie algebra of $\mathrm{Stab}_{G}(\mathbf{x})$, and let $d\ell(\mathbf{x})\in{\bf\mathfrak{g}}^{*}$ be defined by
$d\ell(\mathbf{x})(X)=\left\{\begin{array}[]{cl}d_{e}\ell(\mathbf{x},-)(X)&\textrm{if }X\in\mathfrak{l}\\ 0&\textrm{if }X\in\mathfrak{l}^{\perp}\end{array}\right.$
where the orthogonal is considered with respect to the Killing form.
The $G$-invariance of a period function $\ell\in\mathcal{P}({{\mathcal{F}}^{\pitchfork}})$ makes the infinitesimal period function $G$-equivariant, thus the range of $d\ell:{{\mathcal{F}}^{\pitchfork}}\to{\bf\mathfrak{g}}^{*}$ is a coadjoint orbit
$\mathcal{O}_{\ell}:=\left\{d\ell(\mathbf{x})\,\middle|\,\mathbf{x}\in{{\mathcal{F}}^{\pitchfork}}\right\}\subset{\bf\mathfrak{g}}^{*}\,.$
Proposition 9.2.
The assignment
$\begin{array}[]{ccc}{\mathcal{P}({{\mathcal{F}}^{\pitchfork}})}&\to&{\Omega^{2}({{\mathcal{F}}^{\pitchfork}})}\\ {\ell}&\mapsto&{\omega^{\ell}:=d\ell^{*}\omega^{\mathrm{KKS}}}\end{array}$
is a a linear injective map into closed $G$-invariant 2-forms.
Proof.
Closedness is a consequence of the closedness of $\omega^{\mathrm{KKS}}$, and $G$-invariance comes from the invariance of both $d\ell$ and $\omega^{\mathrm{KKS}}$. Linearity comes from (9.7):
$\omega^{\ell}_{\mathbf{x}}(\overline{X_{1}}(\mathbf{x}),\overline{X_{2}}(\mathbf{x}))=-d\ell(\mathbf{x})\big{(}[X_{1},X_{2}]\big{)}\,,\quad\forall X_{1},X_{2}\in{\bf\mathfrak{g}}\,.$
If $\omega^{\ell}=0$, then $\mathcal{O}_{\ell}$ consists of a single point, i.e. $\ell=0$. ∎
Definition 9.3.
A period function $\ell\in\mathcal{P}({{\mathcal{F}}^{\pitchfork}})$ is regular if $\omega^{\ell}$ is a symplectic form.
Note that a period function $\ell\in\mathcal{P}({{\mathcal{F}}^{\pitchfork}})$ is regular if and only if $d\ell:{{\mathcal{F}}^{\pitchfork}}\to\mathcal{O}_{\ell}$ is a diffeomorphism.
The coadjoint orbit $\mathcal{O}_{\ell}$ is always hyperbolic: the Killing dual of the linear form $d\ell(\mathbf{x})$ at any $\mathbf{x}\in{{\mathcal{F}}^{\pitchfork}}$ must lie in $\mathfrak{z}_{\mathrm{split}}(\mathfrak{l})$, because $d\ell(\mathbf{x})$ vanishes on the Lie algebra of the maximal compact torus of $\mathfrak{l}$ as well as on the derived algebra $[\mathfrak{l},\mathfrak{l}]$. In particular, $\mathcal{O}_{\ell}$ is always $G$-equivariantly diffeomorphic to some transverse flag space.
Remark 9.1.
If $\ell\in\mathcal{P}({{\mathcal{F}}^{\pitchfork}})$ is regular, then $\langle-,-\rangle_{\ell}:=\omega^{\ell}(I_{{\mathcal{F}}^{\pitchfork}}(-),-)$ is a pseudo-Riemannian metric of neutral signature, defining a (generally) different para-Kähler structure on ${{\mathcal{F}}^{\pitchfork}}$.
9.2. Geometry of flow spaces
We now consider a $G$-equivariant affine line bundle $p:{\mathbb{L}}\to{{\mathcal{F}}^{\pitchfork}}$, denote by $\ell\in\mathcal{P}({{\mathcal{F}}^{\pitchfork}})$ its period function, and consider the vertical vector field $\mathcal{X}_{\mathbb{L}}$ on ${\mathbb{L}}$ defined by
$\mathcal{X}_{\mathbb{L}}(\hat{\mathbf{x}})=\left.\frac{d\,}{dt}\right|_{t=0}\hat{\mathbf{x}}\cdot t\,,\quad\forall\hat{\mathbf{x}}\in{\mathbb{L}}\,.$
9.2.1. Pseudo-Riemannian structure
Recall the $G\times\{d\phi^{t}\}$-equivariant splitting
$T{\mathbb{L}}=E^{+}\oplus E^{0}\oplus E^{-}$
from Section 8.3.2, where $E^{0}$ is spanned by the derivative of the flow, and $E^{\pm}$ are lifts of the Lagrangian distributions $\mathcal{L}^{\pm}$ of $T{{\mathcal{F}}^{\pitchfork}}$. There is a unique $G$-invariant pseudo-Riemannian metric $\langle-,-\rangle_{\mathbb{L}}$ of signature $(n+1,n)$ on ${\mathbb{L}}$ with the following properties:
  1. (1)
    $\langle\mathcal{X}_{\mathbb{L}},\mathcal{X}_{\mathbb{L}}\rangle_{\mathbb{L}}\equiv 1\,,$
  2. (2)
    $E^{+}\oplus E^{-}\perp E^{0}\,,$
  3. (3)
    The projection $p:{\mathbb{L}}\to{{\mathcal{F}}^{\pitchfork}}$ is a pseudo-Riemannian submersion, i.e. its differential restricts to an isometry from $(E^{+}\oplus E^{-},\langle-,-\rangle_{\mathbb{L}})$ to $(T{{\mathcal{F}}^{\pitchfork}},\langle-,-\rangle_{{\mathcal{F}}^{\pitchfork}})$.
9.2.2. The canonical 1-form
A similar procedure constructs a 1-form on ${\mathbb{L}}$.
Definition 9.4.
The canonical 1-form $\tau\in\Omega^{1}({\mathbb{L}})$ is defined at $\hat{\mathbf{x}}\in{\mathbb{L}}$ by
$\tau_{\hat{\mathbf{x}}}\big{(}\mathcal{X}_{\mathbb{L}}(\hat{\mathbf{x}})\big{)}=1\textrm{ and }\tau_{\hat{\mathbf{x}}}(v)=0\textrm{ if }v\in E^{+}_{\hat{\mathbf{x}}}\oplus E^{-}_{\hat{\mathbf{x}}}\,.$
Note that $\tau=\langle\mathcal{X}_{\mathbb{L}},-\rangle_{\mathbb{L}}$ is invariant under the flow $\phi^{t}$, and it is non-zero on each fiber, so it can be seen as a principal connection on the principal $\mathbb{R}$-bundle $p:{\mathbb{L}}\to{{\mathcal{F}}^{\pitchfork}}$. Since $\mathbb{R}$ is abelian, the curvature of this connection is simply $d\tau$.
Proposition 9.5.
The canonical 1-form $\tau$ on ${\mathbb{L}}$ satisfies
$d\tau=p^{*}\omega^{\ell}\,.$
It is a contact form if and only if $\ell$ is regular.
Proof.
Fix a base point $\hat{\mathbf{x}}\in{\mathbb{L}}$, and consider the orbit map
$\varphi_{\hat{\mathbf{x}}}:\left\{\begin{array}[]{ccc}G&\to&{\mathbb{L}}\\ g&\mapsto&g\cdot\hat{\mathbf{x}}.\end{array}\right.$
The value of $\varphi_{\hat{\mathbf{x}}}^{*}\tau\in\Omega^{1}(G)$ at $e$ is
$\big{(}\varphi_{\hat{\mathbf{x}}}^{*}\tau\big{)}_{e}(X)=d\ell(\hat{\mathbf{x}})(X)\,,\qquad\forall X\in\mathfrak{g}\,.$
To simplify notations, write $\eta:=d\ell(\hat{\mathbf{x}})\in{\bf\mathfrak{g}}^{*}$. Since $\varphi_{\hat{\mathbf{x}}}^{*}\tau$ is a left-invariant 1-form on $G$, we have that
$\varphi_{\hat{\mathbf{x}}}^{*}\tau=\eta\circ\mathrm{m}$
where $\mathrm{m}\in\Omega^{1}(G,{\bf\mathfrak{g}})$ is the left-invariant Maurer-Cartan form (i.e. $\mathrm{m}_{g}(\mathbf{v})=(d_{e}L_{g})^{-1}(\mathbf{v})\in{\bf\mathfrak{g}}$ for $\mathbf{v}\in T_{g}G$). The Maurer-Cartan equation tells us that for any $\mathbf{v}_{1},\mathbf{v}_{2}\in T_{g}G$
$\big{(}\varphi_{\hat{\mathbf{x}}}^{*}d\tau\big{)}_{g}(\mathbf{v}_{1},\mathbf{v}_{2})=d\big{(}\varphi_{\hat{\mathbf{x}}}^{*}\tau\big{)}_{g}(\mathbf{v}_{1},\mathbf{v}_{2})=\big{(}\eta\circ d\mathrm{m}\big{)}_{g}(\mathbf{v}_{1},\mathbf{v}_{2})=-\eta\big{(}[\mathrm{m}_{g}(\mathbf{v}_{1}),\mathrm{m}_{g}(\mathbf{v}_{2})]\big{)}\,.$
If we now consider the orbit map
$\varphi_{\eta}:\left\{\begin{array}[]{ccc}G&\to&\mathcal{O}_{\eta}\\ g&\mapsto&\eta\circ\mathrm{Ad}(g^{-1}),\end{array}\right.$
then we find that
$\big{(}\varphi_{\hat{\mathbf{x}}}^{*}p^{*}\omega^{\ell}\big{)}_{g}(\mathbf{v}_{1},\mathbf{v}_{2})=\big{(}\varphi_{\eta}^{*}\omega^{\mathrm{KKS}}\big{)}_{g}(\mathbf{v}_{1},\mathbf{v}_{2})=-\eta\big{(}[\mathrm{m}_{g}(\mathbf{v}_{1}),\mathrm{m}_{g}(\mathbf{v}_{2})]\big{)}\,.$
By virtue of $\varphi_{\hat{\mathbf{x}}}$ being a surjective submersion, we find that $d\tau=p^{*}\omega^{\ell}$. In particular, $\tau$ is a contact form iff $\omega^{\ell}$ is a symplectic form, which is equivalent to $\ell$ being regular. ∎
9.3. Geometry of quotient flows
9.3.1. Regularity and multi-densities
Considering opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ corresponding to subsets $\Theta,\iota_{o}(\Theta)\subset\Delta$ and $\mathbf{s}\in\mathbb{R}^{\Theta}$, it is particularly easy to tell when the flow space ${\mathbb{L}}_{\Theta}^{\mathbf{s}}$ over ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ is regular, i.e. when its period function is regular.
Proposition 9.6.
The period function of the flow space ${\mathbb{L}}_{\Theta}^{\mathbf{s}}\to{{\mathcal{F}}^{\pitchfork}}$ is regular if and only if $\mathbf{s}_{\alpha}\neq 0$ for all $\alpha\in\Theta$.
Proof.
Let us fix a restricted root space decomposition, so that we are in the setting of Section 3.4. Fix some $\mathbf{x}\in{{\mathcal{F}}^{\pitchfork}}$ such that $\mathrm{Stab}_{G}(\mathbf{x})=L_{\Theta}$. By (6.5) and Lemma 3.30 we have that
$\ell_{{\mathbb{L}}_{\Theta}^{\mathbf{s}}}(\mathbf{x},\exp(X))=\sum_{\alpha\in\Theta}s_{\alpha}m_{\alpha}\mathrm{B}_{\bf\mathfrak{g}}(X_{\alpha},X)\,,\quad\forall X\in\mathfrak{l}_{\Theta}\,,$
where $X_{\alpha}$ is the co-root of $\alpha\in\Delta$ and $m_{\alpha}>0$. So $X:=\sum_{\alpha\in\Theta}s_{\alpha}m_{\alpha}X_{\alpha}$ is the Killing dual of $d\ell_{{\mathbb{L}}^{\mathbf{s}}_{\Theta}}(\mathbf{x})$, and the map $d\ell_{{\mathbb{L}}^{\mathbf{s}}_{\Theta}}:{{\mathcal{F}}^{\pitchfork}}\to\mathcal{O}_{\ell_{{\mathbb{L}}^{\mathbf{s}}_{\Theta}}}$ is simply the projection $G/L_{\Theta}\to G/L_{X}$, so the flow space ${\mathbb{L}}_{\Theta}^{\mathbf{s}}$ is regular if and only if $L_{X}=L_{\Theta}$. The result follows from $L_{X}=L_{\Theta^{\prime}}$ where $\Theta^{\prime}=\left\{\alpha\in\Theta\,\middle|\,s_{\alpha}\neq 0\right\}$. ∎
9.3.2. Contact realization of refraction flows
Theorem 3.
Consider some transverse flag space ${{\mathcal{F}}^{\pitchfork}}$ and the associated subset $\Theta\subset\Delta$. Let $\Gamma<G$ be a torsion-free $\Theta$-Anosov subgroup, and $\mathbf{s}\in\mathbb{R}^{\Theta}$ a $(\Gamma,\Theta)$-admissible parameter. If $s_{\alpha}\neq 0$ for all $\alpha\in\Theta$, then the flow $\phi^{t}$ on the quotient manifold $\mathcal{M}_{\Gamma,\Theta}^{\mathbf{s}}$ is the Reeb flow of a contact form.
If it is not the case, there is a subset $\Theta^{\prime}\subset\Theta$ and a $(\Gamma,\Theta^{\prime})$-admissible parameter $\mathbf{s}^{\prime}\in\mathbb{R}^{\Theta^{\prime}}$ such that the flow on $\mathcal{M}_{\Gamma,\Theta^{\prime}}^{\mathbf{s}^{\prime}}$ is the Reeb flow of a contact form and there is a flow-equivariant surjective submersion $\mathcal{M}_{\Gamma,\Theta}^{\mathbf{s}}\to\mathcal{M}_{\Gamma,\Theta^{\prime}}^{\mathbf{s}^{\prime}}$ that restricts to a homeomorphism $\mathcal{K}_{\Gamma,\Theta}^{\mathbf{s}}\to\mathcal{K}_{\Gamma,\Theta^{\prime}}^{\mathbf{s}^{\prime}}$.
Remark 9.2.
Note that a $\Theta$-Anosov subgroup is also $\Theta^{\prime}$-Anosov for any $\Theta^{\prime}\subset\Theta$. In particular, Theorem 3 tells us that every refraction flow can be realized as the restriction of a contact Axiom A flow to its non-wandering set.
Proof.
If $s_{\alpha}\neq 0$ for all $\alpha\in\Theta$, the result follows from Propositions 9.5 and 9.6. Now assume that $\Theta^{\prime}=\left\{\alpha\in\Theta\,\middle|\,s_{\alpha}\neq 0\right\}\neq\Theta$, set $\mathbf{s}^{\prime}=(s_{\alpha})_{\alpha\in\Theta^{\prime}}$ and consider the transverse flag space ${\mathcal{F}}_{\Theta^{\prime}}^{\pitchfork}={\mathcal{F}}_{\Theta^{\prime}}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}_{\iota_{o}(\Theta^{\prime})}$. Then $\Gamma$ is $\Theta^{\prime}$-Anosov, $\mathbf{s}^{\prime}$ is $(\Gamma,\Theta^{\prime})$-regular and by the first case the flow on the quotient manifold $\mathcal{M}^{\mathbf{s}^{\prime}}_{\Gamma,\Theta^{\prime}}$ is the Reeb flow of a contact form. The projection $\pi:{{\mathcal{F}}^{\pitchfork}}\to{\mathcal{F}}_{\Theta^{\prime}}^{\pitchfork}$ induces a $G\times\mathbb{R}$-invariant map
$\widetilde{f}:\left\{\begin{array}[]{ccc}{\mathbb{L}}^{\mathbf{s}}_{\Theta}&\to&{\mathbb{L}}^{\mathbf{s}^{\prime}}_{\Theta^{\prime}}\\ \big{(}(\mu^{+}_{\alpha})_{\alpha\in\Theta},(\mu^{-}_{\alpha})_{\alpha\in\Theta}\big{)}&\mapsto&\big{(}(\mu^{+}_{\alpha})_{\alpha\in\Theta^{\prime}},(\mu^{-}_{\alpha})_{\alpha\in\Theta^{\prime}}\big{)}\end{array}\right.$
making the following diagram commutative
${{\mathbb{L}}^{\mathbf{s}}_{\Theta}}$${{\mathbb{L}}_{\Theta^{\prime}}^{\mathbf{s}^{\prime}}}$${{{\mathcal{F}}^{\pitchfork}}}$${{\mathcal{F}}_{\Theta^{\prime}}^{\mathbf{s}^{\prime}}}$$\scriptstyle{p}$$\scriptstyle{\widetilde{f}}$$\scriptstyle{p^{\prime}}$$\scriptstyle{\pi}$
The transverse limit maps $\xi^{\pitchfork}:\partial_{\infty}\Gamma^{(2)}\to{{\mathcal{F}}^{\pitchfork}}$ and $\xi_{\Theta^{\prime}}^{\pitchfork}:\partial_{\infty}\Gamma^{(2)}\to{\mathcal{F}}_{\Theta^{\prime}}^{\pitchfork}$ satify $\xi_{\Theta^{\prime}}^{\pitchfork}=\pi\circ\xi^{\pitchfork}$ (because of Lemma 3.40). This implies that $\widetilde{f}$ maps ~$\widetilde{\mathcal{M}}^{\mathbf{s}}_{\Gamma,\Theta}$ onto ~$\widetilde{\mathcal{M}}^{\mathbf{s}^{\prime}}_{\Gamma,\Theta^{\prime}}$ and $\widetilde{\mathcal{K}}^{\mathbf{s}}_{\Gamma,\Theta}$ onto $\widetilde{\mathcal{K}}^{\mathbf{s}^{\prime}}_{\Gamma,\Theta^{\prime}}$, inducing a flow-equivariant surjective submersion $f:\mathcal{M}_{\Gamma,\Theta}^{\mathbf{s}}\to\mathcal{M}_{\Gamma,\Theta^{\prime}}^{\mathbf{s}^{\prime}}$ that maps $\mathcal{K}^{\mathbf{s}}_{\Gamma,\Theta}$ onto $\mathcal{K}^{\mathbf{s}^{\prime}}_{\Gamma,\Theta^{\prime}}$. This restriction is also injective because of the flow-equivariance, so it must be a homeomorphism by compactness of $\mathcal{K}^{\mathbf{s}}_{\Gamma,\Theta}$. ∎
10. Towers of homogeneous spaces
Given a $\Theta$-Anosov subgroup $\Gamma<G$, the space of $(\Gamma,\Theta)$-admissible parameters $\mathbf{s}\in\mathbb{R}^{\Theta}$ forms an open convex cone parameterizing a family of real analytic Axiom A flows. It is possible to assemble them all into a single dynamical system consisting of an action of the additive group $\mathbb{R}^{\Theta}$, also called a multiflow. If $\lvert\Theta\rvert\geq 2,$ the natural invariant set will be a non-compact principal $\mathbb{R}^{|\Theta|-1}$-bundle over the basic hyperbolic set $\mathcal{K}_{\Gamma,\Theta}^{\mathbf{s}}.$
10.1. The density multiflow space
Consider opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ associated to subsets $\Theta,\iota_{o}(\Theta)\subset\Delta$ respectively, and let ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ be the transverse flag space. Instead of considering multidensities on ${\mathcal{F}}^{+}$ and ${\mathcal{F}}^{-}$ as we did for the flow spaces in Section 6, we will keep track of individual densities on each simple flag manifold dominated by ${\mathcal{F}}^{+}$ and ${\mathcal{F}}^{-}$. Recall from Corollary 3.20 that for every $\alpha\in\Theta$ we have a $G$-equivariant projection $\pi_{\alpha}:{{\mathcal{F}}^{\pitchfork}}\to{\mathcal{F}}^{\pitchfork}_{\alpha}={\mathcal{F}}_{\alpha}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}_{\iota_{o}(\alpha)}$.
Definition 10.1.
The multiflow space over ${{\mathcal{F}}^{\pitchfork}}$ is the $\mathbb{R}^{\Theta}$-principal bundle
$\mathbb{W}_{\Theta}:=\prod_{\alpha\in\Theta}\pi_{\alpha}^{\ast}{\mathbb{L}}^{1}_{\{\alpha\}}\,,$
that is the bundle $\hat{p}:\mathbb{W}_{\Theta}\to{{\mathcal{F}}^{\pitchfork}}$ with fiber
$\big{(}\mathbb{W}_{\Theta}\big{)}_{\mathbf{x}}=\left\{(\nu_{\alpha})_{\alpha\in\Theta}\,\middle|\,\nu_{\alpha}\in\big{(}{\mathbb{L}}_{\{\alpha\}}^{1}\big{)}_{\pi_{\alpha}(\mathbf{x})}\right\}$
over $\mathbf{x}\in{{\mathcal{F}}^{\pitchfork}}$, and $\mathbb{R}^{\Theta}$-action defined by
$\big{(}\nu_{\alpha}\big{)}_{\alpha\in\Theta}\cdot\mathbf{t}=\big{(}\phi^{t_{\alpha}}(\nu_{\alpha})\big{)}_{\alpha\in\Theta}\,,\qquad\forall\,\mathbf{t}=(t_{\alpha})_{\alpha\in\Theta}\in\mathbb{R}^{\Theta}\,.$
Lemma 10.2.
For every $\mathbf{s}\in\mathbb{R}^{\Theta}$, there is a commutative diagram
${\mathbb{W}_{\Theta}}$${{\mathbb{L}}_{\Theta}^{\mathbf{s}}}$${{{\mathcal{F}}^{\pitchfork}}}$$\scriptstyle{\Pi_{\Theta}^{\mathbf{s}}}$$\scriptstyle{\hat{p}}$$\scriptstyle{p}$
of $G$-equivariant fiber bundles, where $\hat{p}$ and $p$ denote the projections, and $\Pi_{\Theta}^{\mathbf{s}}$ satisfies the equivariance property:
$\Pi_{\Theta}^{\mathbf{s}}(\nu)\cdot\mathbf{t}=\phi^{\mathbf{t}\cdot\mathbf{s}}\left(\Pi_{\Theta}^{\mathbf{s}}(\nu)\right)\,,\quad\forall\nu\in\mathbb{W}_{\Theta}\,\forall\,\mathbf{t}\in\mathbb{R}^{\Theta}\,,$
where $\mathbf{t}\cdot\mathbf{s}=\sum_{\alpha\in\Theta}t_{\alpha}s_{\alpha}$.
Proof.
The construction that we present here uses (multi)densities and is an exercise in associativity of various product expansions; an algebraic version on the level of homogeneous coset spaces can be inferred from the description of stabilizers in the next section.
Consider $\nu\in\mathbb{W}_{\Theta}$, let $\mathbf{x}=\hat{p}(\nu)\in{{\mathcal{F}}^{\pitchfork}}$ and write $\pi_{\alpha}(\mathbf{x})=(\mathbf{x}_{\alpha}^{+},\mathbf{x}_{\alpha}^{-})$ for $\alpha\in\Theta$. By definition, we have that $\nu=(\nu_{\alpha})_{\alpha\in\Theta}$ where $\nu_{\alpha}\in{\mathbb{L}}^{1}_{\{\alpha\}}$, that is $\nu_{\alpha}=(\nu_{\alpha}^{+},\nu_{\alpha}^{-})$ where $\nu_{\alpha}^{+}\in\mathscr{D}^{1}_{>0}\mathcal{F}_{\alpha}$ and $\nu_{\alpha}^{-}\in\mathscr{D}^{1}_{>0}\mathcal{F}_{\iota_{o}(\alpha)}$ are positive 1-densities with $\nu_{\alpha}^{+}\cdot\nu_{\alpha}^{-}=1$. In summary:
$\nu=(\nu_{\alpha})_{\alpha\in\Theta}=(\nu_{\alpha}^{+},\nu_{\alpha}^{-})_{\alpha\in\Theta}\,.$
From the $s_{\alpha}$-densities $(\nu_{\alpha}^{+})^{s_{\alpha}}\in\mathscr{D}^{s_{\alpha}}_{>0}{\mathcal{F}}_{\alpha}$ and $(\nu_{\alpha}^{-})^{s_{\alpha}}\in\mathscr{D}^{s_{\alpha}}_{>0}{\mathcal{F}}_{\iota_{o}(\alpha)}$ we assemble the multidensities $\nu^{\pm\mathbf{s}}:=\prod_{\alpha\in\Theta}(\nu_{\alpha}^{\pm})^{s_{\alpha}}\in\mathscr{D}^{\mathbf{s}}_{>0}{\mathcal{F}}^{\pm}_{\bullet}$, which satisfy $\nu^{+\mathbf{s}}\cdot\nu^{-\mathbf{s}}=1$, and set
$\Pi_{\Theta}^{\mathbf{s}}(\nu):=(\nu^{+\mathbf{s}},\nu^{-\mathbf{s}})\in{\mathbb{L}}_{\Theta}^{\mathbf{s}}\,.$
For $\mathbf{t}\in\mathbb{R}^{\Theta}$, expressing $\phi^{t_{\alpha}}(\nu_{\alpha})=(e^{-t_{\alpha}}\nu^{+}_{\alpha},e^{t_{\alpha}}\nu^{-}_{\alpha})$ we find
$\Pi_{\Theta}^{\mathbf{s}}(\nu\cdot\mathbf{t})$ $=\Big{(}\prod_{\alpha\in\Theta}(e^{-t_{\alpha}}\nu_{\alpha}^{+})^{s_{\alpha}},\prod_{\alpha\in\Theta}(e^{t_{\alpha}}\nu_{\alpha}^{-})^{s_{\alpha}}\Big{)}$
$=\Big{(}\prod_{\alpha\in\Theta}e^{-t_{\alpha}s_{\alpha}}(\nu_{\alpha}^{+})^{s_{\alpha}},\prod_{\alpha\in\Theta}e^{t_{\alpha}s_{\alpha}}(\nu_{\alpha}^{-})^{s_{\alpha}}\Big{)}$
$=\Big{(}e^{-\mathbf{t}\cdot\mathbf{s}}\prod_{\alpha\in\Theta}(\nu_{\alpha}^{+})^{s_{\alpha}},e^{\mathbf{t}\cdot\mathbf{s}}\prod_{\alpha\in\Theta}(\nu_{\alpha}^{-})^{s_{\alpha}}\Big{)}$
$=\phi^{\mathbf{t}\cdot\mathbf{s}}\left(\Pi_{\Theta}^{\mathbf{s}}(\nu)\right)\,.\vspace*{-1em}$
Corollary 10.3.
Suppose $\Gamma<G$ is $\Theta$-Anosov. Then $\Gamma$ acts properly discontinuously on the open subset $\hat{p}^{-1}(\Omega_{\Gamma})$.
Proof.
Consider some $(\Gamma,\Theta)$-admissible parameter $\mathbf{s}$ (e.g. $s_{\alpha}=1$ for every $\alpha$). Then $\Pi_{\Theta}^{\mathbf{s}}$ maps $\hat{p}^{-1}(\Omega_{\Gamma})$ onto ~$\widetilde{\mathcal{M}}^{\mathbf{s}}_{\Gamma,\Theta}$ in a $\Gamma$-equivariant way, so the proper discontinuity on ~$\widetilde{\mathcal{M}}^{\mathbf{s}}_{\Gamma,\Theta}$ implies the same property on $\hat{p}^{-1}(\Omega_{\Gamma})$. ∎
Remark 10.1.
As we will see in Section 10.3, the $G$-action on $\mathbb{W}_{\Delta}$ is proper, so any discrete subgroup $\Gamma<G$ acts properly discontinuously on $\mathbb{W}_{\Delta}$. This is not the case when $\Theta\neq\Delta$ and $G$ has rank at least two.
We can therefore consider the quotient manifold $\mathcal{N}_{\Gamma,\Theta}:=\Gamma\backslash\hat{p}^{-1}(\Omega_{\Gamma})$. The equivariance formula in Lemma 10.2 shows that $\mathcal{N}_{\Gamma,\Theta}$ (resp. $\widehat{\mathcal{K}}_{\Gamma,\Theta}:=\Gamma\backslash\hat{p}^{-1}(\Lambda^{\pitchfork}_{\Gamma})$) is a principal $\mathbf{s}^{\perp}\simeq\mathbb{R}^{\lvert\Theta\rvert-1}$-bundle over $\mathcal{M}^{\mathbf{s}}_{\Gamma,\Theta}$ (resp. $\mathcal{K}^{\mathbf{s}}_{\Gamma,\Theta}$). In particular, $\widehat{\mathcal{K}}_{\Gamma,\Theta}$ is never compact when $\lvert\Theta\rvert\geq 2$.
10.2. Algebraic description of homogeneous spaces
The multiflow space $\mathbb{W}_{\Theta}$ is a $G$-homogeneous space: given $\mathbf{x}\in{\mathcal{F}}_{\Theta}^{\pitchfork}$ with $L_{\Theta}=\mathrm{Stab}_{G}(\mathbf{x})$, the action of $g\in L_{\Theta}$ on the fiber over $\mathbf{x}$ is given by
$g\cdot(\nu_{\alpha})_{\alpha\in\Theta}=\big{(}e^{\mathbf{J}_{{{\mathcal{F}}^{\pitchfork}}}^{\alpha}(\mathbf{x},g)}\nu_{\alpha}\big{)}\,,\quad\forall(\nu_{\alpha})_{\alpha\in\Theta}\in\big{(}\mathbb{W}_{\Theta}\big{)}_{\mathbf{x}}\,,$
thus showing that the split center $Z_{\mathrm{split}}(L_{\Theta})$ of $L_{\Theta}=\mathrm{Stab}_{G}(\mathbf{x})$ acts transitively on $\big{(}\mathbb{W}_{\Theta}\big{)}_{\mathbf{x}}$. Recall the split central projection $\lambda_{L_{\Theta}}^{Z}\in\mathrm{Hom}(L_{\Theta},\mathfrak{z}_{\mathrm{split}}(\mathfrak{l}_{\Theta}))$ and its reductive kernel $M_{\Theta}=\ker\lambda_{L_{\Theta}}^{Z}$. We can identify the projection $\hat{p}$ with the natural projection
$\mathbb{W}_{\Theta}\simeq G/M_{\Theta}\longrightarrow G/L_{\Theta}\simeq{\mathcal{F}}_{\Theta}^{\pitchfork}\,.$
The intermediate homogeneous space ${\mathbb{L}}_{\Theta}^{\mathbf{s}}$ corresponds to $G/H_{\mathbf{b}}$ where $H_{\mathbf{b}}=\ker\mathbf{b}$ for some $\mathbf{b}\in\mathrm{Hom}(L_{\Theta},\mathbb{R})$, and the map $\Pi_{\Theta}^{\mathbf{s}}:\mathbb{W}_{\Theta}\to{\mathbb{L}}_{\Theta}^{\mathbf{s}}$ also corresponds to the natural map given by the chain of inclusions $M_{\Theta}<H_{\mathbf{b}}<L_{\Theta}$.
10.3. Connection with the Riemannian symmetric space
When $\Theta=\Delta$, i.e. when the flag manifolds ${\mathcal{F}}^{+}\simeq{\mathcal{F}}^{-}$ are isomorphic to quotients $G/P$ by a minimal parabolic subgroup $P<G$, the reductive subgroup $M=M_{L}<L$ in the Langlands decomposition $P=MAN$ is compact. In this setting, the action of $A\simeq\mathbb{R}^{\Delta}$ on $G/M$ is known as the Weyl chamber flow, and corresponds to the action on the space of parametrized flats of the Riemannian symmetric space $\mathbb{X}$ of $G$.
For a general $\Theta\subset\Delta$, consider a maximal compact subgroup $K_{\Theta}<L_{\Theta}$, and a maximal compact subgroup $K<G$ containing $K_{\Theta}$. Note that $K_{\Theta}<M_{\Theta}$, thus adding another homogeneous space in the chain of projections
$G/K_{\Theta}\to\mathbb{W}_{\Theta}\simeq G/M_{\Theta}\to{\mathbb{L}}_{\Theta}^{\mathbf{s}}\simeq G/H_{\mathbf{b}}\to{\mathcal{F}}_{\Theta}^{\pitchfork}\simeq G/L_{\Theta}\,.$
It is notable that only the initial and terminal spaces in this list have a known geometric interpretation within the Riemannian symmetric space $\mathbb{X}$. Indeed, the subgroup $K<G$ is the stabilizer of a point $x\in\mathbb{X}$, and the subgroup $K_{\Theta}<K$ is the stabilizer of a unit tangent vector $v\in T_{x}^{1}\mathbb{X}$. The geodesic $x_{v}(t)$ starting at $x$ with initial velocity $v$ has distinct endpoints $v^{\pm}=\lim_{t\to\pm\infty}x_{v}(t)$ in the visual boundary $\partial_{\infty}\mathbb{X}$. Their stabilizers $P_{\Theta}^{\pm}=\mathrm{Stab}_{G}(v^{\pm})$ are opposite parabolic subgroups with intersection $L_{\Theta}=\mathrm{Stab}_{G}(v^{+},v^{-})$.
${T^{1}\mathbb{X}\supset G\cdot v\simeq G/K_{\Theta}}$${\mathbb{W}_{\Theta}\simeq G/M_{\Theta}}$${\mathbb{X}\simeq G/K}$${{\mathbb{L}}_{\Theta}^{\mathbf{s}}\simeq G/H_{\mathbf{b}}}$${{\mathcal{F}}_{\Theta}^{\pitchfork}\simeq G/L_{\Theta}}$${\partial_{\infty}\mathbb{X}\supset G\cdot v^{-}\simeq{\mathcal{F}}_{\Theta}^{-}}$${{\mathcal{F}}_{\Theta}^{+}\simeq G\cdot v^{+}\subset\partial_{\infty}\mathbb{X}.}$
Remark 10.2.
This diagram exposes that the families $\mathbb{L}_{\Theta}^{\mathbf{s}}$ may be viewed as equivariant, inhomogeneous affine line bundles supported on the locally closed strata consisting of endpoints $(v^{+},v^{-})\in\partial_{\infty}\mathbb{X}\times\partial_{\infty}\mathbb{X}$ of geodesics of type $\Theta.$ Once $\Theta\subset\Delta$ is not fixed, one must face the stratified geometry of $\partial_{\infty}\mathbb{X}\times\partial_{\infty}\mathbb{X}$ which is a product of spherical buildings.
11. Examples
As is usual in constructions which operate within the language of roots systems and semisimple Lie algebras, consideration of explicit examples can dramatically simplify the nature of the objects under consideration. Therefore, the examples below will not mirror the general construction in Section 6 identically, but will produce isomorphic models of the flow spaces $\mathbb{L}^{\mathbf{s}}_{\Theta}$ constructed therein. We also recommend the first paper [DMS24] in our series which treats the real projective case by an independent method and construction, and also focuses in detail on the family of examples given by $\mathbb{H}^{p,q}$-convex cocompact subgroups. We explain our results in several other examples below.
11.1. Special linear groups over $\mathrm{k}=\mathbb{R},\mathbb{C}$
Consider a vector $\mathbf{d}=(d_{1},...,d_{r})\in\mathbb{N}^{r}$ with $d_{1}+\cdots+d_{r}=d$. Recall from Example 3.10 the space of gradings of type $\mathbf{d}$
${\mathcal{F}}_{\mathbf{d}}^{\pitchfork}(\mathrm{k}^{d})=\left\{E_{\bullet}=(E_{i})_{1\leq i\leq r}\in\prod_{i=1}^{r}\mathcal{G}_{d_{i}}(\mathrm{k}^{d})\,\middle|\,E_{1}\oplus\cdots\oplus E_{r}=\mathrm{k}^{d}\right\}\,,$
which corresponds to the coset space $\operatorname{SL}(d,\mathrm{k})/L_{\mathbf{d}}(\mathrm{k})$ where
$L_{\mathbf{d}}(\mathrm{k})=\left\{\begin{pmatrix}g_{1}&&0\\ &\ddots&\\ 0&&g_{r}\end{pmatrix}\,\middle|\,\mathrm{det}\,(g_{1})\cdots\mathrm{det}\,(g_{r})=1\right\}\subset\operatorname{SL}(d,\mathrm{k})\,.$
Consider the trivial $\mathrm{k}$-vector bundle
$\mathbb{E}_{\mathbf{d}}={\mathcal{F}}_{\mathbf{d}}^{\pitchfork}(\mathrm{k}^{d})\times\mathrm{k}^{d}\rightarrow{\mathcal{F}}_{\mathbf{d}}^{\pitchfork}(\mathrm{k}^{d}).$
The tautological grading
$\mathbb{E}_{\mathbf{d},\bullet}=\bigoplus_{i=1}^{r}\mathbb{E}_{\mathbf{d},i}$
is given at a basepoint $E_{\bullet}\in{\mathcal{F}}_{\mathbf{d}}^{\pitchfork}(\mathrm{k}^{d})$ by the grading $E_{\bullet}$ itself. [^5] This is the graded version of the usual tautological bundle construction in projective geometry. This construction yields an $\mathrm{SL}(d,\mathrm{k})$-equivariant graded vector bundle
$\mathbb{E}_{\mathbf{d},\bullet}\rightarrow{\mathcal{F}}_{\mathbf{d}}^{\pitchfork}(\mathrm{k}^{d}).$
Next, fix a weight vector $\mathbf{q}=(q_{1},...,q_{r})\in\mathbb{R}^{r}$ with $d_{1}q_{1}+\cdots+d_{r}q_{r}=0$ to produce the $\mathbf{q}$-multidensity bundle
$\mathscr{D}^{\mathbf{q}}\left(\mathbb{E}_{\mathbf{d},\bullet}\right)\rightarrow{\mathcal{F}}_{\mathbf{d}}^{\pitchfork}(\mathrm{k}^{d}).$
Regardless of whether $\mathrm{k}=\mathbb{R}$ or $\mathrm{k}=\mathbb{C},$ the multidensity construction treats the input vector bundle as an $\mathbb{R}$-vector bundle, and therefore the $\mathbf{q}$-multidensity bundle is an $\mathrm{SL}(d,\mathrm{k})$-equivariant, orientable real line bundle over ${\mathcal{F}}_{\mathbf{d}}^{\pitchfork}(\mathrm{k}^{d})$, that can also be described as $\operatorname{SL}(d,\mathrm{k})/H_{\mathbf{q}}$ where $H_{\mathbf{q}}<L_{\mathbf{d}}(\mathrm{k})$ is the kernel of $b_{\mathbf{q}}\in\mathrm{Hom}(L_{\mathbf{d}}(\mathrm{k}),\mathbb{R})$ given by
$b_{\mathbf{q}}\begin{pmatrix}g_{1}&&0\\ &\ddots&\\ 0&&g_{r}\end{pmatrix}=\sum_{i=1}^{r}q_{i}\log|\mathrm{det}\,(g_{i})|\,.$
The flow $\phi^{t}$ on $\mathscr{D}_{>0}^{\mathbf{q}}\left(\mathbb{E}_{\mathbf{d},\bullet}\right)$ consists in multiplication by $e^{t}$. On the level of the coset space $\mathrm{SL}(d,\mathrm{k})/H_{\mathbf{q}}$, it is given by right multiplication by the element
$\begin{pmatrix}e^{tq_{1}}\mathbf{1}_{d_{1}}&&0\\ &\ddots&\\ 0&&e^{tq_{r}}\mathbf{1}_{d_{r}}\end{pmatrix}\,.$
Remark 11.1.
In terms of the general construction from Section 6, we find
$\mathscr{D}_{>0}^{\mathbf{q}}\left(\mathbb{E}_{\mathbf{d},\bullet}\right)\simeq\mathrm{SL}(d,\mathrm{k})/H_{\mathbf{q}}\simeq\mathbb{L}_{\Theta_{\mathbf{d}}}^{\mathbf{s}(\mathbf{q})}.$
where $\mathbf{s}(\mathbf{q})=\frac{1}{d\dim_{\mathbb{R}}(\mathrm{k})}(q_{1}-q_{2},\ldots,q_{r-1}-q_{r})$, and $\Theta_{\mathbf{d}}$ was introduced in Example 3.22.
11.1.1. Multiflows
Finally, we discuss the multiflow spaces from Section 10. Consider the sum of line bundles
$\bigoplus_{i=1}^{r}\mathscr{D}^{1}(\mathbb{E}_{i,\mathbf{d}})\rightarrow{\mathcal{F}}_{\mathbf{d}}^{\pitchfork}(\mathrm{k}^{d}).$
In its total space, the open subset of positive $1$-densities $\bigoplus_{i=1}^{r}\mathscr{D}_{>0}^{1}(\mathbb{E}_{i,\mathbf{d}})$ is invariant by the fiber-wise transitive $\mathbb{R}^{r}$-action via
$(\tau_{i})_{1\leq i\leq r}\cdot(\nu_{i})_{1\leq i\leq r}=(e^{\tau_{i}}\nu_{i})_{1\leq i\leq r}.$
The trivial bundle $\mathbb{E}_{\mathbf{d}}={\mathcal{F}}_{\mathbf{d}}^{\pitchfork}(\mathrm{k}^{d})\times\mathrm{k}^{d}$ carries an $\mathrm{SL}(d,\mathrm{k})$-invariant fiberwise volume form $\omega.$ It can be used to define the product $\nu_{1}\cdots\nu_{r}\in\mathbb{R}$ of a collection $(\nu_{i})$ of 1-densities on the spaces $E_{i}$ of $E_{\bullet}\in{\mathcal{F}}_{\mathbf{d}}^{\pitchfork}(\mathrm{k}^{d})$ by letting
$\nu_{1}\cdots\nu_{r}:=\frac{\nu_{1}(\mathbf{e}_{1})\cdots\nu_{r}(\mathbf{e}_{r})}{|\omega(\mathbf{e}_{1},\dots,\mathbf{e}_{r})|}\,$
where each basis $\mathbf{e}_{i}=(e_{i,1},\dots,e_{i,d_{i}})\in\mathcal{B}_{E_{i}}$ is concatenated into $(\mathbf{e}_{1},\dots,\mathbf{e}_{r})\in\mathcal{B}_{E}$. This quantity is independent from the chosen bases, and we can now consider the codimension one submanifold
$\mathbb{W}_{\mathbf{d}}:=\left\{(\nu_{i})_{1\leq i\leq r}\in\bigoplus_{i=1}^{r}\mathscr{D}_{>0}^{1}(\mathbb{E}_{i,\mathbf{d}})\,\middle|\,\nu_{1}\cdots\nu_{r}=1\right\}\subset\bigoplus_{i=1}^{r}\mathscr{D}_{>0}^{1}(\mathbb{E}_{i,\mathbf{d}}).$
This manifold is invariant under the subgroup $\left\{\tau_{\bullet}\in\mathbb{R}^{r}\,\middle|\,d_{1}\tau_{1}+\cdots+d_{r}\tau_{r}=0\right\}\simeq\mathbb{R}^{r-1}$, and homogeneous under the $\mathrm{SL}(d,\mathrm{k})$-action, with stabilizer conjugate to
$M_{\mathbf{d}}(\mathrm{k})=\left\{\begin{pmatrix}g_{1}&&0\\ &\ddots&\\ 0&&g_{r}\end{pmatrix}\in L_{\mathbf{d}}(\mathrm{k})\,\middle|\,|\mathrm{det}\,(g_{i})|=1\right\}\subset L_{\mathbf{d}}(\mathrm{k})\,.$
Concerning the $\mathbb{R}^{r-1}$-action, it corresponds to multiplication on the right by the subgroup
$A_{\mathbf{d}}=\left\{\begin{pmatrix}e^{\tau_{1}}\mathbf{1}_{d_{1}}&&0\\ &\ddots&\\ 0&&e^{\tau_{d}}\mathbf{1}_{d_{r}}\end{pmatrix}\in\operatorname{SL}(d,\mathrm{k})\,\middle|\,\tau_{i}\in\mathbb{R}\right\}\,.$
Given a tuple $\mathbf{q}=(q_{1},...,q_{r}),$ the fiberwise monomial map
$\begin{split}\bigoplus_{i=1}^{r}\mathscr{D}^{1}_{>0}(\mathbb{E}_{i,\mathbf{d}})&\rightarrow\mathscr{D}^{\mathbf{q}}_{>0}(\mathbb{E}_{\bullet,\mathbf{d}})\\ (\nu_{i})_{1\leq i\leq r}&\mapsto\prod_{1\leq i\leq r}\ \nu_{i}^{q_{i}}\end{split}$
corresponds to the projection $\operatorname{SL}(d,\mathrm{k})/M_{\mathbf{d}}(\mathrm{k})\to\operatorname{SL}(d,\mathrm{k})/H_{\mathbf{q}}$.
11.2. Lagrangians in $(E,\omega)$
We next study the case of transverse Lagrangian subspaces in a symplectic real vector space. Let $E$ be a finite dimensional real vector space of dimension $2d\geq 2$ and let $\omega\in\Lambda^{2}E^{*}$ be a symplectic 2-form. The space of Lagrangian subspaces
$\mathrm{Lag}=\left\{L\in\mathcal{G}_{d}(E)\,\middle|\,\omega|_{L\times L}=0\right\}$
is a simple flag manifold for the symplectic group $\mathrm{Sp}(E,\omega)$. It is opposite to itself, and the corresponding transverse flag space is
$\mathrm{Lag}^{\pitchfork}=\left\{(L_{1},L_{2})\in\mathrm{Lag}\times\mathrm{Lag}\,\middle|\,L_{1}\oplus L_{2}=E\right\}\,.$
Consider $(L_{1},L_{2})\in\mathrm{Lag}^{\pitchfork}$, as well as positive 1-densities $\mu_{i}\in\mathscr{D}^{1}_{>0}(L_{i})$ (that is $\mu_{i}=|\omega_{i}|$ for some non-zero $\omega^{i}\in\Lambda^{d}L_{i}^{\ast}$). Using the symplectic volume instead of the pseudo-Riemannian volume used in Section 6, we can define the pairing $\mu_{1}\cdot\mu_{2}\in\mathbb{R}_{>0}$ as the quantity
$\mu_{1}\cdot\mu_{2}=\frac{\mu_{1}(\mathbf{e}_{1})\mu_{2}(\mathbf{e}_{2})}{\left|\mathrm{vol}_{\omega}(\mathbf{e}_{1},\mathbf{e}_{2})\right|}$
independent of the ordered bases $\mathbf{e}_{i}=(e_{i,1},\cdots,e_{i,d})\in\mathcal{B}_{L_{i}}$, where
$(\mathbf{e}_{1},\mathbf{e}_{2})=(e_{1,1},\cdots,e_{1,d},e_{2,1},\cdots,e_{2,d})\in\mathcal{B}_{E}$
stands for the concatenation and $\mathrm{vol}_{\omega}=\frac{1}{d!}\omega^{d}$. From there we can define the flow space ${\mathbb{L}}\to\mathrm{Lag}^{\pitchfork}$ with fiber
${\mathbb{L}}_{(L_{1},L_{2})}=\left\{(\mu_{1},\mu_{2})\in\mathscr{D}^{1}_{>0}(L_{1})\times\mathscr{D}^{1}_{>0}(L_{2})\,\middle|\,\mu_{1}\cdot\mu_{2}=1\right\}$
and endowed with the flow
$\phi^{t}(\mu_{1},\mu_{2})=(e^{-t}\mu_{1},e^{t}\mu_{2})\,.$
Now let $\Gamma<\mathrm{Sp}(E,\omega)$ be Anosov with respect to the simple restricted root defining the flag manifold $\mathrm{Lag}$ and torsion-free. As $\mathrm{Lag}$ is opposite to itself, we have a single limit map $\xi:\partial_{\infty}\Gamma\to\mathrm{Lag}$ and the corresponding invariant subsets
$\Lambda_{\Gamma}^{\pitchfork}$ =$=\left\{(\xi(z^{+}),\xi(z^{-}))\,\middle|\,z^{+}\neq z^{-}\right\}\subset\mathrm{Lag}^{\pitchfork}\,,$
$\Omega_{\Gamma}$ =$=\left\{(L_{1},L_{2})\in\mathrm{Lag}^{\pitchfork}\,\middle|\,\forall z\in\partial_{\infty}\Gamma~{}L_{1}\pitchfork\xi(z)\textrm{ or }\xi(z)\pitchfork L_{2}\right\}$
lift respectively to invariant subsets $\widetilde{\mathcal{K}}_{\Gamma}\subset\widetilde{\mathcal{M}}_{\Gamma}\subset{\mathbb{L}}$ where the action of $\Gamma$ on ~$\widetilde{\mathcal{M}}_{\Gamma}$ is properly discontinuous and the restriction to $\widetilde{\mathcal{K}}_{\Gamma}$ is cocompact, leading to a real analytic contact Axiom A flow on the quotient $\mathcal{M}_{\Gamma}=\Gamma\backslash\widetilde{\mathcal{M}}_{\Gamma}$ with non-wandering set $\mathcal{K}_{\Gamma}=\Gamma\backslash\widetilde{\mathcal{K}}_{\Gamma}$.
Remark 11.2.
It is possible to use non-zero volume forms $\omega_{i}\in\Lambda^{d}L_{i}^{*}$ instead of densities, but in the absence of a natural orientation of a Lagrangrian subspace, one has to consider them up to multiplication by $-1$. This leads either to this density construction, or to considering the Plücker embedding $\mathrm{Lag}\subset\mathcal{G}_{d}(E)\hookrightarrow\mathbb{P}(\Lambda^{d}E)$ then using the flow considered in [DMS24].
There is an injective homomorphism $h:\mathrm{SL}(2,\mathbb{R})\rightarrow\mathrm{Sp}(E,\omega)$ defined by
$\begin{pmatrix}a&b\\ c&d\end{pmatrix}\mapsto\begin{pmatrix}a\mathbf{1}_{d}&b\mathbf{1}_{d}\\ c\mathbf{1}_{d}&d\mathbf{1}_{d}\end{pmatrix}$
preserving a standard splitting $E\simeq L\oplus L^{*},$ and centralized by the compact subgroup
$\begin{pmatrix}A&0\\ 0&A\end{pmatrix}$
where $A\in\mathrm{O}(d).$ If $\Gamma_{0}<\mathrm{SL}(2,\mathbb{R})$ is a cocompact Fuchsian group and $\rho:\Gamma\rightarrow\mathrm{O}(d)$ is a homomorphism, the product $\gamma\mapsto\rho(\gamma)h(\gamma)$ defines a $\mathrm{Lag}^{\pitchfork}$-Anosov subgroup $\Gamma_{\rho}<\mathrm{Sp}(E,\omega)$ with the property that all of the associated Axiom A systems $\mathcal{M}_{\rho}^{s}$ have the same periods for fixed $s\in\mathbb{R}.$ Since the parameter space $\mathrm{Hom}(\Gamma,\mathrm{O}(d))$ is positive dimensional once $d>1,$ this shows that a general Anosov subgroup is not determined (up to conjugacy) by the periods of its associated Axiom A flows. This is in contrast to geodesic flows in negative curvature where length spectrum rigidity is a common phenomenon. The corresponding refraction flows are all conjugate, and therefore indistinguishable using classical Ergodic theory. In this case, microlocal objects (e.g. Ruelle resonant states and invariant distributions) should be helpful to distinguish the corresponding Axiom A systems.